AutoGen - The Automated Program Generator

For version 5.8, September 2006

Bruce Korb

bkorb@gnu.org

AutoGen copyright \begin{matrix}c \\ ◯ \end{matrix} 1992-2006 Bruce Korb

This is the second edition of the GNU AutoGen documentation,

Published by Bruce Korb, 910 Redwood Dr., Santa Cruz, CA 95060 AutoGen is free software. You may redistribute it and/or modify it under the terms of the GNU General Public License, as published by the Free Software Foundation; either version 2 of the License, or (at your option) any later version. AutoGen is distributed in the hope that it will be useful, but WITHOUT ANY WAR-RANTY; without even the implied warranty of MERCHANTABILITY or FITNESS FOR A PARTICULAR PURPOSE. See the GNU General Public License for more details. You should have received a copy of the GNU General Public License along with AutoGen. If not, write to: The Free Software Foundation, Inc., 51 Franklin Street, Fifth Floor Boston, MA 02110-1301, USA.

1 Introduction

AutoGen is a tool designed for generating program files that contain repetitive text with varied substitutions. Its goal is to simplify the maintenance of programs that contain large amounts of repetitious text. This is especially valuable if there are several blocks of such text that must be kept synchronized in parallel tables.

One common example is the problem of maintaining the code required for processing program options. Processing options requires a minimum of four different constructs be kept in proper order in different places in your program. You need at least:

1. The flag character in the flag string,

2. code to process the flag when it is encountered,

3. a global state variable or two, and

4. a line in the usage text.

You will need more things besides this if you choose to implement long option names, rc/ini/config file processing, environment variables and so on. All of this can be done mechanically; with the proper templates and this program. In fact, it has already been done and AutoGen itself uses it See Chapter 7 [AutoOpts], page 68. For a simple example of Automated Option processing, See Section 7.3 [Quick Start], page 71. For a full list of the Automated Option features, See Section 7.1 [Features], page 68.

1.1 The Purpose of AutoGen

The idea of this program is to have a text file, a template if you will, that contains the general text of the desired output file. That file includes substitution expressions and sections of text that are replicated under the control of separate definition files.

AutoGen was designed with the following features:

1. The definitions are completely separate from the template. By completely isolating the definitions from the template it greatly increases the flexibility of the template implementation. A secondary goal is that a template user only needs to specify those data that are necessary to describe his application of a template.

2. Each datum in the definitions is named. Thus, the definitions can be rearranged, augmented and become obsolete without it being necessary to go back and clean up older definition files. Reduce incompatibilities!

3. Every definition name defines an array of values, even when there is only one entry. These arrays of values are used to control the replication of sections of the template.

4. There are named collections of definitions. They form a nested hierarchy. Associated values are collected and associated with a group name. These associated data are used collectively in sets of substitutions.

5. The template has special markers to indicate where substitutions are required, much like the ${VAR} construct in a shell here doc. These markers are not fixed strings. They are specified at the start of each template. Template designers know best what fits into their syntax and can avoid marker conflicts. We did this because it is burdensome and difficult to avoid conflicts using either M4 tokenization or C preprocessor substitution rules. It also makes it easier to specify expressions that transform the value. Of course, our expressions are less cryptic than the shell methods.

6. These same markers are used, in conjunction with enclosed keywords, to indicate sections of text that are to be skipped and for sections of text that are to be repeated. This is a major improvement over using C preprocessing macros. With the C preprocessor, you have no way of selecting output text because it is an unvarying, mechanical substitution process.

7. Finally, we supply methods for carefully controlling the output. Sometimes, it is just simply easier and clearer to compute some text or a value in one context when its application needs to be later. So, functions are available for saving text or values for later use.

1.2 A Simple Example

This is just one simple example that shows a few basic features. If you are interested, you also may run "make check" with the VERBOSE environment variable set and see a number of other examples in the ‘agen5/test/testdir’ directory.

Assume you have an enumeration of names and you wish to associate some string with each name. Assume also, for the sake of this example, that it is either too complex or too large to maintain easily by hand. We will start by writing an abbreviated version of what the result is supposed to be. We will use that to construct our output templates. In a header file, ‘list.h’, you define the enumeration and the global array containing the associated strings:

     typedef enum {
             IDX_ALPHA,
             IDX_BETA,
             IDX_OMEGA }  list_enum;
     extern char const* az_name_list[ 3 ];

Then you also have ‘list.c’ that defines the actual strings:

     #include "list.h"
     char const* az_name_list[] = {
             "some alpha stuff",
             "more beta stuff",
             "final omega stuff" };

First, we will define the information that is unique for each enumeration name/string pair. This would be placed in a file named, ‘list.def’, for example.

     autogen definitions list;
     list = { list_element = alpha;
              list_info    = "some alpha stuff"; };
     list = { list_info    = "more beta stuff";
              list_element = beta; };
     list = { list_element = omega;
              list_info    = "final omega stuff"; };

The autogen definitions list; entry defines the file as an AutoGen definition file that uses a template named list. That is followed by three list entries that define the associations between the enumeration names and the strings. The order of the differently named elements inside of list is unimportant. They are reversed inside of the beta entry and the output is unaffected.

Now, to actually create the output, we need a template or two that can be expanded into the files you want. In this program, we use a single template that is capable of multiple output files. The definitions above refer to a ‘list’ template, so it would normally be named, ‘list.tpl’.

It looks something like this. (For a full description, See Chapter 3 [Template File], page 17.)

     [+ AutoGen5 template h c +]
     [+ CASE (suffix) +][+
        ==  h  +]
     typedef enum {[+
        FOR list "," +]
             IDX_[+ (string-upcase! (get "list_element")) +][+
        ENDFOR list +] }  list_enum;
     extern char const* az_name_list[ [+ (count "list") +] ];
     [+
        ==  c  +]
     #include "list.h"
     char const* az_name_list[] = {[+
       FOR list "," +]
             "[+list_info+]"[+
       ENDFOR list +] };[+
     ESAC +]

The [+ AutoGen5 template h c +] text tells AutoGen that this is an AutoGen version 5 template file; that it is to be processed twice; that the start macro marker is [+; and the end marker is +]. The template will be processed first with a suffix value of h and then with c. Normally, the suffix values are appended to the ‘base-name’ to create the output file name.

The [+ == h +] and [+ == c +] CASE selection clauses select different text for the two different passes. In this example, the output is nearly disjoint and could have been put in two separate templates. However, sometimes there are common sections and this is just an example.

The [+FOR list "," +] and [+ ENDFOR list +] clauses delimit a block of text that will be repeated for every definition of list. Inside of that block, the definition name-value pairs that are members of each list are available for substitutions.

The remainder of the macros are expressions. Some of these contain special expression functions that are dependent on AutoGen named values; others are simply Scheme expressions, the result of which will be inserted into the output text. Other expressions are names of AutoGen values. These values will be inserted into the output text. For example, [+list_info+] will result in the value associated with the name list_info being inserted

between the double quotes and (string-upcase! (get "list_element")) will first "get"

the value associated with the name list_element, then change the case of all the letters to upper case. The result will be inserted into the output document.

If you have compiled AutoGen, you can copy out the template and definitions as described above and run autogen list.def. This will produce exactly the hypothesized desired output.

One more point, too. Lets say you decided it was too much trouble to figure out how to use AutoGen, so you created this enumeration and string list with thousands of entries. Now, requirements have changed and it has become necessary to map a string containing the enumeration name into the enumeration number. With AutoGen, you just alter the template to emit the table of names. It will be guaranteed to be in the correct order, missing none of the entries. If you want to do that by hand, well, good luck.

1.3 csh/zsh caveat

AutoGen tries to use your normal shell so that you can supply shell code in a manner you are accustomed to using. If, however, you use csh or zsh, you cannot do this. Csh is sufficiently difficult to program that it is unsupported. Zsh, though largely programmable, also has some anomalies that make it incompatible with AutoGen usage. Therefore, when invoking AutoGen from these environments, you must be certain to set the SHELL environment variable to a Bourne-derived shell, e.g., sh, ksh or bash.

Any shell you choose for your own scripts need to follow these basic requirements:

1. It handles trap $sig ":" without output to standard out. This is done when the server shell is first started. If your shell does not handle this, then it may be able to by loading functions from its start up files.

2. At the beginning of each scriptlet, the command \\cd $PWD is inserted. This ensures that cd is not aliased to something peculiar and each scriptlet starts life in the execution directory.

3. At the end of each scriptlet, the command echo mumble is appended. The program you use as a shell must emit the single argument mumble on a line by itself.

1.4 A User’s Perspective

Alexandre wrote:

>

> I’d appreciate opinions from others about advantages/disadvantages of > each of these macro packages.

I am using AutoGen in my pet project, and find one of its best points to be that it separates the operational data from the implementation.

Indulge me for a few paragraphs, and all will be revealed: In the manual, Bruce cites the example of maintaining command line flags inside the source code; traditionally spreading usage information, flag names, letters and processing across several functions (if not files). Investing the time in writing a sort of boiler plate (a template in AutoGen terminology) pays by moving all of the option details (usage, flags names etc.) into a well structured table (a definition file if you will), so that adding a new command line option becomes a simple matter of adding a set of details to the table.

So far so good! Of course, now that there is a template, writing all of that tedious optargs processing and usage functions is no longer an issue. Creating a table of the options needed for the new project and running AutoGen generates all of the option processing code in C automatically from just the tabular data. AutoGen in fact already ships with such a template... AutoOpts.

One final consequence of the good separation in the design of AutoGen is that it is retargetable to a greater extent. The egcs/gcc/fixinc/inclhack.def can equally be used (with different templates) to create a shell script (inclhack.sh) or a c program (fixincl.c).

This is just the tip of the iceberg. AutoGen is far more powerful than these examples might indicate, and has many other varied uses. I am certain Bruce or I could supply you with many and varied examples, and I would heartily recommend that you try it for your project and see for yourself how it compares to m4.

As an aside, I would be interested to see whether someone might be persuaded to rationalise autoconf with AutoGen in place of m4... Ben, are you listening? autoconf-3.0! ‘kay? =)O| Sincerely,

Gary V. Vaughan

2 Definitions File

This chapter describes the syntax and semantics of the AutoGen definition file. In order to instantiate a template, you normally must provide a definitions file that identifies itself and contains some value definitions. Consequently, we keep it very simple. For "advanced" users, there are preprocessing directives, sparse arrays, named indexes and comments that may be used as well.

The definitions file is used to associate values with names. Every value is implicitly an array of values, even if there is only one value. Values may be either simple strings or compound collections of name-value pairs. An array may not contain both simple and compound members. Fundamentally, it is as simple as:

     prog-name = "autogen";
     flag = {
         name      = templ_dirs;
         value     = L;
         descrip   = "Template search directory list";
     };

For purposes of commenting and controlling the processing of the definitions, C-style comments and most C preprocessing directives are honored. The major exception is that the #if directive is ignored, along with all following text through the matching #endif directive. The C preprocessor is not actually invoked, so C macro substitution is not performed.

2.1 The Identification Definition

The first definition in this file is used to identify it as a AutoGen file. It consists of the two keywords, ‘autogen’ and ‘definitions’ followed by the default template name and a terminating semi-colon (;). That is:

             AutoGen Definitions template-name;

Note that, other than the name template-name, the words ‘AutoGen’ and ‘Definitions’ are searched for without case sensitivity. Most lookups in this program are case insensitive.

Also, if the input contains more identification definitions, they will be ignored. This is done so that you may include (see Section 2.5 [Directives], page 10) other definition files without an identification conflict.

AutoGen uses the name of the template to find the corresponding template file. It searches for the file in the following way, stopping when it finds the file:

1. It tries to open ‘./template-name’. If it fails,

 2. it tries ‘./template-name.tpl’.

3. It searches for either of these files in the directories listed in the templ-dirs command line option.

If AutoGen fails to find the template file in one of these places, it prints an error message and exits.

2.2 Named Definitions

Any name may have multiple values associated with it in the definition file. If there is more than one instance, the only way to expand all of the copies of it is by using the FOR (see Section 3.6.13 [FOR], page 49) text function on it, as described in the next chapter.

There are two kinds of definitions, ‘simple’ and ‘compound’. They are defined thus (see Section 2.9 [Full Syntax], page 14):

     compound_name '=' '{' definition-list '}' ';'
     simple_name '=' string ';'
     no_text_name ';'

No_text_name is a simple definition with a shorthand empty string value. The string values for definitions may be specified in any of several formation rules.

2.2.1 Definition List

definition-list is a list of definitions that may or may not contain nested compound definitions. Any such definitions may only be expanded within a FOR block iterating over the containing compound definition. See Section 3.6.13 [FOR], page 49.

Here is, again, the example definitions from the previous chapter, with three additional name value pairs. Two with an empty value assigned (first and last), and a "global" group name.

     autogen definitions list;
     group_name = example;
     list = { list_element = alpha;  first;
              list_info    = "some alpha stuff"; };
     list = { list_info    = "more beta stuff";
              list_element = beta; };
     list = { list_element = omega;  last;
              list_info    = "final omega stuff"; };

2.2.2 Double Quote String

The string follows the C-style escaping (\, \n, \f, \v, etc.), plus octal character numbers specified as \ooo. The difference from "C" is that the string may span multiple lines. Like ANSI "C", a series of these strings, possibly intermixed with single quote strings, will be concatenated together.

2.2.3 Single Quote String

This is similar to the shell single-quote string. However, escapes \ are honored before another escape, single quotes ' and hash characters #. This latter is done specifically to disambiguate lines starting with a hash character inside of a quoted string. In other words,

     fumble = '
     #endif
     ';

could be misinterpreted by the definitions scanner, whereas this would not:

     fumble = '
     \#endif
     ';

As with the double quote string, a series of these, even intermixed with double quote strings, will be concatenated together.

2.2.4 Shell Output String

This is assembled according to the same rules as the double quote string, except that there is no concatenation of strings and the resulting string is written to a shell server process. The definition takes on the value of the output string.

NB The text is interpreted by a server shell. There may be left over state from previous server shell processing. This scriptlet may also leave state for subsequent processing. However, a cd to the original directory is always issued before the new command is issued.

2.2.5 An Unquoted String

A simple string that does not contain white space may be left unquoted. The string must not contain any of the characters special to the definition text (i.e., ", #, ', (, ), ,, ;, <, =, >, [, ], `, {, or }). This list is subject to change, but it will never contain underscore (_), period (.), slash (/), colon (:), hyphen (-) or backslash (\\). Basically, if the string looks like it is a normal DOS or UNIX file or variable name, and it is not one of two keywords (‘autogen’ or ‘definitions’) then it is OK to not quote it, otherwise you should.

2.2.6 Scheme Result String

A scheme result string must begin with an open parenthesis (. The scheme expression will be evaluated by Guile and the value will be the result. The AutoGen expression functions are disabled at this stage, so do not use them.

2.2.7 A Here String

A ‘here string’ is formed in much the same way as a shell here doc. It is denoted with a doubled less than character and, optionally, a hyphen. This is followed by optional horizontal white space and an ending marker-identifier. This marker must follow the syntax rules for identifiers. Unlike the shell version, however, you must not quote this marker. The resulting string will start with the first character on the next line and continue up to but not including the newline that precedes the line that begins with the marker token. No backslash or any other kind of processing is done on this string. The characters are copied directly into the result string.

Here are two examples:

     str1 = <<-  STR_END
             $quotes = " ' `
             STR_END;
     str2 = <<   STR_END
             $quotes = " ' `
             STR_END;

     STR_END;

The first string contains no new line characters. The first character is the dollar sign, the last the back quote.

The second string contains one new line character. The first character is the tab character preceding the dollar sign. The last character is the semicolon after the STR_END. That STR_ END does not end the string because it is not at the beginning of the line. In the preceding case, the leading tab was stripped.

2.2.8 Concatenated Strings

If single or double quote characters are used, then you also have the option, a la ANSI-C syntax, of implicitly concatenating a series of them together, with intervening white space ignored.

NB You cannot use directives to alter the string content. That is,

     str = "fumble"
     #ifdef LATER
           "stumble"
     #endif
           ;

will result in a syntax error. The preprocessing directives are not carried out by the C preprocessor. However,

     str = '"fumble\n"
     #ifdef LATER
     "     stumble\n"
     #endif
     ';

Will work. It will enclose the ‘#ifdef LATER’ and ‘#endif’ in the string. But it may also wreak havoc with the definition processing directives. The hash characters in the first column should be disambiguated with an escape \ or join them with previous lines:

"fumble\n#ifdef LATER....

2.3 Assigning an Index to a Definition

In AutoGen, every name is implicitly an array of values. When assigning values, they are usually implicitly assigned to the next highest slot. They can also be specified explicitly:

     mumble[9] = stumble;
     mumble[0] = grumble;

If, subsequently, you assign a value to mumble without an index, its index will be 10, not 1. If indexes are specified, they must not cause conflicts.

#define-d names may also be used for index values. This is equivalent to the above:

     #define FIRST 0
     #define LAST  9
     mumble[LAST]  = stumble;
     mumble[FIRST] = grumble;

All values in a range do not have to be filled in. If you leave gaps, then you will have a sparse array. This is fine (see Section 3.6.13 [FOR], page 49). You have your choice of iterating over all the defined values, or iterating over a range of slots. This:

     [+ FOR mumble +][+ ENDFOR +]

iterates over all and only the defined entries, whereas this:

     [+ FOR mumble (for-by 1) +][+ ENDFOR +]

will iterate over all 10 "slots". Your template will likely have to contain something like this:

     [+ IF (exist? (sprintf "mumble[%d]" (for-index))) +]

or else "mumble" will have to be a compound value that, say, always contains a "grumble" value:

     [+ IF (exist? "grumble") +]

2.4 Dynamic Text

There are several methods for including dynamic content inside a definitions file. Three of them are mentioned above (Section 2.2.4 [shell-generated], page 8 and see Section 2.2.6 [scheme-generated], page 8) in the discussion of string formation rules. Another method uses the #shell processing directive. It will be discussed in the next section (see Section 2.5 [Directives], page 10). Guile/Scheme may also be used to yield to create definitions.

When the Scheme expression is preceded by a backslash and single quote, then the expression is expected to be an alist of names and values that will be used to create AutoGen definitions. This method can be be used as follows:

     \'( (name  (value-expression))
         (name2 (another-expr))  )

This is entirely equivalent to:

     name  = (value-expression);
     name2 = (another-expr);

Under the covers, the expression gets handed off to a Guile function named alist->autogen-def in an expression that looks like this:

     (alist->autogen-def
         ( (name (value-expression))  (name2 (another-expr)) ) )

2.5 Controlling What Gets Processed

Definition processing directives can only be processed if the ’#’ character is the first character on a line. Also, if you want a ’#’ as the first character of a line in one of your string assignments, you should either escape it by preceding it with a backslash ‘\’, or by embedding it in the string as in "\n#".

All of the normal C preprocessing directives are recognized, though several are ignored. There is also an additional #shell - #endshell pair. Another minor difference is that AutoGen directives must have the hash character (#) in column 1.

The final tweak is that #! is treated as a comment line. Using this feature, you can use: ‘#! /usr/local/bin/autogen’ as the first line of a definitions file, set the mode to executable and "run" the definitions file as if it were a direct invocation of AutoGen. This was done for its hack value.

The ignored directives are: ‘#ident’, ‘#let’, ‘#pragma’, and ‘#if’. Note that when ignoring the #if directive, all intervening text through its matching #endif is also ignored, including the #else clause.

The AutoGen directives that affect the processing of definitions are:

#assert `shell-script` | (scheme-expr) | <anything else>

If the shell-script or scheme-expr do not yield true valued results, autogen will be aborted. If <anything else> or nothing at all is provided, then this directive is ignored. When writing the shell script, remember this is on a preprocessing line. Multiple lines must be backslash continued and the result is a single long line. Separate multiple commands with semi-colons. The result is false (and fails) if the result is empty, the number zero, or a string that starts with the letters ’n’ or ’f’ ("no" or "false").

#define name [ <text> ]

Will add the name to the define list as if it were a DEFINE program argument. Its value will be the first non-whitespace token following the name. Quotes are not processed. After the definitions file has been processed, any remaining entries in the define list will be added to the environment.

#elif

This must follow an #if otherwise it will generate an error. It will be ignored.

#else

This must follow an #if, #ifdef or #ifndef. If it follows the #if, then it will be ignored. Otherwise, it will change the processing state to the reverse of what it was.

#endif

This must follow an #if, #ifdef or #ifndef. In all cases, this will resume normal processing of text.

#endmac

This terminates a "macdef", but must not ever be encountered directly.

#endshell

Ends the text processed by a command shell into autogen definitions.

#error [ <descriptive text> ]

This directive will cause AutoGen to stop processing and exit with a status of EXIT FAILURE.

#if [ <ignored conditional expression> ]

#if expressions are not analyzed. Everything from here to the matching #endif is skipped.

#ifdef name-to-test

The definitions that follow, up to the matching #endif will be processed only if there is a corresponding -Dname command line option or if a #define of that name has been previously encountered.

#ifndef name-to-test

The definitions that follow, up to the matching #endif will be processed only if there is not a corresponding -Dname command line option or there was a canceling -Uname option.

#include unadorned-file-name

This directive will insert definitions from another file into the current collection. If the file name is adorned with double quotes or angle brackets (as in a C program), then the include is ignored.

#line

Alters the current line number and/or file name. You may wish to use this directive if you extract definition source from other files. getdefs uses this mechanism so AutoGen will report the correct file and approximate line number of any errors found in extracted definitions.

#macdef

This is a new AT&T research preprocessing directive. Basically, it is a multi-line #define that may include other preprocessing directives.

#option opt-name [ <text> ]

This directive will pass the option name and associated text to the AutoOpts optionLoadLine routine (see Section 7.6.28.8 [libopts-optionLoadLine], page 104). The option text may span multiple lines by continuing them with a backslash. The backslash/newline pair will be replaced with two space characters. This directive may be used to set a search path for locating template files For example, this:

     #option templ-dirs $ENVVAR/dirname

will direct autogen to use the ENVVAR environment variable to find a directory named dirname that (may) contain templates. Since these directories are searched in most recently supplied first order, search directories supplied in this way will be searched before any supplied on the command line.

#shell

Invokes $SHELL or ‘/bin/sh’ on a script that should generate AutoGen definitions. It does this using the same server process that handles the back-quoted ` text. CAUTION let not your $SHELL be csh.

#undef name-to-undefine

Will remove any entries from the define list that match the undef name pattern.

2.6 Pre-defined Names

When AutoGen starts, it tries to determine several names from the operating environment and put them into environment variables for use in both #ifdef tests in the definitions files and in shell scripts with environment variable tests. __autogen__is always defined. For other names, AutoGen will first try to use the POSIX version of the sysinfo(2) system call. Failing that, it will try for the POSIX uname(2) call. If neither is available, then only "__autogen__" will be inserted into the environment. In all cases, the associated names are converted to lower case, surrounded by doubled underscores and non-symbol characters are replaced with underscores.

With Solaris on a sparc platform, sysinfo(2) is available. The following strings are used:

• SI_SYSNAME (e.g., " sunos ")

  • SI_HOSTNAME (e.g., " ellen ")
  • SI_ARCHITECTURE (e.g., " sparc ")

• SI_HW_PROVIDER (e.g., " sun microsystems ")

• SI_PLATFORM (e.g., " sun ultra 5 10 ")

• SI_MACHINE (e.g., " sun4u ")

For Linux and other operating systems that only support the uname(2) call, AutoGen will use these values:

• sysname (e.g., " linux ")

• machine (e.g., " i586 ")

• nodename (e.g., " bach ")

By testing these pre-defines in my definitions, you can select pieces of the definitions without resorting to writing shell scripts that parse the output of uname(1). You can also segregate real C code from autogen definitions by testing for "__autogen__".

     #ifdef __bach__
       location = home;
     #else
       location = work;
     #endif

2.7 Commenting Your Definitions

The definitions file may contain C and C++ style comments.

     /*
      *  This is a comment.  It continues for several lines and closes
      *  when the characters '*' and '/' appear together.
      */
     // this comment is a single line comment

2.8 What it all looks like.

This is an extended example:

     autogen definitions `template-name';
     /*
      *  This is a comment that describes what these
      *  definitions are all about.

*/
global = "value for a global text definition.";
/*
*  Include a standard set of definitions
*/
#include standards.def
a_block = {
a_field;
a_subblock = {
    sub_name  = first;
    sub_field = "sub value.";
};
#ifdef FEATURE
a_subblock = {
    sub_name  = second;
};
#endif
};

2.9 Finite State Machine Grammar

The preprocessing directives and comments are not part of the grammar. They are handled by the scanner/lexer. The following was extracted directly from the generated defParsefsm.c source file. The "EVT:" is the token seen, the "STATE:" is the current state and the entries in this table describe the next state and the action to take. Invalid transitions were removed from the table.

dp_trans_table[ DP_STATE_CT ][ DP_EVENT_CT ] = {
/* STATE 0:  DP_ST_INIT */
{ { DP_ST_NEED_DEF, NULL },                       /* EVT:  autogen */
/* STATE 1:  DP_ST_NEED_DEF */
{ DP_ST_NEED_TPL, NULL },                       /* EVT:  definitions */
/* STATE 2:  DP_ST_NEED_TPL */
{ DP_ST_NEED_SEMI, &dp_do_tpl_name },           /* EVT:  var_name */
{ DP_ST_NEED_SEMI, &dp_do_tpl_name },           /* EVT:  other_name */
{ DP_ST_NEED_SEMI, &dp_do_tpl_name },           /* EVT:  string */
/* STATE 3:  DP_ST_NEED_SEMI */
{ DP_ST_NEED_NAME, NULL },                      /* EVT:  ; */
/* STATE 4:  DP_ST_NEED_NAME */

{ { DP_ST_NEED_DEF, NULL },                       /* EVT:  autogen */
{ DP_ST_DONE, &dp_do_need_name_end },           /* EVT:  End-Of-File */
{ DP_ST_HAVE_NAME, &dp_do_need_name_var_name }, /* EVT:  var_name */
{ DP_ST_HAVE_VALUE, &dp_do_end_block },         /* EVT:  } */
/* STATE 5:  DP_ST_HAVE_NAME */
{ DP_ST_NEED_NAME, &dp_do_empty_val },          /* EVT:  ; */
{ DP_ST_NEED_VALUE, &dp_do_have_name_lit_eq },  /* EVT:  = */
{ DP_ST_NEED_IDX, NULL },                       /* EVT:  [ */
/* STATE 6:  DP_ST_NEED_VALUE */
{ DP_ST_HAVE_VALUE, &dp_do_str_value },         /* EVT:  var_name */
{ DP_ST_HAVE_VALUE, &dp_do_str_value },         /* EVT:  other_name */
{ DP_ST_HAVE_VALUE, &dp_do_str_value },         /* EVT:  string */
{ DP_ST_HAVE_VALUE, &dp_do_str_value },         /* EVT:  here_string */
{ DP_ST_HAVE_VALUE, &dp_do_str_value },         /* EVT:  number */
{ DP_ST_NEED_NAME, &dp_do_start_block },        /* EVT:  { */
/* STATE 7:  DP_ST_NEED_IDX */
{ DP_ST_NEED_CBKT, &dp_do_indexed_name },       /* EVT:  var_name */
{ DP_ST_NEED_CBKT, &dp_do_indexed_name },       /* EVT:  number */
/* STATE 8:  DP_ST_NEED_CBKT */
{ DP_ST_INDX_NAME, NULL }                       /* EVT:  ] */
/* STATE 9:  DP_ST_INDX_NAME */
{ DP_ST_NEED_NAME, &dp_do_empty_val },          /* EVT:  ; */
{ DP_ST_NEED_VALUE, NULL },                     /* EVT:  = */
/* STATE 10:  DP_ST_HAVE_VALUE */
{ DP_ST_NEED_NAME, NULL },                      /* EVT:  ; */
{ DP_ST_NEED_VALUE, &dp_do_next_val },          /* EVT:  , */

2.10 Alternate Definition Forms

There are several methods for supplying data values for templates.

‘no definitions’

It is entirely possible to write a template that does not depend upon external definitions. Such a template would likely have an unvarying output, but be convenient nonetheless because of an external library of either AutoGen or Scheme functions, or both. This can be accommodated by providing the - -override-tpl and --no-definitions options on the command line. See Chapter 5 [autogen Invocation], page 56.

‘CGI’ AutoGen behaves as a CGI server if the definitions input is from stdin and the environment variable REQUEST_METHOD is defined and set to either "GET" or "POST", See Section 6.2 [AutoGen CGI], page 65. Obviously, all the values are constrained to strings because there is no way to represent nested values.

‘XML’ AutoGen comes with a program named, xml2ag. Its output can either be redirected to a file for later use, or the program can be used as an AutoGen wrapper. See Section 8.6 [xml2ag Invocation], page 152. The introductory template example (see Section 1.2 [Example Usage], page 2) can be rewritten in XML as follows:

     <EXAMPLE  template="list.tpl">
     <LIST list_element="alpha"
           list_info="some alpha stuff"/>
     <LIST list_info="more beta stuff"
           list_element="beta"/>
     <LIST list_element="omega"
           list_info="final omega stuff"/>
     </EXAMPLE>

A more XML-normal form might look like this:

     <EXAMPLE  template="list.tpl">
     <LIST list_element="alpha">some alpha stuff</LIST>
     <LIST list_element="beta" >more beta stuff</LIST>
     <LIST list_element="omega">final omega stuff</LIST>
     </EXAMPLE>

but you would have to change the template list_info references into text references.

‘standard AutoGen definitions’

Of course. :-)

3 Template File

The AutoGen template file defines the content of the output text. It is composed of two parts. The first part consists of a pseudo macro invocation and commentary. It is followed by the template proper.

This pseudo macro is special. It is used to identify the file as a AutoGen template file, fixing the starting and ending marks for the macro invocations in the rest of the file, specifying the list of suffixes to be generated by the template and, optionally, the shell to use for processing shell commands embedded in the template.

AutoGen-ing a file consists of copying text from the template to the output file until a start macro marker is found. The text from the start marker to the end marker constitutes the macro text. AutoGen macros may cause sections of the template to be skipped or processed several times. The process continues until the end of the template is reached. The process is repeated once for each suffix specified in the pseudo macro.

This chapter describes the format of the AutoGen template macros and the usage of the AutoGen native macros. Users may augment these by defining their own macros, See Section 3.6.4 [DEFINE], page 48.

3.1 Format of the Pseudo Macro

The pseudo macro is used to tell AutoGen how to process a template. It tells autogen:

1. The start macro marker. It consists of punctuation characters used to demarcate the start of a macro. It may be up to seven characters long and must be the first nonwhitespace characters in the file. It is generally a good idea to use some sort of opening bracket in the starting macro and closing bracket in the ending macro (e.g. {, (, [, or even < in the starting macro). It helps both visually and with editors capable of finding a balancing parenthesis.

2. That start marker must be immediately followed by the identifier strings "AutoGen5" and then "template", though capitalization is not important.

The next several components may be intermingled:

3. Zero, one or more suffix specifications tell AutoGen how many times to process the template file. No suffix specifications mean that it is to be processed once and that the generated text is to be written to stdout. The current suffix for each pass can be determined with the (suffix) scheme function (see Section 3.4.38 [SCM suffix], page 29). The suffix specification consists of a sequence of POSIX compliant file name characters and, optionally, an equal sign and a file name formatting specification. That specification may be either an ordinary sequence of file name characters with zero, one or two "%s" formatting sequences in it, or else it may be a Scheme expression that, when evaluated, produces such a string. The two string arguments allowed for that string are the base name of the definition file, and the current suffix (that being the text to the left of the equal sign). (Note: "POSIX compliant file name characters" consist of alphanumerics plus the period (.), hyphen (-) and underscore (_) characters.) If the suffix begins with one of these three latter characters and a formatting string is not specified, then that character is presumed to be the suffix separator. Otherwise, without a specified format string, a single period will separate the suffix from the base name in constructing the output file name.

4. Comments: blank lines, lines starting with a hash mark [#]), and edit mode comments (text between pairs of -*- strings).

5. Scheme expressions may be inserted in order to make configuration changes before template processing begins. It is used, for example, to allow the template writer to specify the shell program that must be used to interpret the shell commands in the template. It can have no effect on any shell commands in the definitions file, as that file will have been processed by the time the pseudo macro is interpreted.

     (setenv "SHELL" "/bin/sh")

This is extremely useful to ensure that the shell used is the one the template was written to use. By default, AutoGen determines the shell to use by user preferences. Sometimes, that can be the "csh", though. The scheme expression can also be used to save a pre-existing output file for later text extraction (see Section 3.5.7 [SCM extract], page 32).

     (shellf "mv -f %1$s.c %1$s.sav" (base-name))

After these must come the end macro marker:

6. The punctuation characters used to demarcate the end of a macro. Like the start marker, it must consist of seven or fewer punctuation characters.

The ending macro marker has a few constraints on its content. Some of them are just advisory, though. There is no special check for advisory restrictions.

• It must not begin with a POSIX file name character (hyphen -, underscore _or period .), the backslash (\) or open parenthesis ((). These are used to identify a suffix specification, indicate Scheme code and trim white space.

• If it begins with an equal sign, then it must be separated from any suffix specification by white space.

• The closing marker may not begin with an open parenthesis, as that is used to enclose a scheme expression.

• It cannot begin with a backslash, as that is used to indicate white space trimming after the end macro mark. If, in the body of the template, you put the backslash character (\) before the end macro mark, then any white space characters after the mark and through the newline character are trimmed.

• It is also helpful to avoid using the comment marker (#). It might be seen as a comment within the pseudo macro.

• You should avoid using any of the quote characters double, single or back-quote. It won’t confuse AutoGen, but it might well confuse you and/or your editor.

As an example, assume we want to use [+ and +] as the start and end macro markers, and we wish to produce a ‘.c’ and a ‘.h’ file, then the pseudo macro might look something like this:

 [+ AutoGen5 template -*- Mode: emacs-mode-of-choice -*-
 h=chk-%s.h

     # make sure we don't use csh:
     (setenv "SHELL" "/bin/sh")  +]

The template proper starts after the pseudo-macro. The starting character is either the first non-whitespace character or the first character after the newline that follows the end macro marker.

3.2 Naming a value

When an AutoGen value is specified in a template, it is specified by name. The name may be a simple name, or a compound name of several components. Since each named value in AutoGen is implicitly an array of one or more values, each component may have an index associated with it. It looks like this:

     comp-name-1 . comp-name-2 [ 2 ]

Note that if there are multiple components to a name, each component name is separated by a dot (.). Indexes follow a component name, enclosed in square brackets ([ and ]). The index may be either an integer or an integer-valued define name. The first component of the name is searched for in the current definition level. If not found, higher levels will be searched until either a value is found, or there are no more definition levels. Subsequent components of the name must be found within the context of the newly-current definition level. Also, if the named value is prefixed by a dot (.), then the value search is started in the current context only. No higher levels are searched.

If someone rewrites this, I’ll incorporate it. :-)

3.3 Macro Expression Syntax

AutoGen has two types of expressions: full expressions and basic ones. A full AutoGen expression can appear by itself, or as the argument to certain AutoGen built-in macros: CASE, IF, ELIF, INCLUDE, INVOKE (explicit invocation, see Section 3.6.16 [INVOKE], page 51), and WHILE. If it appears by itself, the result is inserted into the output. If it is an argument to one of these macros, the macro code will act on it sensibly.

You are constrained to basic expressions only when passing arguments to user defined macros, See Section 3.6.4 [DEFINE], page 48.

The syntax of a full AutoGen expression is:

     [[ <apply-code> ] <value-name> ] [ <basic-expr-1> [ <basic-expr-2> ]]

How the expression is evaluated depends upon the presence or absence of the apply code and value name. The "value name" is the name of an AutoGen defined value, or not. If it does not name such a value, the expression result is generally the empty string. All expressions must contain either a value-name or a basic-expr.

3.3.1 Apply Code

The "apply code" selected determines the method of evaluating the expression. There are five apply codes, including the non-use of an apply code.

‘no apply code’

This is the most common expression type. Expressions of this sort come in three flavors:

‘<value-name>’

The result is the value of value-name, if defined. Otherwise it is the empty string.

‘<basic-expr>’

The result of the basic expression is the result of the full expression, See Section 3.3.2 [basic expression], page 20.

‘<value-name> <basic-expr>’

If there is a defined value for value-name, then the basic-expr is evaluated. Otherwise, the result is the empty string.

‘% <value-name> <basic-expr>’

If value-name is defined, use basic-expr as a format string for sprintf. Then, if the basic-expr is either a back-quoted string or a parenthesized expression, then hand the result to the appropriate interpreter for further evaluation. Otherwise, for single and double quote strings, the result is the result of the sprintf operation. Naturally, if value-name is not defined, the result is the empty string. For example, assume that fumble had the string value, stumble:

     [+ % fumble `printf '%%x\\n' $%s` +]

This would cause the shell to evaluate "printf '%x\n' $stumble". Assuming that the shell variable stumble had a numeric value, the expression result would be that number, in hex. Note the need for doubled percent characters and backslashes.

‘? <value-name> <basic-expr-1> <basic-expr-2>’

Two basic-expr-s are required. If the value-name is defined, then the first

basic-expr-1 is evaluated, otherwise basic-expr-2 is.
‘- <value-name> <basic-expr>’

Evaluate basic-expr only if value-name is not defined.

‘?% <value-name> <basic-expr-1> <basic-expr-2>’

This combines the functions of ‘?’ and ‘%’. If value-name is defined, it behaves exactly like ‘%’, above, using basic-expr-1. If not defined, then basic-expr-2 is evaluated. For example, assume again that fumble had the string value, stumble:

     [+ ?% fumble `cat $%s` `pwd` +]

This would cause the shell to evaluate "cat $stumble". If fumble were not defined, then the result would be the name of our current directory.

3.3.2 Basic Expression

A basic expression can have one of the following forms:

‘'STRING'’

A single quoted string. Backslashes can be used to protect single quotes ('), hash characters (#), or backslashes (\) in the string. All other characters of STRING are output as-is when the single quoted string is evaluated. Backslashes are processed before the hash character for consistency with the definition syntax. It is needed there to avoid preprocessing conflicts.

‘"STRING"’

A double quoted string. This is a cooked text string as in C, except that they are not concatenated with adjacent strings. Evaluating "STRING" will output STRING with all backslash sequences interpreted.

‘`STRING`’

A back quoted string. When this expression is evaluated, STRING is first interpreted as a cooked string (as in ‘"STRING"’) and evaluated as a shell expression by the AutoGen server shell. This expression is replaced by the stdout output of the shell.

‘(STRING)’

A parenthesized expression. It will be passed to the Guile interpreter for evaluation and replaced by the resulting value. If there is a Scheme error in this expression, Guile 1.4 and Guile 1.6 will report the template line number where the error occurs. Guile 1.7 has lost this capability. Additionally, other than in the % and ?% expressions, the Guile expressions may be introduced with the Guile comment character (;) and you may put a series of Guile expressions within a single macro. They will be implicitly evaluated as if they were arguments to the (begin ...) expression. The result will be the result of the last Guile expression evaluated.

3.4 AutoGen Scheme Functions

AutoGen uses Guile to interpret Scheme expressions within AutoGen macros. All of the normal Guile functions are available, plus several extensions (see Section 3.5 [Common Functions], page 31) have been added to augment the repertoire of string manipulation functions and manage the state of AutoGen processing.

This section describes those functions that are specific to AutoGen. Please take note that these AutoGen specific functions are not loaded and thus not made available until after the command line options have been processed and the AutoGen definitions have been loaded. They may, of course, be used in Scheme functions that get defined at those times, but they cannot be invoked.

3.4.1 ‘ag-function?’ - test for function

Usage: (ag-function? ag-name) return SCM BOOL T if a specified name is a user-defined AutoGen macro, otherwise return SCM BOOL F.

Arguments: ag-name - name of AutoGen macro

3.4.2 ‘base-name’ - base output name

Usage: (base-name) Returns a string containing the base name of the output file(s). Generally, this is also the base name of the definitions file.

This Scheme function takes no arguments.

3.4.3 ‘chdir’ - Change current directory

Usage: (chdir dir) Sets the current directory for AutoGen. Shell commands will run from this directory as well. This is a wrapper around the Guile native function. It returns its directory name argument and fails the program on failure.

Arguments: dir - new directory name

3.4.4 ‘count’ - definition count

Usage: (count ag-name) Count the number of entries for a definition. The input argument must be a string containing the name of the AutoGen values to be counted. If there is no value associated with the name, the result is an SCM immediate integer value of zero.

Arguments: ag-name - name of AutoGen value

3.4.5 ‘def-file’ - definitions file name

Usage: (def-file) Get the name of the definitions file. Returns the name of the source file containing the AutoGen definitions.

This Scheme function takes no arguments.

3.4.6 ‘def-file-line’ - get a definition file+line number

Usage: (def-file-line ag-name [ msg-fmt ]) Returns the file and line number of a AutoGen defined value, using either the default format, "from %s line %d", or else the format you supply. For example, if you want to insert a "C" language file-line directive, you would supply the format "# %2$d \"%1$s\"", but that is also already supplied with the scheme variable See Section 3.4.42 [SCM c-file-line-fmt], page 30. You may use it thus:

     (def-file-line "ag-def-name" c-file-line-fmt)

It is also safe to use the formatting string, "%2$d". AutoGen uses an argument vector version of printf: See Section 8.7 [snprintfv], page 157.

Arguments: ag-name - name of AutoGen value msg-fmt - Optional - formatting for line message

3.4.7 ‘dne’ - "Do Not Edit" warning

Usage: (dne prefix [ first prefix ] [ optpfx ]) Generate a "DO NOT EDIT" or "EDIT WITH CARE" warning string. Which depends on whether or not the --writable command line option was set. The first argument is a per-line string prefix. The optional second argument is a prefix for the first-line and, in read-only mode, activates the editor hints.

     -*- buffer-read-only: t -*- vi: set ro:

The warning string also includes information about the template used to construct the file and the definitions used in its instantiation.

The optional third argument is used when the first argument is actually an invocation option and the prefix arguments get shifted. The first argument must be, specifically, "-d". That is used to signify that the date stamp should not be inserted into the output.

Arguments: prefix - string for starting each output line first prefix - Optional - for the first output line optpfx - Optional - shifted prefix

3.4.8 ‘error’ - display message and exit

Usage: (error message) The argument is a string that printed out as part of an error message. The message is formed from the formatting string:

     DEFINITIONS ERROR in %s line %d for %s:  %s\n

The first three arguments to this format are provided by the routine and are: The name of the template file, the line within the template where the error was found, and the current output file name.

After displaying the message, the current output file is removed and autogen exits with the EXIT FAILURE error code. IF, however, the argument begins with the number 0 (zero), or the string is the empty string, then processing continues with the next suffix.

Arguments: message - message to display before exiting

3.4.9 ‘exist?’ - test for value name

Usage: (exist? ag-name) return SCM BOOL T iff a specified name has an AutoGen value. The name may include indexes and/or member names. All but the last member name must be an aggregate definition. For example:

     (exist? "foo[3].bar.baz")

will yield true if all of the following is true: There is a member value of either group or string type named baz for some group value bar that is a member of the foo group with index 3. There may be multiple entries of bar within foo, only one needs to contain a value for baz.

Arguments: ag-name - name of AutoGen value

3.4.10 ‘find-file’ - locate a file in the search path

Usage: (find-file file-name [ suffix ]) AutoGen has a search path that it uses to locate template and definition files. This function will search the same list for ‘file-name’, both with and without the ‘.suffix’, if provided.

Arguments: file-name - name of file with text suffix - Optional - file suffix to try, too

3.4.11 ‘first-for?’ - detect first iteration

Usage: (first-for? [ for var ]) Returns SCM BOOL T if the named FOR loop (or, if not named, the current innermost loop) is on the first pass through the data. Outside of any FOR loop, it returns SCM UNDEFINED. See Section 3.6.13 [FOR], page 49.

Arguments: for var - Optional - which for loop

3.4.12 ‘for-by’ - set iteration step

Usage: (for-by by) This function records the "step by" information for an AutoGen FOR function. Outside of the FOR macro itself, this function will emit an error. See Section 3.6.13 [FOR], page 49.

Arguments: by - the iteration increment for the AutoGen FOR macro

3.4.13 ‘for-from’ - set initial index

Usage: (for-from from) This function records the initial index information for an AutoGen FOR function. Outside of the FOR macro itself, this function will emit an error. See Section 3.6.13 [FOR], page 49.

Arguments: from - the initial index for the AutoGen FOR macro

3.4.14 ‘for-index’ - get current loop index

Usage: (for-index [ for var ]) Returns the current index for the named FOR loop. If not named, then the index for the innermost loop. Outside of any FOR loop, it returns SCM UNDEFINED. See Section 3.6.13 [FOR], page 49.

Arguments: for var - Optional - which for loop

3.4.15 ‘for-sep’ - set loop separation string

Usage: (for-sep separator) This function records the separation string that is to be inserted between each iteration of an AutoGen FOR function. This is often nothing more than a comma. Outside of the FOR macro itself, this function will emit an error.

Arguments: separator - the text to insert between the output of each FOR iteration

3.4.16 ‘for-to’ - set ending index

Usage: (for-to to) This function records the terminating value information for an AutoGen FOR function. Outside of the FOR macro itself, this function will emit an error. See Section 3.6.13 [FOR], page 49.

Arguments: to - the final index for the AutoGen FOR macro

3.4.17 ‘get’ - get named value

Usage: (get ag-name [ alt-val ]) Get the first string value associated with the name. It will either return the associated string value (if the name resolves), the alternate value (if one is provided), or else the empty string.

Arguments: ag-name - name of AutoGen value alt-val - Optional - value if not present

3.4.18 ‘high-lim’ - get highest value index

Usage: (high-lim ag-name) Returns the highest index associated with an array of definitions. This is generally, but not necessarily, one less than the count value. (The indexes may be specified, rendering a non-zero based or sparse array of values.)

This is very useful for specifying the size of a zero-based array of values where not all values are present. For example:

     tMyStruct myVals[ [+ (+ 1 (high-lim "my-val-list")) +] ];

Arguments: ag-name - name of AutoGen value

3.4.19 ‘last-for?’ - detect last iteration

Usage: (last-for? [ for var ]) Returns SCM BOOL T if the named FOR loop (or, if not named, the current innermost loop) is on the last pass through the data. Outside of any FOR loop, it returns SCM UNDEFINED. See Section 3.6.13 [FOR], page 49.

Arguments: for var - Optional - which for loop

3.4.20 ‘len’ - get count of values

Usage: (len ag-name) If the named object is a group definition, then "len" is the same as "count". Otherwise, if it is one or more text definitions, then it is the sum of their string lengths. If it is a single text definition, then it is equivalent to (string-length (get "ag-name")).

Arguments: ag-name - name of AutoGen value

3.4.21 ‘low-lim’ - get lowest value index

Usage: (low-lim ag-name) Returns the lowest index associated with an array of definitions.

Arguments: ag-name - name of AutoGen value

3.4.22 ‘make-header-guard’ - make self-inclusion guard

Usage: (make-header-guard name) This function will create a #ifndef/#define sequence for protecting a header from multiple evaluation. It will also set the Scheme variable header-file to the name of the file being protected and it will set header-guard to the name of the #define being used to protect it. It is expected that this will be used as follows:

     [+ (make-header-guard "group_name") +]
     ...
     #endif /* [+ (. header-guard) +]
     #include "[+ (. header-file)  +]"

The #define name is composed as follows:

1. The first element is the string argument and a separating underscore.

2. That is followed by the name of the header file with illegal characters mapped to underscores.

3. The end of the name is always, "_GUARD".

4. Finally, the entire string is mapped to upper case.

The final #define name is stored in an SCM symbol named header-guard. Consequently, the concluding #endif for the file should read something like:

     #endif /* [+ (. header-guard) +] */

The name of the header file (the current output file) is also stored in an SCM symbol, header-file. Therefore, if you are also generating a C file that uses the previously generated header file, you can put this into that generated file:

     #include "[+ (. header-file) +]"

Obviously, if you are going to produce more than one header file from a particular template, you will need to be careful how these SCM symbols get handled.

Arguments: name - header group name

3.4.23 ‘match-value?’ - test for matching value

Usage: (match-value? op ag-name test-str) This function answers the question, "Is there an AutoGen value named ag-name with a value that matches the pattern test-str using the match function op?" Return SCM BOOL T iff at least one occurrence of the specified name has such a value. The operator can be any function that takes two string arguments and yields a boolean. It is expected that you will use one of the string matching functions provided by AutoGen. The value name must follow the same rules as the ag-name argument for exist? (see Section 3.4.9 [SCM exist?], page 24).

Arguments: op - boolean result operator ag-name - name of AutoGen value test-str - string to test against

3.4.24 ‘out-delete’ - delete current output file

Usage: (out-delete) Remove the current output file. Cease processing the template for the current suffix. It is an error if there are push-ed output files. Use the (error "0") scheme function instead. See Section 3.7 [output controls], page 52.

This Scheme function takes no arguments.

3.4.25 ‘out-depth’ - output file stack depth

Usage: (out-depth) Returns the depth of the output file stack. See Section 3.7 [output controls], page 52.

This Scheme function takes no arguments.

3.4.26 ‘out-line’ - output file line number

Usage: (out-line) Returns the current line number of the output file. It rewinds and reads the file to count newlines.

This Scheme function takes no arguments.

3.4.27 ‘out-move’ - change name of output file

Usage: (out-move new-name) Rename current output file. See Section 3.7 [output controls], page 52. Please note: changing the name will not save a temporary file from being deleted. It may, however, be used on the root output file.

Arguments: new-name - new name for the current output file

3.4.28 ‘out-name’ - current output file name

Usage: (out-name) Returns the name of the current output file. If the current file is a temporary, unnamed file, then it will scan up the chain until a real output file name is found. See Section 3.7 [output controls], page 52.

This Scheme function takes no arguments.

3.4.29 ‘out-pop’ - close current output file

Usage: (out-pop [ disp ]) If there has been a push on the output, then close that file and go back to the previously open file. It is an error if there has not been a push. See Section 3.7 [output controls], page 52.

If there is no argument, no further action is taken. Otherwise, the argument should be #t and the contents of the file are returned by the function.

Arguments: disp - Optional - return contents of the file

3.4.30 ‘out-push-add’ - append output to file

Usage: (out-push-add file-name) Identical to push-new, except the contents are not purged, but appended to. See Section 3.7 [output controls], page 52.

Arguments: file-name - name of the file to append text to

3.4.31 ‘out-push-new’ - purge and create output file

Usage: (out-push-new [ file-name ]) Leave the current output file open, but purge and create a new file that will remain open until a popdelete or switch closes it. The file name is optional and, if omitted, the output will be sent to a temporary file that will be deleted when it is closed. See Section 3.7 [output controls], page 52.

Arguments: file-name - Optional - name of the file to create

3.4.32 ‘out-resume’ - resume suspended output file

Usage: (out-resume suspName) If there has been a suspended output, then make that output descriptor current again. That output must have been suspended with the same tag name given to this routine as its argument.

Arguments: suspName - A name tag for reactivating

3.4.33 ‘out-suspend’ - suspend current output file

Usage: (out-suspend suspName) If there has been a push on the output, then set aside the output descriptor for later reactiviation with (out-resume "xxx"). The tag name need not reflect the name of the output file. In fact, the output file may be an anonymous temporary file. You may also change the tag every time you suspend output to a file, because the tag names are forgotten as soon as the file has been "resumed".

Arguments: suspName - A name tag for reactivating

3.4.34 ‘out-switch’ - close and create new output

Usage: (out-switch file-name) Switch output files - close current file and make the current file pointer refer to the new file. This is equivalent to out-pop followed by out-push-new, except that you may not pop the base level output file, but you may switch it. See Section 3.7 [output controls], page 52.

Arguments: file-name - name of the file to create

3.4.35 ‘set-option’ - Set a command line option

Usage: (set-option opt) The text argument must be an option name followed by any needed option argument. Returns SCM UNDEFINED.

Arguments: opt - AutoGen option name + its argument

3.4.36 ‘set-writable’ - Make the output file be writable

Usage: (set-writable [ set? ]) This function will set the current output file to be writable (or not). This is only effective if neither the --writable nor --not-writable have been specified. This state is reset when the current suffix’s output is complete.

Arguments: set? - Optional - boolean arg, false to make output non-writable

3.4.37 ‘stack’ - make list of AutoGen values

Usage: (stack ag-name) Create a scheme list of all the strings that are associated with a name. They must all be text values or we choke.

Arguments: ag-name - AutoGen value name

3.4.38 ‘suffix’ - get the current suffix

Usage: (suffix) Returns the current active suffix (see Section 3.1 [pseudo macro], page 17).

This Scheme function takes no arguments.

3.4.39 ‘tpl-file’ - get the template file name

Usage: (tpl-file [ full path ]) Returns the name of the current template file. If #t is passed in as an argument, then the template file is hunted for in the template search path. Otherwise, just the unadorned name.

Arguments: full path - Optional - include full path to file

3.4.40 ‘tpl-file-line’ - get the template file+line number

Usage: (tpl-file-line [ msg-fmt ]) Returns the file and line number of the current template macro using either the default format, "from %s line %d", or else the format you supply. For example, if you want to insert a "C" language file-line directive, you would supply the format "# %2$d \"%1$s\"", but that is also already supplied with the scheme variable See Section 3.4.42 [SCM c-file-line-fmt], page 30. You may use it thus:

     (tpl-file-line c-file-line-fmt)

It is also safe to use the formatting string, "%2$d". AutoGen uses an argument vector version of printf: See Section 8.7 [snprintfv], page 157.

Arguments: msg-fmt - Optional - formatting for line message

3.4.41 ‘autogen-version’ - autogen version number

This is a symbol defining the current AutoGen version number string. It was first defined in AutoGen-5.2.14. It is currently “5.8.6”.

3.4.42 format file info as, “#line nn "file"”

This is a symbol that can easily be used with the functions See Section 3.4.40 [SCM tpl-file-line], page 30, and See Section 3.4.6 [SCM def-file-line], page 23. These will emit C program #line directives pointing to template and definitions text, respectively.

3.5 Common Scheme Functions

This section describes a number of general purpose functions that make the kind of string processing that AutoGen does a little easier. Unlike the AutoGen specific functions (see Section 3.4 [AutoGen Functions], page 22), these functions are available for direct use during definition load time. The equality test (see Section 3.5.42 [SCM =], page 41) is “overloaded” to do string equivalence comparisons. If you are looking for inequality, the Scheme/Lisp way of spelling that is, “(not (= ...))”.

3.5.1 ‘ag-fprintf’ - format to autogen stream

Usage: (ag-fprintf ag-diversion format [ format-arg ... ]) Format a string using arguments from the alist. Write to a specified AutoGen diversion. That may be either a specified suspended output stream (see Section 3.4.33 [SCM outsuspend], page 29) or an index into the output stack (see Section 3.4.31 [SCM out-push-new],

page 28). (ag-fprintf 0 ...) is equivalent to (emit (sprintf ...)), and (ag-fprintf

1 ...) sends output to the most recently suspended output stream.

Arguments: ag-diversion - AutoGen diversion name or number format - formatting string format-arg - Optional - list of arguments to formatting string

3.5.2 ‘bsd’ - BSD Public License

Usage: (bsd prog name owner prefix) Emit a string that contains the Free BSD Public License. It takes three arguments: prefix contains the string to start each output line. owner contains the copyright owner. prog_ name contains the name of the program the copyright is about.

Arguments: prog name - name of the program under the BSD owner - Grantor of the BSD License prefix - String for starting each output line

3.5.3 ‘c-string’ - emit string for ANSI C

Usage: (c-string string) Reform a string so that, when printed, the C compiler will be able to compile the data and construct a string that contains exactly what the current string contains. Many non-printing characters are replaced with escape sequences. Newlines are replaced with a backslash, an n, a closing quote, a newline, seven spaces and another re-opening quote. The compiler will implicitly concatenate them. The reader will see line breaks.

A K&R compiler will choke. Use kr-string for that compiler.

Arguments: string - string to reformat

3.5.4 ‘emit’ - emit the text for each argument

Usage: (emit alist ...) Walk the tree of arguments, displaying the values of displayable SCM types.

Arguments: alist - list of arguments to stringify and emit

3.5.5 ‘emit-string-table’ - output a string table

Usage: (emit-string-table st-name) Emit into the current output stream a static char const array named st-name that will have NUL bytes between each inserted string.

Arguments: st-name - the name of the array of characters

3.5.6 ‘error-source-line’ - display of file & line

Usage: (error-source-line) This function is only invoked just before Guile displays an error message. It displays the file name and line number that triggered the evaluation error. You should not need to invoke this routine directly. Guile will do it automatically.

This Scheme function takes no arguments.

3.5.7 ‘extract’ - extract text from another file

Usage: (extract file-name marker-fmt [ caveat ] [ default ]) This function is used to help construct output files that may contain text that is carried from one version of the output to the next.

The first two arguments are required, the second are optional:

• The file-name argument is used to name the file that contains the demarcated text.

• The marker-fmt is a formatting string that is used to construct the starting and ending demarcation strings. The sprintf function is given the marker-fmt with two arguments. The first is either "START" or "END". The second is either "DO NOT CHANGE THIS COMMENT" or the optional caveat argument.

• caveat is presumed to be absent if it is the empty string (""). If absent, “DO NOT CHANGE THIS COMMENT” is used as the second string argument to the marker-

    fmt.

• When a default argument is supplied and no pre-existing text is found, then this text will be inserted between the START and END markers.

The resulting strings are presumed to be unique within the subject file. As a simplified example:

     [+ (extract "fname" "// %s - SOMETHING - %s" ""
     "example default") +]

will result in the following text being inserted into the output:

     // START - SOMETHING - DO NOT CHANGE THIS COMMENT
     example default
     // END   - SOMETHING - DO NOT CHANGE THIS COMMENT

The “example default” string can then be carried forward to the next generation of the output, provided the output is not named "fname" and the old output is renamed to "fname" before AutoGen-eration begins.

NB: You can set aside previously generated source files inside the pseudo macro with a Guile/scheme function, extract the text you want to keep with this extract function. Just remember you should delete it at the end, too. Here is an example from my Finite State Machine generator:

     [+ AutoGen5 Template  -*- Mode: text -*-
     h=%s-fsm.h   c=%s-fsm.c
     (shellf
     "[ -f %1$s-fsm.h ] && mv -f %1$s-fsm.h .fsm.head
     [ -f %1$s-fsm.c ] && mv -f %1$s-fsm.c .fsm.code" (base-name)) +]

This code will move the two previously produced output files to files named ".fsm.head" and ".fsm.code". At the end of the ’c’ output processing, I delete them.

Arguments: file-name - name of file with text marker-fmt - format for marker text caveat - Optional - warn about changing marker default - Optional - default initial text

3.5.8 ‘format-arg-count’ - count the args to a format

Usage: (format-arg-count format) Sometimes, it is useful to simply be able to figure out how many arguments are required by a format string. For example, if you are extracting a format string for the purpose of generating a macro to invoke a printf-like function, you can run the formatting string through this function to determine how many arguments to provide for in the macro. e.g. for this extraction text:

      /*=fumble bumble
       * fmt: 'stumble %s: %d\n'
      =*/

You may wish to generate a macro:

      #define BUMBLE(a1,a2) printf_like(something,(a1),(a2))

You can do this by knowing that the format needs two arguments.

Arguments: format - formatting string

3.5.9 ‘fprintf’ - format to a file

Usage: (fprintf port format [ format-arg ... ]) Format a string using arguments from the alist. Write to a specified port. The result will NOT appear in your output. Use this to print information messages to a template user.

Arguments: port - Guile-scheme output port format - formatting string format-arg - Optional - list of arguments to formatting string

3.5.10 ‘gperf’ - perform a perfect hash function

Usage: (gperf name str) Perform the perfect hash on the input string. This is only useful if you have previously created a gperf program with the make-gperf function See Section 3.5.19 [SCM make-gperf], page 35. The name you supply here must match the name used to create the program and the string to hash must be one of the strings supplied in the make-gperf string list. The result will be a perfect hash index.

See the documentation for gperf(1GNU) for more details.

Arguments: name - name of hash list str - string to hash

3.5.11 ‘gpl’ - GNU General Public License

Usage: (gpl prog-name prefix) Emit a string that contains the GNU General Public License. It takes two arguments: prefix contains the string to start each output line, and prog_name contains the name of the program the copyright is about.

Arguments: prog-name - name of the program under the GPL prefix - String for starting each output line

3.5.12 ‘hide-email’ - convert eaddr to javascript

Usage: (hide-email display eaddr) Hides an email address as a java scriptlett. The ’mailto:’ tag and the email address are coded bytes rather than plain text. They are also broken up.

Arguments: display - display text eaddr - email address

3.5.13 ‘html-escape-encode’ - encode html special characters

Usage: (html-escape-encode str) This function will replace replace the characters '&', '<' and '>' characters with the HTML/XML escape-encoded strings ("&amp;", "&lt;", and "&gt;", respectively).

Arguments: str - string to make substitutions in

3.5.14 ‘in?’ - test for string in list

Usage: (in? test-string string-list ...) Return SCM BOOL T if the first argument string is found in one of the entries in the second (list-of-strings) argument.

Arguments: test-string - string to look for string-list - list of strings to check

3.5.15 ‘join’ - join string list with separator

Usage: (join separator list ...) With the first argument as the separator string, joins together an a-list of strings into one long string. The list may contain nested lists, partly because you cannot always control that.

Arguments: separator - string to insert between entries list - list of strings to join

3.5.16 ‘kr-string’ - emit string for K&R C

Usage: (kr-string string) Reform a string so that, when printed, a K&R C compiler will be able to compile the data and construct a string that contains exactly what the current string contains. Many non-printing characters are replaced with escape sequences. New-lines are replaced with a backslash-n-backslash and newline sequence,

Arguments: string - string to reformat

3.5.17 ‘lgpl’ - GNU Library General Public License

Usage: (lgpl prog name owner prefix) Emit a string that contains the GNU Library General Public License. It takes three arguments: prefix contains the string to start each output line. owner contains the copyright owner. prog_name contains the name of the program the copyright is about.

Arguments: prog name - name of the program under the LGPL owner - Grantor of the LGPL prefix - String for starting each output line

3.5.18 ‘license’ - an arbitrary license

Usage: (license lic name prog name owner prefix) Emit a string that contains the named license. The license text is read from a file named, lic_name.lic, searching the standard directories. The file contents are used as a format argument to printf(3), with prog_name and owner as the two string formatting arguments. Each output line is automatically prefixed with the string prefix.

Arguments: lic name - file name of the license prog name - name of the licensed program or library owner - Grantor of the License prefix - String for starting each output line

3.5.19 ‘make-gperf’ - build a perfect hash function program

Usage: (make-gperf name strings ...) Build a program to perform perfect hashes of a known list of input strings. This function produces no output, but prepares a program named, ‘gperf_<name>’ for use by the gperf function See Section 3.5.10 [SCM gperf], page 34.

This program will be obliterated as AutoGen exits. However, you may incorporate the generated hashing function into your C program with commands something like the following:

     [+ (shellf "sed '/^int main(/,$d;/^#line/d' ${gpdir}/%s.c"
     name ) +]

where name matches the name provided to this make-perf function. gpdir is the variable used to store the name of the temporary directory used to stash all the files.

Arguments: name - name of hash list strings - list of strings to hash

3.5.20 ‘makefile-script’ - create makefile script

Usage: (makefile-script text) This function will take ordinary shell script text and reformat it so that it will work properly inside of a makefile shell script. Not every shell construct can be supported; the intent is to have most ordinary scripts work without much, if any, alteration.

The following transformations are performed on the source text:

1. Trailing whitespace on each line is stripped.

2. Except for the last line, the string, " ; \\" is appended to the end of every line that does not end with a backslash, semi-colon, conjunction operator or pipe. Note that this will mutilate multi-line quoted strings, but make renders it impossible to use multi-line constructs anyway.

3. If the line ends with a backslash, it is left alone.

4. If the line ends with one of the excepted operators, then a space and backslash is added.

5. The dollar sign character is doubled, unless it immediately precedes an opening parenthesis or the single character make macros ’*’, ’<’, ’@’, ’?’ or ’%’. Other single character make macros that do not have enclosing parentheses will fail. For shell usage of the "$@", "$?" and "$*" macros, you must enclose them with curly braces, e.g., "${?}". The ksh construct $(<command>) will not work. Though some makes accept ${var} constructs, this function will assume it is for shell interpretation and double the dollar character. You must use $(var) for all make substitutions.

6. Double dollar signs are replaced by four before the next character is examined.

7. Every line is prefixed with a tab, unless the first line already starts with a tab.

8. The newline character on the last line, if present, is suppressed.

9. Blank lines are stripped.

This function is intended to be used approximately as follows:

     $(TARGET) : $(DEPENDENCIES)
     <+ (out-push-new) +>
     ....mostly arbitrary shell script text....
     <+ (makefile-script (out-pop #t)) +>

Arguments: text - the text of the script

3.5.21 ‘max’ - maximum value in list

Usage: (max list ...) Return the maximum value in the list

Arguments: list - list of values. Strings are converted to numbers

3.5.22 ‘min’ - minimum value in list

Usage: (min list ...) Return the minimum value in the list

Arguments: list - list of values. Strings are converted to numbers

3.5.23 ‘prefix’ - prefix lines with a string

Usage: (prefix prefix text) Prefix every line in the second string with the first string.

For example, if the first string is "# " and the second contains:

     two
     lines

The result string will contain:

     # two
     # lines

Arguments: prefix - string to insert at start of each line text - multi-line block of text

3.5.24 ‘printf’ - format to stdout

Usage: (printf format [ format-arg ... ]) Format a string using arguments from the alist. Write to the standard out port. The result will NOT appear in your output. Use this to print information messages to a template user. Use “(sprintf ...)” to add text to your document.

Arguments: format - formatting string format-arg - Optional - list of arguments to formatting string

3.5.25 ‘raw-shell-str’ - single quote shell string

Usage: (raw-shell-str string) Convert the text of the string into a singly quoted string that a normal shell will process into the original string. (It will not do macro expansion later, either.) Contained single quotes become tripled, with the middle quote escaped with a backslash. Normal shells will reconstitute the original string.

Notice: some shells will not correctly handle unusual non-printing characters. This routine works for most reasonably conventional ASCII strings.

Arguments: string - string to transform

3.5.26 ‘shell’ - invoke a shell script

Usage: (shell command) Generate a string by writing the value to a server shell and reading the output back in. The template programmer is responsible for ensuring that it completes within 10 seconds. If it does not, the server will be killed, the output tossed and a new server started.

Arguments: command - shell command - the result value is stdout

3.5.27 ‘shell-str’ - double quote shell string

Usage: (shell-str string) Convert the text of the string into a double quoted string that a normal shell will process into the original string, almost. It will add the escape character \\ before two special characters to accomplish this: the backslash \\ and double quote ".

NOTE: some shells will not correctly handle unusual non-printing characters. This routine works for most reasonably conventional ASCII strings.

WARNING: This function omits the extra backslash in front of a backslash, however, if it is followed by either a backquote or a dollar sign. It must do this because otherwise it would be impossible to protect the dollar sign or backquote from shell evaluation. Consequently, it is not possible to render the strings "\\$" or "\\‘". The lesser of two evils.

All others characters are copied directly into the output.

The sub-shell-str variation of this routine behaves identically, except that the extra backslash is omitted in front of " instead of `. You have to think about it. I’m open to suggestions.

Meanwhile, the best way to document is with a detailed output example. If the backslashes make it through the text processing correctly, below you will see what happens with three example strings. The first example string contains a list of quoted foos, the second is the same with a single backslash before the quote characters and the last is with two backslash escapes. Below each is the result of the raw-shell-str, shell-str and

sub-shell-str functions.
     foo[0]           ''foo'' 'foo' "foo" `foo` $foo
     raw-shell-str -> \'\''foo'\'\'' '\''foo'\'' "foo" `foo` $foo'
     shell-str     -> "''foo'' 'foo' \"foo\" `foo` $foo"
     sub-shell-str -> `''foo'' 'foo' "foo" \`foo\` $foo`
     foo[1]           \'bar\' \"bar\" \`bar\` \$bar
     raw-shell-str -> '\'\''bar\'\'' \"bar\" \`bar\` \$bar'
     shell-str     -> "\\'bar\\' \\\"bar\\\" \`bar\` \$bar"
     sub-shell-str -> `\\'bar\\' \"bar\" \\\`bar\\\` \$bar`
     foo[2]           \\'BAZ\\' \\"BAZ\\" \\`BAZ\\` \\$BAZ
     raw-shell-str -> '\\'\''BAZ\\'\'' \\"BAZ\\" \\`BAZ\\` \\$BAZ'
     shell-str     -> "\\\\'BAZ\\\\' \\\\\"BAZ\\\\\" \\\`BAZ\\\` \\\$BAZ"
     sub-shell-str -> `\\\\'BAZ\\\\' \\\"BAZ\\\" \\\\\`BAZ\\\\\` \\\$BAZ`

There should be four, three, five and three backslashes for the four examples on the last line, respectively. The next to last line should have four, five, three and three backslashes. If this was not accurately reproduced, take a look at the agen5/test/shell.test test. Notice the backslashes in front of the dollar signs. It goes from zero to one to three for the "cooked" string examples.

Arguments: string - string to transform

3.5.28 ‘shellf’ - format a string, run shell

Usage: (shellf format [ format-arg ... ]) Format a string using arguments from the alist, then send the result to the shell for interpretation.

Arguments: format - formatting string format-arg - Optional - list of arguments to formatting string

3.5.29 ‘sprintf’ - format a string

Usage: (sprintf format [ format-arg ... ]) Format a string using arguments from the alist.

Arguments: format - formatting string format-arg - Optional - list of arguments to formatting string

3.5.30 ‘string-capitalize’ - capitalize a new string

Usage: (string-capitalize str) Create a new SCM string containing the same text as the original, only all the first letter of each word is upper cased and all other letters are made lower case.

Arguments: str - input string

3.5.31 ‘string-capitalize!’ - capitalize a string

Usage: (string-capitalize! str) capitalize all the words in an SCM string.

Arguments: str - input/output string

3.5.32 ‘string-contains-eqv?’ - caseless substring

Usage: (*=* text match) string-contains-eqv?: Test to see if a string contains an equivalent string. ‘equivalent’ means the strings match, but without regard to character case and certain characters are considered ‘equivalent’. Viz., ’-’, ’ ’ and ’^’ are equivalent.

Arguments: text - text to test for pattern match - pattern/substring to search for

3.5.33 ‘string-contains?’ - substring match

Usage: (*==* text match) string-contains?: Test to see if a string contains a substring. "strstr(3)" will find an address.

Arguments: text - text to test for pattern match - pattern/substring to search for

3.5.34 ‘string-downcase’ - lower case a new string

Usage: (string-downcase str) Create a new SCM string containing the same text as the original, only all the upper case letters are changed to lower case.

Arguments: str - input string

3.5.35 ‘string-downcase!’ - make a string be lower case

Usage: (string-downcase! str) Change to lower case all the characters in an SCM string.

Arguments: str - input/output string

3.5.36 ‘string-end-eqv-match?’ - caseless regex ending

Usage: (*~ text match) string-end-eqv-match?: Test to see if a string ends with a pattern. Case is not significant.

Arguments: text - text to test for pattern match - pattern/substring to search for

3.5.37 ‘string-end-match?’ - regex match end

Usage: (*~~ text match) string-end-match?: Test to see if a string ends with a pattern. Case is significant.

Arguments: text - text to test for pattern match - pattern/substring to search for

3.5.38 ‘string-ends-eqv?’ - caseless string ending

Usage: (*= text match) string-ends-eqv?: Test to see if a string ends with an equivalent string.

Arguments: text - text to test for pattern match - pattern/substring to search for

3.5.39 ‘string-ends-with?’ - string ending

Usage: (*== text match) string-ends-with?: Test to see if a string ends with a substring. strcmp(3) returns zero for comparing the string ends.

Arguments: text - text to test for pattern match - pattern/substring to search for

3.5.40 ‘string-equals?’ - string matching

Usage: (== text match) string-equals?: Test to see if two strings exactly match.

Arguments: text - text to test for pattern match - pattern/substring to search for

3.5.41 ‘string-eqv-match?’ - caseless regex match

Usage: (~ text match) string-eqv-match?: Test to see if a string fully matches a pattern. Case is not significant, but any character equivalences must be expressed in your regular expression.

Arguments: text - text to test for pattern match - pattern/substring to search for

3.5.42 ‘string-eqv?’ - caseless string match

Usage: (= text match) string-eqv?: Test to see if two strings are equivalent. ‘equivalent’ means the strings match, but without regard to character case and certain characters are considered ‘equivalent’. Viz., ’-’, ’ ’ and ’^’ are equivalent. If the arguments are not strings, then the result of the numeric comparison is returned.

This is an overloaded operation. If the arguments are not both strings, then the query is passed through to scm_num_eq_p().

Arguments: text - text to test for pattern match - pattern/substring to search for

3.5.43 ‘string-has-eqv-match?’ - caseless regex contains

Usage: (*~* text match) string-has-eqv-match?: Test to see if a string contains a pattern. Case is not significant.

Arguments: text - text to test for pattern match - pattern/substring to search for

3.5.44 ‘string-has-match?’ - contained regex match

Usage: (*~~* text match) string-has-match?: Test to see if a string contains a pattern. Case is significant.

Arguments: text - text to test for pattern match - pattern/substring to search for

3.5.45 ‘string-match?’ - regex match

Usage: (~~ text match) string-match?: Test to see if a string fully matches a pattern. Case is significant.

Arguments: text - text to test for pattern match - pattern/substring to search for

3.5.46 ‘string-start-eqv-match?’ - caseless regex start

Usage: (~* text match) string-start-eqv-match?: Test to see if a string starts with a pattern. Case is not significant.

Arguments: text - text to test for pattern match - pattern/substring to search for

3.5.47 ‘string-start-match?’ - regex match start

Usage: (~~* text match) string-start-match?: Test to see if a string starts with a pattern. Case is significant.

Arguments: text - text to test for pattern match - pattern/substring to search for

3.5.48 ‘string-starts-eqv?’ - caseless string start

Usage: (=* text match) string-starts-eqv?: Test to see if a string starts with an equivalent string.

Arguments: text - text to test for pattern match - pattern/substring to search for

3.5.49 ‘string-starts-with?’ - string starting

Usage: (==* text match) string-starts-with?: Test to see if a string starts with a substring.

Arguments: text - text to test for pattern match - pattern/substring to search for

3.5.50 ‘string-substitute’ - multiple global replacements

Usage: (string-substitute source match repl) match and repl may be either a single string or a list of strings. Either way, they must have the same structure and number of elements. For example, to replace all less than and all greater than characters, do something like this:

     (string-substitute source
     ("&"     "<"    ">")
     ("&amp;" "&lt;" "&gt;"))

Arguments: source - string to transform match - substring or substring list to be replaced repl - replacement strings or substrings

3.5.51 ‘string-table-add’ - Add an entry to a string table

Usage: (string-table-add st-name str-val) Check for a duplicate string and, if none, then insert a new string into the string table. In all cases, returns the character index of the beginning of the string in the table.

The returned index can be used in expressions like:

     string_array + <returned-value>

that will yield the address of the first byte of the inserted string. See the ‘strtable.test’ AutoGen test for a usage example.

Arguments: st-name - the name of the array of characters str-val - the (possibly) new value to add

3.5.52 ‘string-table-new’ - create a string table

Usage: (string-table-new st-name) This function will create an array of characters. The companion functions, (See Section 3.5.51 [SCM string-table-add], page 43, and see Section 3.5.5 [SCM emit-string-table], page 32) will insert text and emit the populated table, respectively.

With these functions, it should be much easier to construct structures containing string offsets instead of string pointers. That can be very useful when transmitting, storing or sharing data with different address spaces. Here is a brief example copied from the strtable.test test:

     [+ (string-table-new "scribble")
        (out-push-new)
        (define ix 0)
        (define ct 1)  +][+
     FOR str IN that was the week that was +][+
       (set! ct (+ ct 1))
       (set! ix (string-table-add "scribble" (get "str")))
     +]
         scribble + [+ (. ix) +],[+
     ENDFOR  +]
         NULL };
     [+ (out-suspend "main")
        (emit-string-table "scribble")
        (ag-fprintf 0 "\nchar const *ap[%d] = {" ct)
        (out-resume "main")
        (out-pop #t) +]

Some explanation: I added the (out-push-new) because the string table text is diverted into an output stream named, “scribble” and I want to have the string table emitted before the string table references. The string table references are also emitted inside the FOR loop. So, when the loop is done, the current output is suspended under the name, “main” and the “scribble” table is then emitted into the primary output. (emit-string-table inserts its output directly into the current output stream. It does not need to be the last function in an AutoGen macro block.) Next I ag-fprintf the array-of-pointer declaration directly into the current output. Finally I restore the “main” output stream and (out-pop #t)-it into the main output stream.

Here is the result. Note that duplicate strings are not repeated in the string table:

     static char const scribble[18] =
         "that\0" "was\0"  "the\0"  "week\0";
     char const *ap[7] = {
         scribble + 0,
         scribble + 5,
         scribble + 9,
         scribble + 13,
         scribble + 0,
         scribble + 5,
         NULL };

These functions use the global name space stt-* in addition to the function names.

Arguments: st-name - the name of the array of characters

3.5.53 ‘string->c-name!’ - map non-name chars to underscore

Usage: (string->c-name! str) Change all the graphic characters that are invalid in a C name token into underscores. Whitespace characters are ignored. Any other character type (i.e. non-graphic and nonwhite) will cause a failure.

Arguments: str - input/output string

3.5.54 ‘string-tr’ - convert characters with new result

Usage: (string-tr source match translation) This is identical to string-tr!, except that it does not over-write the previous value.

Arguments: source - string to transform match - characters to be converted translation - conversion list

3.5.55 ‘string-tr!’ - convert characters

Usage: (string-tr! source match translation) This is the same as the tr(1) program, except the string to transform is the first argument. The second and third arguments are used to construct mapping arrays for the transformation of the first argument.

It is too bad this little program has so many different and incompatible implementations!

Arguments: source - string to transform match - characters to be converted translation - conversion list

3.5.56 ‘string-upcase’ - upper case a new string

Usage: (string-upcase str) Create a new SCM string containing the same text as the original, only all the lower case letters are changed to upper case.

Arguments: str - input string

3.5.57 ‘string-upcase!’ - make a string be upper case

Usage: (string-upcase! str) Change to upper case all the characters in an SCM string.

Arguments: str - input/output string

3.5.58 ‘sub-shell-str’ - back quoted (sub-)shell string

Usage: (sub-shell-str string) This function is substantially identical to shell-str, except that the quoting character is ` and the "leave the escape alone" character is ".

Arguments: string - string to transform

3.5.59 ‘sum’ - sum of values in list

Usage: (sum list ...) Compute the sum of the list of expressions.

Arguments: list - list of values. Strings are converted to numbers

3.5.60 ‘version-compare’ - compare two version numbers

Usage: (version-compare op v1 v2) Converts v1 and v2 strings into 64 bit values and returns the result of running ’op’ on those values. It assumes that the version is a 1 to 4 part dot-separated series of numbers. Suffixes like, "5pre4" or "5-pre4" will be interpreted as two numbers. The first number ("5" in this case) will be decremented and the number after the "pre" will be added to 0xC000. (Unless your platform is unable to support 64 bit integer arithmetic. Then it will be added to 0xC0.) Consequently, these yield true:

     (version-compare > "5.8.5"       "5.8.5-pre4")
     (version-compare > "5.8.5-pre10" "5.8.5-pre4")

Arguments: op - comparison operator v1 - first version v2 - compared-to version

3.6 AutoGen Native Macros

This section describes the various AutoGen natively defined macros. Unlike the Scheme functions, some of these macros are "block macros" with a scope that extends through a terminating macro. Block macros must not overlap. That is to say, a block macro started within the scope of an encompassing block macro must have its matching end macro appear before the encompassing block macro is either ended or subdivided.

The block macros are these:

CASE This macro has scope through the ESAC macro. The scope is subdivided by SELECT macros. You must have at least one SELECT macro.

DEFINE This macro has scope through the ENDDEF macro. The defined user macro can never be a block macro. This macro is extracted from the template before the template is processed.

FOR This macro has scope through the ENDFOR macro.

IF This macro has scope through the ENDIF macro. The scope may be subdivided by ELIF and ELSE macros. Obviously, there may be only one ELSE macro and it must be the last of these subdivisions.

INCLUDE This macro has the scope of the included file. It is a block macro in the sense that the included file must not contain any incomplete block macros.

WHILE This macro has scope through the ENDWHILE macro.

3.6.1 AutoGen Macro Syntax

The general syntax is:

     [ { <native-macro-name> | <user-defined-name> } ] [ <arg> ... ]

The syntax for <arg> depends on the particular macro, but is generally a full expression (see Section 3.3 [expression syntax], page 19). Here are the exceptions to that general rule:

1. INVOKE macros, implicit or explicit, must be followed by a list of name/string value pairs. The string values are simple expressions, as described above. That is, the INVOKE syntax is one of these two:

         <user-macro-name> [ <name> [ = <expression> ] ... ]
         INVOKE <name-expression> [ <name> [ = <expression> ] ... ]

2. AutoGen FOR macros must be in one of three forms:

         FOR <name> [ <separator-string> ]
         FOR <name> (...Scheme expression list)
         FOR <name> IN <string-entry> [ ... ]

where:

‘<name>’ must be a simple name.

    ‘<separator-string>’

is inserted between copies of the enclosed block. Do not try to use “IN” as your separator string. It won’t work.

‘<string-entry>’

is an entry in a list of strings. “<name>” is assigned each value from the “IN” list before expanding the FOR block.

‘(...Scheme expression list)’

is expected to contain one or more of the for-from, for-to, for-by, and for-sep functions. (See Section 3.6.13 [FOR], page 49, and Section 3.4 [AutoGen Functions], page 22)

The first two forms iterate over the FOR block if <name> is found in the AutoGen values. The last form will create the AutoGen value named <name>.

3. AutoGen DEFINE macros must be followed by a simple name. Anything after that is ignored. Consequently, that “comment space” may be used to document any named values the macro expects to have set up as arguments. See Section 3.6.4 [DEFINE], page 48.

4. The AutoGen COMMENT, ELSE, ESAC and the END* macros take no arguments and ignore everything after the macro name (e.g. see Section 3.6.3 [COMMENT], page 48)

3.6.2 CASE - Select one of several template blocks

The arguments are evaluated and converted to a string, if necessary. A simple name will be interpreted as an AutoGen value name and its value will be used by the SELECT macros (see the example below and the expression evaluation function, see Section 3.6.12 [EXPR], page 49). The scope of the macro is up to the matching ESAC macro. Within the scope of a CASE, this string is matched against case selection macros. There are sixteen match macros that are derived from four different ways matches may be performed, plus an "always true", "true if the AutoGen value was found", and "true if no AutoGen value was found" matches. The codes for the nineteen match macros are formed as follows:

1. Must the match start matching from the beginning of the string? If not, then the match macro code starts with an asterisk (*).

2. Must the match finish matching at the end of the string? If not, then the match macro code ends with an asterisk (*).

3. Is the match a pattern match or a string comparison? If a comparison, use an equal sign (=). If a pattern match, use a tilde (~).

4. Is the match case sensitive? If alphabetic case is important, double the tilde or equal sign.

5. Do you need a default match when none of the others match? Use a single asterisk (*).

6. Do you need to distinguish between an empty string value and a value that was not found? Use the non-existence test (!E) before testing a full match against an empty string (== ''). There is also an existence test (+E), more for symmetry than for practical use.

For example:

 [+ CASE <full-expression> +]
 [+ ~~*  "[Tt]est" +]reg exp must match at start, not at end
 [+ ==   "TeSt"    +]a full-string, case sensitive compare
 [+ =    "TEST"    +]a full-string, case insensitive compare

     [+ !E             +]not exists - matches if no AutoGen value found
     [+ ==   ""        +]expression yielded a zero-length string
     [+ +E             +]exists - matches if there is any value result
     [+ *              +]always match - no testing
     [+ ESAC +]

<full-expression> (see Section 3.3 [expression syntax], page 19) may be any expression, including the use of apply-codes and value-names. If the expression yields a number, it is converted to a decimal string.

These case selection codes have also been implemented as Scheme expression functions using the same codes. They are documented in this texi doc as “string-*?” predicates (see Section 3.5 [Common Functions], page 31).

3.6.3 COMMENT - A block of comment to be ignored

This function can be specified by the user, but there will never be a situation where it will be invoked at emit time. The macro is actually removed from the internal representation.

If the native macro name code is #, then the entire macro function is treated as a comment and ignored.

3.6.4 DEFINE - Define a user AutoGen macro

This function will define a new macro. You must provide a name for the macro. You do not specify any arguments, though the invocation may specify a set of name/value pairs that are to be active during the processing of the macro.

     [+ define foo +]
     ... macro body with macro functions ...
     [+ enddef +]
     ... [+ foo bar='raw text' baz=<<text expression>> +]

Once the macro has been defined, this new macro can be invoked by specifying the macro name as the first token after the start macro marker. Alternatively, you may make the invocation explicitly invoke a defined macro by specifying INVOKE (see Section 3.6.16 [INVOKE], page 51) in the macro invocation. If you do that, the macro name can be computed with an expression that gets evaluated every time the INVOKE macro is encountered.

Any remaining text in the macro invocation will be used to create new name/value pairs that only persist for the duration of the processing of the macro. The expressions are evaluated the same way basic expressions are evaluated. See Section 3.3 [expression syntax], page 19.

The resulting definitions are handled much like regular definitions, except:

1. The values may not be compound. That is, they may not contain nested name/value pairs.

2. The bindings go away when the macro is complete.

3. The name/value pairs are separated by whitespace instead of semi-colons.

4. Sequences of strings are not concatenated.

NB: The macro is extracted from the template as the template is scanned. You cannot conditionally define a macro by enclosing it in an IF/ENDIF (see Section 3.6.14 [IF], page 50) macro pair. If you need to dynamically select the format of a DEFINEd macro, then put the flavors into separate template files that simply define macros. INCLUDE (see Section 3.6.15 [INCLUDE], page 51) the appropriate template when you have computed which you need.

Due to this, it is acceptable and even a good idea to place all the DEFINE macros at the end of the template. That puts the main body of the template at the beginning of the file.

3.6.5 ELIF - Alternate Conditional Template Block

This macro must only appear after an IF function, and before any associated ELSE or ENDIF functions. It denotes the start of an alternate template block for the IF function. Its expression argument is evaluated as are the arguments to IF. For a complete description See Section 3.6.14 [IF], page 50.

3.6.6 ELSE - Alternate Template Block

This macro must only appear after an IF function, and before the associated ENDIF function. It denotes the start of an alternate template block for the IF function. For a complete description See Section 3.6.14 [IF], page 50.

3.6.7 ENDDEF - Ends a macro definition.

This macro ends the DEFINE function template block. For a complete description See Section 3.6.4 [DEFINE], page 48.

3.6.8 ENDFOR - Terminates the FOR function template block

This macro ends the FOR function template block. For a complete description See Section 3.6.13 [FOR], page 49.

3.6.9 ENDIF - Terminate the IF Template Block

This macro ends the IF function template block. For a complete description See Section 3.6.14 [IF], page 50.

3.6.10 ENDWHILE - Terminate the WHILE Template Block

This macro ends the WHILE function template block. For a complete description See Section 3.6.19 [WHILE], page 52.

3.6.11 ESAC - Terminate the CASE Template Block

This macro ends the CASE function template block. For a complete description, See Section 3.6.2 [CASE], page 47.

3.6.12 EXPR - Evaluate and emit an Expression

This macro does not have a name to cause it to be invoked explicitly, though if a macro starts with one of the apply codes or one of the simple expression markers, then an expression macro is inferred. The result of the expression evaluation (see Section 3.3 [expression syntax], page 19) is written to the current output.

3.6.13 FOR - Emit a template block multiple times

This macro has a slight variation on the standard syntax:

     FOR <value-name> [ <separator-string> ]
     FOR <value-name> (...Scheme expression list)
     FOR <value-name> IN "string" [ ... ]

Other than for the last form, the first macro argument must be the name of an AutoGen value. If there is no value associated with the name, the FOR template block is skipped entirely. The scope of the FOR macro extends to the corresponding ENDFOR macro. The last form will create an array of string values named <value-name> that only exists within the context of this FOR loop. With this form, in order to use a separator-string, you must code it into the end of the template block using the (last-for?) predicate function (see Section 3.4.19 [SCM last-for?], page 26).

If there are any arguments after the value-name, the initial characters are used to determine the form. If the first character is either a semi-colon (;) or an opening parenthesis ((), then it is presumed to be a Scheme expression containing the FOR macro specific functions for-from, for-by, for-to, and/or for-sep. See Section 3.4 [AutoGen Functions], page 22. If it consists of an ’i’ an ’n’ and separated by white space from more text, then the FOR x IN form is processed. Otherwise, the remaining text is presumed to be a string for inserting between each iteration of the loop. This string will be emitted one time less than the number of iterations of the loop. That is, it is emitted after each loop, excepting for the last iteration.

If the from/by/to functions are invoked, they will specify which copies of the named value are to be processed. If there is no copy of the named value associated with a particular index, the FOR template block will be instantiated anyway. The template must use methods for detecting missing definitions and emitting default text. In this fashion, you can insert entries from a sparse or non-zero based array into a dense, zero based array.

NB: the for-from, for-to, for-by and for-sep functions are disabled outside of the context of the FOR macro. Likewise, the first-for, last-for and for-index functions are disabled outside of the range of a FOR block.

Also: the <value-name> must be a single level name, not a compound name (see Section 3.2 [naming values], page 19).

     [+FOR var (for-from 0) (for-to <number>) (for-sep ",") +]
     ... text with various substitutions ...[+
     ENDFOR var+]
this will repeat the ... text with various substitutions ... <number>+1 times. Each

repetition, except for the last, will have a comma , after it.

     [+FOR var ",\n" +]
     ... text with various substitutions ...[+
     ENDFOR var +]

This will do the same thing, but only for the index values of var that have actually been defined.

3.6.14 IF - Conditionally Emit a Template Block

Conditional block. Its arguments are evaluated (see Section 3.6.12 [EXPR], page 49) and if the result is non-zero or a string with one or more bytes, then the condition is true and the text from that point until a matched ELIF, ELSE or ENDIF is emitted. ELIF introduces a conditional alternative if the IF clause evaluated FALSE and ELSE introduces an unconditional alternative.

     [+IF <full-expression> +]
     emit things that are for the true condition[+
     ELIF <full-expression-2> +]
     emit things that are true maybe[+
     ELSE "This may be a comment" +]
     emit this if all but else fails[+
     ENDIF "This may *also* be a comment" +]

<full-expression> may be any expression described in the EXPR expression function, including the use of apply-codes and value-names. If the expression yields an empty string, it is interpreted as false.

3.6.15 INCLUDE - Read in and emit a template block

The entire contents of the named file is inserted at this point. The contents of the file are processed for macro expansion. The arguments are eval-ed, so you may compute the name of the file to be included. The included file must not contain any incomplete function blocks. Function blocks are template text beginning with any of the macro functions ‘CASE’, ‘DEFINE’, ‘FOR’, ‘IF’ and ‘WHILE’; extending through their respective terminating macro functions.

3.6.16 INVOKE - Invoke a User Defined Macro

User defined macros may be invoked explicitly or implicitly. If you invoke one implicitly, the macro must begin with the name of the defined macro. Consequently, this may not be a computed value. If you explicitly invoke a user defined macro, the macro begins with the macro name INVOKE followed by a basic expression that must yield a known user defined macro. A macro name must be found, or AutoGen will issue a diagnostic and exit.

Arguments are passed to the invoked macro by name. The text following the macro name must consist of a series of names each of which is followed by an equal sign (=) and a basic expression that yields a string.

The string values may contain template macros that are parsed the first time the macro is processed and evaluated again every time the macro is evaluated.

3.6.17 SELECT - Selection block for CASE function

This macro selects a block of text by matching an expression against the sample text expression evaluated in the CASE macro. See Section 3.6.2 [CASE], page 47.

You do not specify a SELECT macro with the word “select”. Instead, you must use one of the 19 match operators described in the CASE macro description.

3.6.18 UNKNOWN - Either a user macro or a value name.

The macro text has started with a name not known to AutoGen. If, at run time, it turns out to be the name of a defined macro, then that macro is invoked. If it is not, then it is a conditional expression that is evaluated only if the name is defined at the time the macro is invoked.

You may not specify UNKNOWN explicitly.

3.6.19 WHILE - Conditionally loop over a Template Block

Conditionally repeated block. Its arguments are evaluated (see Section 3.6.12 [EXPR], page 49) and as long as the result is non-zero or a string with one or more bytes, then the condition is true and the text from that point until a matched ENDWHILE is emitted.

[+WHILE <full-expression> +]
emit things that are for the true condition[+
ENDWHILE +]

<full-expression> may be any expression described in the EXPR expression function, including the use of apply-codes and value-names. If the expression yields an empty string, it is interpreted as false.

3.7 Redirecting Output

AutoGen provides a means for redirecting the template output to different files or, in ‘M4’ parlance, to various diversions. It is accomplished by providing a set of Scheme functions named out-* (see Section 3.4 [AutoGen Functions], page 22).

‘out-push-new (see Section 3.4.31 [SCM out-push-new], page 28)’

This allows you to logically "push" output files onto a stack. If you supply a string name, then a file by that name is created to hold the output. If you do not supply a name, then the text is written to a scratch pad and retrieved by passing a “#t” argument to the out-pop (see Section 3.4.29 [SCM out-pop], page 28) function.

‘out-pop (see Section 3.4.29 [SCM out-pop], page 28)’

This function closes the current output file and resumes output to the next one in the stack. At least one output must have been pushed onto the output stack with the out-push-new (see Section 3.4.31 [SCM out-push-new], page 28) function. If “#t” is passed in as an argument, then the entire contents of the diversion (or file) is returned.

‘out-suspend (see Section 3.4.33 [SCM out-suspend], page 29)’

This function does not close the current output, but instead sets it aside for resumption by the given name with out-resume. The current output must have been pushed on the output queue with out-push-new (see Section 3.4.31 [SCM out-push-new], page 28).

‘out-resume (see Section 3.4.32 [SCM out-resume], page 28)’

This will put a named file descriptor back onto the top of stack so that it becomes the current output again.

‘out-switch (see Section 3.4.34 [SCM out-switch], page 29)’

This closes the current output and creates a new file, purging any preexisting one. This is a shortcut for "pop" followed by "push", but this can also be done at the base level.

‘out-move (see Section 3.4.27 [SCM out-move], page 27)’

Renames the current output file without closing it.

There are also several functions for determining the output status. See Section 3.4 [AutoGen Functions], page 22.

4 Augmenting AutoGen Features

AutoGen was designed to be simple to enhance. You can do it by providing shell commands, Guile/Scheme macros or callout functions that can be invoked as a Guile macro. Here is how you do these.

4.1 Shell Output Commands

Shell commands are run inside of a server process. This means that, unlike ‘make’, context is kept from one command to the next. Consequently, you can define a shell function in one place inside of your template and invoke it in another. You may also store values in shell variables for later reference. If you load functions from a file containing shell functions, they will remain until AutoGen exits.

If your shell script should determine that AutoGen should stop processing, the recommended method for stopping AutoGen is:

     die "some error text"

That is a shell function added by AutoGen. It will send a SIGTERM to autogen and exit from the "persistent" shell.

4.2 Guile Macros

Guile also maintains context from one command to the next. This means you may define functions and variables in one place and reference them elsewhere. You also may load Guile macro definitions from a Scheme file by using the --load-scheme command line option (see Section 5.7 [autogen load-scheme], page 59). Beware, however, that the AutoGen specific scheme functions have not been loaded at this time, so though you may define functions that reference them, do not invoke the AutoGen functions at this time.

If your Scheme script should determine that AutoGen should stop processing, the recommended method for stopping AutoGen is:

     (error "some error text")

4.3 Guile Callout Functions

Callout functions must be registered with Guile to work. This can be accomplished either by putting your routines into a shared library that contains a void scm_init( void ) routine that registers these routines, or by building them into AutoGen.

To build them into AutoGen, you must place your routines in the source directory and name the files ‘exp*.c’. You also must have a stylized comment that ‘getdefs’ can find that conforms to the following:

     /*=gfunc <function-name>
      *
      *  what:    <short one-liner>
      *  general_use:
      *  string:  <invocation-name-string>
      *  exparg:  <name>, <description> [, ['optional'] [, 'list']]
      *  doc:     A long description telling people how to use

      *           this function.
     =*/
     SCM
     ag_scm_<function-name>( SCM arg_name[, ...] )
     { <code> }

‘gfunc’ You must have this exactly thus.

‘<function-name>’

This must follow C syntax for variable names

‘<short one-liner>’

This should be about a half a line long. It is used as a subsection title in this document.

‘general_use:’

You must supply this unless you are an AutoGen maintainer and are writing a function that queries or modifies the state of AutoGen.

‘<invocation-name-string>’

Normally, the function-name string will be transformed into a reasonable invocation name. However, that is not always true. If the result does not suit your needs, then supply an alternate string.

‘exparg:’ You must supply one for each argument to your function. All optional arguments must be last. The last of the optional arguments may be a list, if you choose.

‘doc:’ Please say something meaningful.

‘[, ...]’ Do not actually specify an ANSI ellipsis here. You must provide for all the arguments you specified with exparg.

See the Guile documentation for more details. More information is also available in a large comment at the beginning of the ‘agen5/snarf.tpl’ template file.

4.4 AutoGen Macros

There are two kinds those you define yourself and AutoGen native. The user-defined macros may be defined in your templates or loaded with the --lib-template option (See Section 3.6.4 [DEFINE], page 48 and Section 5.4 [autogen lib-template], page 58).

As for AutoGen native macros, do not add any. It is easy to do, but I won’t like it. The basic functions needed to accomplish looping over and selecting blocks of text have proved to be sufficient over a period of several years. New text transformations can be easily added via any of the AutoGen extension methods, as discussed above.

5 Invoking autogen

AutoGen creates text files from templates using external definitions. The definitions file (‘<def-file>’) can be specified with the ‘definitions’ option or as the command argument, but not both. Omitting it or specifying ‘-’ will result in reading definitions from standard input.

The output file names are based on the template, but generally use the base name of the definition file. If standard in is read for the definitions, then ‘stdin’ will be used for that base name. The suffixes to the base name are gotten from the template. However, the template file may specify the entire output file name. The generated files are always created in the current directory. If you need to place output in an alternate directory, ‘cd’ to that directory and use the ‘–templ dirs’ option to search the original directory.

‘loop-limit’ is used in debugging to stop runaway expansions.

This chapter was generated by AutoGen, the aginfo template and the option descriptions for the autogen program. It documents the autogen usage text and option meanings.

This software is released under the GNU General Public License.

5.1 autogen usage help (-?)

This is the automatically generated usage text for autogen:

autogen (GNU AutoGen) - The Automated Program Generator - Ver. 5.8.6
USAGE:  autogen [ -<flag> [<val>] | --<name>[{=| }<val>] ]... [ <def-file> ]
  Flg Arg Option-Name    Description
   -L Str templ-dirs     Template search directory list
                                - may appear multiple times
   -T Str override-tpl   Override template file
                                - may not be preset
   -l Str lib-template   Library template file
                                - may appear multiple times
   -b Str base-name      Base name for output file(s)
                                - may not be preset
      Str definitions    Definitions input file
                                - disabled as --no-definitions
                                - enabled by default
                                - may not be preset
   -S Str load-scheme    Scheme code file to load
   -F Str load-functions Load scheme function library
   -s Str skip-suffix    Omit the file with this suffix
                                - may not be preset
                                - may appear multiple times
   -o opt select-suffix  specify this output suffix
                                - may not be preset
                                - may appear multiple times
      no  source-time    set mod times to latest source
                                - disabled as --no-source-time
   -m no  no-fmemopen    Do not use in-mem streams

      Str equate         characters considered equivalent
      no  writable       Allow output files to be writable
                                - disabled as --not-writable
                                - may not be preset
The following options are often useful while debugging new templates:
  Flg Arg Option-Name    Description
      Num loop-limit     Limit on increment loops
                                  it must lie in one of the ranges:
                                  -1 exactly, or
                                  1 to 16777216
   -t Num timeout        Time limit for servers
                                  it must lie in the range: 0 to 3600
      KWd trace          tracing level of detail
      Str trace-out      tracing output file or filter
These options can be used to control what gets processed
in the definitions files and template files.
  Flg Arg Option-Name    Description
   -D Str define         name to add to definition list
                                - may appear multiple times
   -U Str undefine       definition list removal pattern
                                - an alternate for define
version and help options:
  Flg Arg Option-Name    Description
   -v opt version        Output version information and exit
   -? no  help           Display usage information and exit
   -! no  more-help      Extended usage information passed thru pager
   -> opt save-opts      Save the option state to a config file
   -< Str load-opts      Load options from a config file
                                - disabled as --no-load-opts
                                - may appear multiple times
Options are specified by doubled hyphens and their name
or by a single hyphen and the flag character.
AutoGen creates text files from templates using external definitions.
The following option preset mechanisms are supported:
 - reading file /dev/null
 - reading file /home/bkorb/ag/ag/agen5/.autogenrc
 - examining environment variables named AUTOGEN_*

The valid "trace" option keywords are:
  nothing server-shell templates block-macros expressions everything
The definitions file (`<def-file>') can be specified with the
`definitions' option or as the command argument, but not both.
Omitting it or specifying `-' will result in reading definitions from
standard input.
The output file names are based on the template, but generally use the
base name of the definition file.  If standard in is read for the
definitions, then `stdin' will be used for that base name.  The
suffixes to the base name are gotten from the template.  However, the
template file may specify the entire output file name.  The generated
files are always created in the current directory.  If you need to
place output in an alternate directory, `cd' to that directory and use
the `--templ_dirs' option to search the original directory.
`loop-limit' is used in debugging to stop runaway expansions.
please send bug reports to:  autogen-users@lists.sourceforge.net

5.2 templ-dirs option (-L)

This is the “template search directory list” option.

This option has some usage constraints. It:

• may appear an unlimited number of times.

Add a directory to the list of directories to search when opening a template, either as the primary template or an included one. The last entry has the highest priority in the search list. That is to say, they are searched in reverse order.

5.3 override-tpl option (-T)

This is the “override template file” option.

This option has some usage constraints. It:

• may not be preset with environment variables or in initialization (rc) files.

Definition files specify the standard template that is to be expanded. This option will override that name and expand a different template.

5.4 lib-template option (-l)

This is the “library template file” option.

This option has some usage constraints. It:

• may appear an unlimited number of times.

DEFINE macros are saved from this template file for use in processing the main macro file. Template text aside from the DEFINE macros is is ignored.

5.5 base-name option (-b)

This is the “base name for output file(s)” option.

This option has some usage constraints. It:

• may not be preset with environment variables or in initialization (rc) files.

A template may specify the exact name of the output file. Normally, it does not. Instead, the name is composed of the base name of the definitions file with suffixes appended. This option will override the base name derived from the definitions file name. This is required if there is no definitions file and advisable if definitions are being read from stdin. If the definitions are being read from standard in, the base name defaults to ‘stdin’. Any leading directory components in the name will be silently removed. If you wish the output file to appear in a particular directory, it is recommended that you "cd" into that directory first, or use directory names in the format specification for the output suffix lists, See Section 3.1 [pseudo macro], page 17.

5.6 definitions option

This is the “definitions input file” option.

This option has some usage constraints. It:

• is enabled by default.

• may not be preset with environment variables or in initialization (rc) files.

Use this argument to specify the input definitions file with a command line option. If you do not specify this option, then there must be a command line argument that specifies the file, even if only to specify stdin with a hyphen (-). Specify, --no-definitions when you wish to process a template without any active AutoGen definitions.\n

5.7 load-scheme option (-S)

This is the “scheme code file to load” option. Use this option to pre-load Scheme scripts into the Guile interpreter before template processing begins. Please note that the AutoGen specific functions are not loaded until after argument processing. So, though they may be specified in lambda functions you define, they may not be invoked until after option processing is complete.

5.8 load-functions option (-F)

This is the “load scheme function library” option.

This option has some usage constraints. It:

• must be compiled in by defining HAVE_DLOPEN during the compilation.

This option is used to load Guile-scheme functions. The automatically called initialization routine scm_init must be used to register these routines or data. This routine can be generated by using the following command and the ‘snarf.tpl’ template. Read the introductory comment in ‘snarf.tpl’ to see what the ‘getdefs(1AG)’ comment must contain.

First, create a config file for getdefs, and then invoke getdefs loading that file:

        cat > getdefs.cfg <<EOF
        subblock    exparg=arg_name,arg_desc,arg_optional,arg_list
        defs-to-get gfunc
        template    snarf
        srcfile
        linenum
        assign      group = name_of_some_group
        assign      init  = _init
        EOF
        getdefs load=getdefs.cfg <<source-file-list>>

Note, however, that your functions must be named:

        name_of_some_group_scm_<<function_name>>(...)

so you may wish to use a shorter group name.

5.9 skip-suffix option (-s)

This is the “omit the file with this suffix” option.

This option has some usage constraints. It:

• may appear an unlimited number of times.

• may not be preset with environment variables or in initialization (rc) files.

Occasionally, it may not be desirable to produce all of the output files specified in the template. (For example, only the ‘.h’ header file, but not the ‘.c’ program text.) To do this specify --skip-suffix=c on the command line.

5.10 select-suffix option (-o)

This is the “specify this output suffix” option.

This option has some usage constraints. It:

• may appear an unlimited number of times.

• may not be preset with environment variables or in initialization (rc) files.

If you wish to override the suffix specifications in the template, you can use one or more copies of this option. See the suffix specification in the Section 3.1 [pseudo macro], page 17 section of the info doc.

5.11 source-time option

This is the “set mod times to latest source” option. If you stamp your output files with the ‘DNE’ macro output, then your output files will always be different, even if the content has not really changed. If you use this option, then the modification time of the output files will change only if the input files change. This will help reduce unneeded builds.

5.12 no-fmemopen option (-m)

This is the “do not use in-mem streams” option.

This option has some usage constraints. It:

• must be compiled in by defining ENABLE_FMEMOPEN during the compilation.

If the local C library supports "fopencookie(3GNU)", or "funopen(3BSD)" then Auto-Gen prefers to use in-memory stream buffer opens instead of anonymous files. This may lead to problems if there is a shortage of virtual memory. If, for a particular application, you run out of memory, then specify this option. This is unlikely in a modern virtual memory environment.

5.13 equate option

This is the “characters considered equivalent” option. This option will alter the list of characters considered equivalent. The default are the three characters, " -^". (The last is conventional on a Tandem/HP-NonStop, and I used to do a lot of work on Tandems.)

5.14 writable option

This is the “allow output files to be writable” option.

This option has some usage constraints. It:

• may not be preset with environment variables or in initialization (rc) files.

This option will leave output files writable. Normally, output files are read-only.

5.15 loop-limit option

This is the “limit on increment loops” option. This option prevents runaway loops. For example, if you accidentally specify, "FOR x (for-from 1) (for-to -1) (for-by 1)", it will take a long time to finish. If you do have more than 256 entries in tables, you will need to specify a new limit with this option.

5.16 timeout option (-t)

This is the “time limit for servers” option.

This option has some usage constraints. It:

• must be compiled in by defining SHELL_ENABLED during the compilation.

AutoGen works with a shell server process. Most normal commands will complete in less than 10 seconds. If, however, your commands need more time than this, use this option.

The valid range is 0 to 3600 seconds (1 hour). Zero will disable the server time limit.

5.17 trace option

This is the “tracing level of detail” option.

This option has some usage constraints. It:

• This option takes a keyword as its argument. The argument sets an enumeration value that can be tested by comparing the option value macro (OPT VALUE TRACE). The available keywords are:

           nothing      server-shell templates
           block-macros expressions  everything

This option will cause AutoGen to display a trace of its template processing. There are six levels, each level including messages from the previous levels:

‘nothing’ Does no tracing at all (default)

‘server-shell’

Traces all input and output to the server shell. This includes a shell "independent" initialization script about 30 lines long. Its output is discarded and not inserted into any template.

‘templates’

Traces the invocation of DEFINEd macros and INCLUDEs

‘block-macros’

Traces all block macros. The above, plus IF, FOR, CASE and WHILE.

‘expressions’

Displays the results of expression evaluations.

‘everything’

Displays the invocation of every AutoGen macro, even TEXT macros (i.e. the text outside of macro quotes).

5.18 trace-out option

This is the “tracing output file or filter” option. The output specified may be either a file name, or, if the option argument begins with the pipe operator (|), a command that will receive the tracing output as standard in. For example, --traceout='| less' will run the trace output through the less program.

5.19 show-defs option

This is the “show the definition tree” option.

This option has some usage constraints. It:

• must be compiled in by defining DEBUG_ENABLED during the compilation.

• may not be preset with environment variables or in initialization (rc) files.

This will print out the complete definition tree before processing the template.

5.20 define option (-D)

This is the “name to add to definition list” option.

This option has some usage constraints. It:

• may appear an unlimited number of times.

The AutoGen define names are used for the following purposes:

1. Sections of the AutoGen definitions may be enabled or disabled by using C-style #ifdef and #ifndef directives.

2. When defining a value for a name, you may specify the index for a particular value. That index may be a literal value, a define option or a value #define-d in the definitions themselves.

3. The name of a file may be prefixed with $NAME/. The $NAME part of the name string will be replaced with the define-d value for NAME.

4. When AutoGen is finished loading the definitions, the defined values are exported to the environment with, putenv(3). These values can then be used in shell scripts with ${NAME} references and in templates with (getenv "NAME").

5. While processing a template, you may specify an index to retrieve a specific value. That index may also be a define-d value.

5.21 undefine option (-U)

This is the “definition list removal pattern” option.

This option has some usage constraints. It:

• may appear an unlimited number of times.

• may not be preset with environment variables or in initialization (rc) files.

Just like ’C’, AutoGen uses #ifdef/#ifndef preprocessing directives. This option will cause the matching names to be removed from the list of defined values.

6 Configuring and Installing

6.1 Configuring AutoGen

AutoGen is configured and built using Libtool, Automake and Autoconf. Consequently, you can install it wherever you wish using the various ‘--prefix’ options. To the various configuration options supplied by these tools, AutoGen adds a few of its own:

‘--disable-shell’

AutoGen is now capable of acting as a CGI forms server, See Section 6.2 [AutoGen CGI], page 65. As such, it will gather its definitions using either ‘GET’ or ‘POST’ methods. All you need to do is have a template named ‘cgi.tpl’ handy or specify a different one with a command line option. However, doing this without disabling the server shell brings considerable risk. If you were to pass user input to a script that contained, say, the classic "‘`rm -rf /`’", you might have a problem. This configuration option will cause shell template commands to simply return the command string as the result. No mistakes. Much safer. Strongly recommended. The default is to have server shell scripting enabled. Disabling the shell will have some build side effects, too.

• Many of the make check tests will fail, since they assume a working server shell.

• The getdefs and columns programs are not built. The options are distributed as definition files and they cannot be expanded with a shelldisabled AutoGen.

• Similarly, the documentation cannot be regenerated because the documentation templates depend on subshell functionality.

‘--enable-debug’

Turning on AutoGen debugging enables very detailed inspection of the input definitions and monitoring shell script processing. These options are not particularly useful to anyone not directly involved in maintaining AutoGen. If you do choose to enable AutoGen debugging, be aware that the usage page was generated without these options, so when the build process reaches the documentation rebuild, there will be a failure. ‘cd’ into the ‘agen5’ build directory, ‘make’ the ‘autogen.texi’ file and all will be well thereafter.

‘--with-regex-header’
‘--with-header-path’
‘--with-regex-lib’

These three work together to specify how to compile with and link to a particular POSIX regular expression library. The value for ‘--with-regex-header=value’ must be the name of the relevant header file. The AutoGen sources will attempt to include that source with a #include <value> C preprocessing statement. The path from the

‘--with-header-path=path’ will be added to CPPFLAGS as ‘-Ipath’. The
lib-specs from ‘--with-regex-lib=lib-specs’ will be added to LDFLAGS

without any adornment.

6.2 AutoGen as a CGI server

AutoGen is now capable of acting as a CGI forms server. It behaves as a CGI server if the definitions input is from stdin and the environment variable REQUEST_METHOD is defined and set to either "GET" or "POST". If set to anything else, AutoGen will exit with a failure message. When set to one of those values, the CGI data will be converted to AutoGen definitions (see Chapter 2 [Definitions File], page 6) and the template named "cgi.tpl" will be processed.

This works by including the name of the real template to process in the form data and having the "cgi.tpl" template include that template for processing. I do this for

processing the form http://autogen.sourceforge.net/conftest.html. The "cgi.tpl"

looks approximately like this:

        <? AutoGen5 Template ?>
        <?
        IF (not (exist? "template"))                       ?><?
          form-error                                       ?><?
        ELIF (=* (get "template") "/")                     ?><?
          form-error                                       ?><?
        ELIF (define tpl-file (string-append "cgi-tpl/"
                              (get "template")))
             (access? tpl-file R_OK)                       ?><?
          INCLUDE (. tpl-file)                             ?><?
        ELIF (set! tpl-file (string-append tpl-file ".tpl"))
             (access? tpl-file R_OK)                       ?><?
          INCLUDE (. tpl-file)                             ?><?
        ELSE                                               ?><?
          form-error                                       ?><?
        ENDIF                                              ?>

This forces the template to be found in the "cgi-tpl/" directory. Note also that there is no suffix specified in the pseudo macro (see Section 3.1 [pseudo macro], page 17). That tells AutoGen to emit the output to stdout.

The output is actually spooled until it is complete so that, in the case of an error, the output can be discarded and a proper error message can be written in its stead.

Please also note that it is advisable, especially for network accessible machines, to configure AutoGen (see Section 6.1 [configuring], page 64) with shell processing disabled (-- disable-shell). That will make it impossible for any referenced template to hand data to a subshell for interpretation.

6.3 Signal Names

When AutoGen is first built, it tries to use psignal(3), sys_siglist, strsigno(3) and strsignal(3) from the host operating system. If your system does not supply these, the AutoGen distribution will. However, it will use the distributed mapping and this mapping is unlikely to match what your system uses. This can be fixed. Once you have installed autogen, the mapping can be rebuilt on the host operating system. To do so, you must perform the following steps:

1. Build and install AutoGen in a place where it will be found in your search path.

 2. cd ${top_srcdir}/compat
 3. autogen strsignal.def

4. Verify the results by examining the ‘strsignal.h’ file produced.

5. Re-build and re-install AutoGen.

If you have any problems or peculiarities that cause this process to fail on your platform, please send me copies of the header files containing the signal names and numbers, along with the full path names of these files. I will endeavor to fix it. There is a shell script inside of ‘strsignal.def’ that tries to hunt down the information.

6.4 Installing AutoGen

There are several files that get installed. The number depend whether or not both shared and archive libraries are to be installed. The following assumes that everything is installed relative to $prefix. You can, of course, use configure to place these files where you wish.

NB AutoGen does not contain any compiled-in path names. All support directories are located via option processing, the environment variable HOME or finding the directory where the executable came from.

The installed files are:

1. The executables in ‘bin’ (autogen, getdefs and columns).

2. The AutoOpts link libraries as ‘lib/libopts.*’.

3. An include file in ‘include/options.h’, needed for Automated Option Processing (see next chapter).

4. Several template files and a scheme script in ‘share/autogen’, needed for Automated Option Processing (see Chapter 7 [AutoOpts], page 68), parsing definitions written with scheme syntax (see Section 2.4 [Dynamic Text], page 10), the templates for producing documentation for your program (see Section 7.5.6 [documentation attributes], page 90), autoconf test macros, and AutoFSM.

5. Info-style help files as ‘info/autogen.info*’. These files document AutoGen, the option processing library AutoOpts, and several add-on components.

6. The three man pages for the three executables are installed in man/man1.

This program, library and supporting files can be installed with three commands:

• <src-dir>/configure [ <configure-options> ]

• make

• make install

However, you may wish to insert make check before the make install command.

If you do perform a make check and there are any failures, you will find the results in <module>/test/FAILURES. Needless to say, I would be interested in seeing the contents of those files and any associated messages. If you choose to go on and analyze one of these failures, you will need to invoke the test scripts individually. You may do so by specifying the test (or list of test) in the TESTS make variable, thus:

        gmake TESTS=test-name.test check

I specify gmake because most makes will not let you override internal definitions with command line arguments. gmake does.

All of the AutoGen tests are written to honor the contents of the VERBOSE environment variable. Normally, any commentary generated during a test run is discarded unless the VERBOSE environment variable is set. So, to see what is happening during the test, you might invoke the following with bash or ksh:

        VERBOSE=1 gmake TESTS="for.test forcomma.test" check

Or equivalently with csh:

        env VERBOSE=1 gmake TESTS="for.test forcomma.test" check

7 Automated Option Processing

AutoOpts 27.4 is bundled with AutoGen. It is a tool that virtually eliminates the hassle of processing options and keeping man pages, info docs and usage text up to date. This package allows you to specify several program attributes, up to a hundred option types and many option attributes. From this, it then produces all the code necessary to parse and handle the command line and configuration file options, and the documentation that should go with your program as well.

All the features notwithstanding, some applications simply have well-established command line interfaces. Even still, those programs may use the configuration file parsing portion of the library. See the “AutoOpts Features” and “Configuration File Format” sections.

7.1 AutoOpts Features

AutoOpts supports option processing; option state saving; and program documentation with innumerable features. Here, we list a few obvious ones and some important ones, but the full list is really defined by all the attributes defined in the Section 7.5 [Option Definitions], page 73 section.

1. POSIX-compliant short (flag) option processing.

2. GNU-style long options processing. Long options are recognized without case sensitivity, and they may be abbreviated.

3. Environment variable initializations, See Section 7.9.4 [environrc], page 114.

4. Initialization from configuration files (aka RC or INI files), and saving the option state back into one, See Section 7.9.1 [loading rcfile], page 114.

5. Config files may be partitioned. One config file may be used by several programs by partitioning it with lines containing, “[PROGRAM_NAME]”, See Section 7.9.1 [loading rcfile], page 114.

6. Options may be marked as dis-abled with a disablement prefix. Such options may default to either an enabled or a disabled state. You may also provide an enablement

    prefix, too, e.g., --allow-mumble and --prevent-mumble.

7. Verify that required options are present between the minimum and maximum number of times on the command line.

8. Verify that conflicting options do not appear together, and that options that require the presence of other options are, in fact, used in the presence of other options.

9. Provides a callable routine to parse a text string as if it were from one of the rc/ini/config files, hereafter referred to as a configuration file.

10. --help and --version are automatically supported. --more-help will page the generated help.

11. By adding a ‘doc’ and ‘arg-name’ attributes to each option, AutoGen will also be able to produce a man page and the ‘invoking’ section of a texinfo document.

12. Insert the option processing state into Scheme-defined variables. Thus, Guile based applications that are linked with private main() routines can take advantage of all of AutoOpts’ functionality.

13. Various forms of main procedures can be added to the output, See Section 7.5.3 [Generated main], page 76. There are four basic forms: a. A program that processes the arguments and writes to standard out portable shell

commands containing the digested options. b. A program that will generate portable shell commands to parse the defined options.

The expectation is that this result will be copied into a shell script and used there. c. A “for-each” main that will invoke a named function once for either each non-

option argument on the command line or, if there are none, then once for each non-blank, non-comment input line read from stdin.

d. A main procedure of your own design. Its code can be supplied in the option

description template or by incorporating another template.

14. Library suppliers can specify command line options that their client programs will accept. They specify option definitions that get #include-d into the client option definitions and they specify an "anchor" option that has a callback and must be invoked. That will give the library access to the option state for their options.

15. The generated usage text can be emitted in either AutoOpts standard format (maximizing the information about each option), or GNU-ish normal form. The default form is selected by either specifying or not specifying the gnu-usage attribute (see Section 7.5.4 [information attributes], page 80). This can be overridden by the user himself with the AUTOOPTS_USAGE environment variable. If it exists and is set to the string gnu, it will force GNU-ish style format; if it is set to the string autoopts, it will force AutoOpts standard format; otherwise, it will have no effect.

16. If you compile with ENABLE_NLS defined and _() defined to a localization function such as gettext(3GNU), then the option processing code will be localizable (see Section 7.15 [i18n], page 139).

17. Intermingled option processing. AutoOpts options may be intermingled with command line operands and options processed with other parsing techniques. This is accomplished by setting the allow-errors (see Section 7.5.1 [program attributes], page 73) attribute. When processing reaches a point where optionProcess (see Section 7.6.28.11 [libopts-optionProcess], page 106) needs to be called again, the current option can be set with RESTART_OPT(n) (see Section 7.6.15 [RESTART OPT], page 97) before calling optionProcess. See: See Section 7.5.2 [library attributes], page 75.

18. library options. An AutoOpt-ed library may export its options for use in an AutoOpted program. This is done by providing an option definition file that client programs #include into their own option definitions. See “AutoOpt-ed Library for AutoOpt-ed Program” (see Section 7.5.2.1 [lib and program], page 75) for more details.

7.2 AutoOpts Licensing

When AutoGen is installed, the AutoOpts project is installed with it. AutoOpts includes various AutoGen templates and a pair of shared libraries. These libraries may be used under the terms of the GNU Lesser General Public License (LGPL).

One of these libraries (libopts) is needed by programs that are built using AutoOpts generated code. This library is available as a separate “tear-off” source tarball. It is redistributable for use under either of two licenses: The GNU Lesser General Public License ("Lesser" meaning you have greater license with it and may link it into commercial programs), and the advertising-clause-free BSD license. Both of these license terms are incorporated into appropriate COPYING files included with the libopts source tarball. This source may be incorporated into your package with the following simple commands:

        rm -rf libopts libopts-*
        gunzip -c `autoopts-config libsrc` | \
           tar -xvf -
        mv libopts-*.*.* libopts

View the ‘libopts/README’ file for further integration information.

7.3 Quick Start

Since it is generally easier to start with a simple example than it is to look at the options that AutoGen uses itself, here is a very simple AutoOpts example. You can copy this example out of the Info file and into a source file to try it. You can then embellish it into what you really need. For more extensive examples, you can also examine the help output and option definitions for the commands columns, getdefs and autogen itself.

For our simple example, assume you have a program named check that takes two options:

1. A list of directories to check over for whatever it is check does. You want this option available as a POSIX-style flag option and a GNU long option. You want to allow as many of these as the user wishes.

2. An option to show or not show the definition tree being used. Only one occurrence is to be allowed, specifying one or the other.

First, specify your program attributes and its options to AutoOpts, as with the following example.

        AutoGen Definitions options;
        prog-name     = check;
        prog-title    = "Checkout Automated Options";
        long-opts;
        main = { main-type = shell-process; };
        flag = {
            name      = check-dirs;
            value     = L;        /* flag style option character */
            arg-type  = string;   /* option argument indication  */
            max       = NOLIMIT;  /* occurrence limit (none)     */
            stack-arg;            /* save opt args in a stack    */
            descrip   = "Checkout directory list";
        };
        flag = {
            name      = show_defs;
            descrip   = "Show the definition tree";
            disable   = dont;     /* mark as enable/disable type */
                                  /* option.  Disable as `dont-' */
        };

Then perform the following steps:

 1. cflags="-DTEST_CHECK_OPTS `autoopts-config cflags`"
 2. ldflags="`autoopts-config ldflags`"
 3. autogen checkopt.def
 4. cc -o check -g ${cflags} checkopt.c ${ldflags}
 5. ./check --help

Running those commands yields:

        check - Checkout Automated Options
        USAGE:  check [ -<flag> [<val>] | --<name>[{=| }<val>] ]...
          Flg Arg Option-Name    Description
           -L Str check-dirs     Checkout directory list
                                        - may appear multiple times
              no  show-defs      Show the definition tree
                                        - disabled as --dont-show-defs
           -? no  help           Display usage information and exit
           -! no  more-help      Extended usage information passed thru pager
        Options are specified by doubled hyphens and their name
        or by a single hyphen and the flag character.

Normally, however, you would compile ‘checkopt.c’ as in:

        cc -o checkopt.o -I$prefix/include -c checkopt.c

and link ‘checkopt.o’ with the rest of your program. The main program causes the options to be processed by calling optionProcess (see Section 7.6.28.11 [libopts-optionProcess], page 106):

        main( int argc, char** argv )
        {
          {
            int optct = optionProcess( &checkOptions, argc, argv );
            argc -= optct;
            argv += optct;
          }

The options are tested and used as in the following fragment. “ENABLED_OPT” is used instead of “HAVE_OPT” for the show-defs option because it is an enabled/disabled option type:

          if (  ENABLED_OPT( SHOW_DEFS )
             && HAVE_OPT( CHECK_DIRS )) {
            int    dirct = STACKCT_OPT( CHECK_DIRS );
            char** dirs  = STACKLST_OPT( CHECK_DIRS );
            while (dirct-- > 0) {
              char* dir = *dirs++;
              ...

A lot of magic happens to make this happen. The rest of this chapter will describe the myriad of option attributes supported by AutoOpts. However, keep in mind that, in general, you won’t need much more than what was described in this "quick start" section.

7.4 Multi-Threading

AutoOpts was designed to configure a program for running. This generally happens before much real work has been started. Consequently, it is expected to be run before multithreaded applications have started multiple threads. However, this is not always the case. Some applications may need to reset and reload their running configuration, and some may use SET_OPT_xxx() macros during processing. If you need to dynamically change your option configuration in your multi-threaded application, it is your responsibility to prevent all threads from accessing the option configuration state, except the one altering the configuration.

The various accessor macros (HAVE_OPT(), etc.) do not modify state and are safe to use in a multi-threaded application. It is safe as long as no other thread is concurrently modifying state, of course.

7.5 Option Definitions

AutoOpts uses an AutoGen definitions file for the definitions of the program options and overall configuration attributes. The complete list of program and option attributes is quite extensive, so if you are reading to understand how to use AutoOpts, I recommend reading the "Quick Start" section (see Section 7.3 [Quick Start], page 71) and paying attention to the following:

1. prog-name, prog-title, and argument, program attributes, See Section 7.5.1 [program attributes], page 73.

2. name and descrip option attributes, See Section 7.5.5.1 [Required Attributes], page 82.

3. value (flag character) and min (occurrence counts) option attributes, See Section 7.5.5.2 [Common Attributes], page 83.

4. arg-type from the option argument specification section, See Section 7.5.5.6 [Option Arguments], page 86.

5. Read the overall how to, See Section 7.8 [Using AutoOpts], page 111.

6. Highly recommended, but not required, are the several "man" and "info" documentation attributes, See Section 7.5.6 [documentation attributes], page 90.

Keep in mind that the majority are rarely used and can be safely ignored. However, when you have special option processing requirements, the flexibility is there.

7.5.1 Program Description Attributes

The following global definitions are used to define attributes of the entire program. These generally alter the configuration or global behavior of the AutoOpts option parser. The first two are required of every program. The rest have been alphabetized. Except as noted, there may be only one copy of each of these definitions:

‘prog-name’

This attribute is required. Variable names derived from this name are derived using string->c_name! (see Section 3.5.53 [SCM string->c-name!], page 44).

‘prog-title’

This attribute is required and may be any descriptive text.

‘allow-errors’

The presence of this attribute indicates ignoring any command line option errors. This may also be turned on and off by invoking the macros ERRSKIP_ OPTERR and ERRSTOP_OPTERR from the generated interface file.

‘argument’

Specifies the syntax of the arguments that follow the options. It may not be empty, but if it is not supplied, then option processing must consume all the arguments. If it is supplied and starts with an open bracket ([), then there is no requirement on the presence or absence of command line arguments following the options. Lastly, if it is supplied and does not start with an open bracket, then option processing must not consume all of the command line arguments.

‘environrc’

Indicates looking in the environment for values of variables named, PROGRAM_ OPTNAME or PROGRAM, where PROGRAM is the upper cased C-name of the program and OPTNAME is the upper cased C-name of a specific option.

‘export’ This string is inserted into the .h interface file. Generally used for global variables or #include directives required by flag_code text and shared with other program text. Do not specify your configuration header (‘config.h’) in this attribute or the include attribute, however. Instead, use config-header, below.

‘config-header’

The contents of this attribute should be just the name of the configuration file. A "#include" naming this file will be inserted at the top of the generated header.

‘homerc’ Specifies either a directory or a file using a specific path (like . or

‘/usr/local/share/progname’) or an environment variable (like ‘$HOME/rc/’

or ‘$PREFIX/share/progname’) or the directory where the executable was found (‘$$[/...]’) to use to try to find the rcfile. Use as many as you like. The presence of this attribute activates the --save-opts and --load-opts options. See Section 7.9.1 [loading rcfile], page 114.

‘include’ This string is inserted into the .c file. Generally used for global variables required only by flag_code program text.

‘long-opts’

Presence indicates GNU-standard long option processing. If any options do not have an option value (flag character) specified, and least one does specify such a value, then you must specify long-opts. If none of your options specify an option value (flag character) and you do not specify long-opts, then command line arguments are processed in "named option mode". This means that:

• Every command line argument must be a long option.

• The flag markers - and -- are completely optional.

• The argument program attribute is disallowed.

• One of the options may be specified as the default (as long as it has a required option argument).

‘prefix’ This value is inserted into all global names. This will disambiguate them if more than one set of options are to be compiled into a single program.

‘rcfile’ Specifies the configuration file name. This is only useful if you have provided at least one homerc attribute. default: .<prog-name>rc

‘version’ Specifies the program version and activates the VERSION option, See Section 7.5.7 [automatic options], page 90.

‘guard-option-names’

AutoOpts generates macros that presume that there are no cpp macros with the same name as the option name. For example, if you have an option named, debug, then you must not use #ifdef DEBUG in your code. If you specify this attribute, every option name will be guarded. If the name is #define-d, then a warning will be issued and the name undefined. If you do not specify this and there is a conflict, you will get strange error messages. This will produce code that will warn you when conflicts get hidden. The builder of your program may suppress these warnings by adding this define to the compile command line:

        -DNO_OPTION_NAME_WARNINGS

7.5.2 Options for Library Code

Some libraries provide their own code for processing command line options, and this may be used by programs that utilize AutoOpts. You may also wish to write a library that gets configured with AutoOpts options and config files. Such a library may either supply its own configury routine and process its own options, or it may export its option descriptions to programs that also use AutoOpts. This section will describe how to do all of these different things.

7.5.2.1 AutoOpt-ed Library for AutoOpt-ed Program

The library source code must provide an option definition file that consists of only the flag entries. The first flag entry must contain the following attributes:

‘name’ This name is used in the construction of a global pointer of type tOptDesc const*. It is always required.

‘documentation’

It tells AutoOpts that this option serves no normal purpose. It will be used to add usage clarity and to locate option descriptors in the library code.

‘descrip’ This is a string that is inserted in the extended usage display before the options specific to the current library. It is always required.

‘lib-name’

This should match the name of the library. This string is also used in the construction of the option descriptor pointer name. In the end, it looks like this:

        extern tOptDesc const* <<lib-name>>_<<name>>_optDesc_p;

and is used in the macros generated for the library’s .h file.

In order to compile this AutoOpts using library, you must create a special header that is not used by the client program. This is accomplished by creating an option definition file that contains essentially exactly the following:

AutoGen definitions options;
prog-name  = does-not-matter;  // but is always required
prog-title = 'also does not matter';  // also required
config-header = 'config.h'; // optional, but common

        library;
        #include library-options-only.def

and nothing else. AutoGen will produce only the .h file. You may now compile your library, referencing just this .h file. The macros it creates will utilize a global variable that will be defined by the AutoOpts-using client program. That program will need to have the following #include in its option definition file:

        #include library-options-only.def

All the right things will magically happen so that the global variables named <<lib-name>>_ <<name>>_optDesc_p are initialized correctly. For an example, please see the AutoOpts test

script: ‘autoopts/test/library.test’.

7.5.2.2 AutoOpt-ed Library for Regular Program

In this case, your library must provide an option processing function to a calling program. This is accomplished by setting the allow-errors global option attribute. Each time your option handling function is called, you must determine where your scan is to resume and tell the AutoOpts library by invoking:

        RESTART_OPT(next_arg_index);
and then invoke not_opt_index = optionProcess(...). The not_opt_index value can

be used to set optind, if that is the global being used to scan the program argument array.

In this method, do NOT utilize the global library attribute. Your library must specify its options as if it were a complete program. You may choose to specify an alternate usage() function so that usage for other parts of the option interface may be displayed as well. See “Program Information Attributes” (see Section 7.5.4 [information attributes], page 80).

At the moment, there is no method for calling optionUsage() telling it to produce just the information about the options and not the program as a whole. Some later revision after somebody asks.

7.5.2.3 AutoOpt-ed Program Calls Regular Library

As with providing an AutoOpt-ed library to a non-AutoOpt-ed program, you must write the option description file as if you were writing all the options for the program, but you should specify the allow-errors global option attribute and you will likely want an alternate usage() function (see “Program Information Attributes” see Section 7.5.4 [information attributes], page 80). In this case, though, when optionProcess() returns, you need to test to see if there might be library options. If there might be, then call the library’s exported routine for handling command line options, set the next-option-to-process with the RESTART_OPT() macro, and recall optionProcess(). Repeat until done.

7.5.3 Generating main procedures

When AutoOpts generates the code to parse the command line options, it has the ability to produce any of several types of main() procedures. This is done by specifying a global structured value for main. The values that it contains are dependent on the value set for the one value it must have: main-type.

The recognized values for main-type are:

Here is an example of an include variation:

        main = {
          main-type = include;
          tpl       = "main-template.tpl";
        };

7.5.3.1 guile: main and inner main procedures

When the main-type is specified to be guile, a main() procedure is generated that calls gh_enter(), providing it with a generated inner_main() to invoke. If you must perform certain tasks before calling gh_enter(), you may specify such code in the value for the

before-guile-boot attribute.

The inner_main() procedure itself will process the command line arguments (by calling optionProcess(), see Section 7.6.28.11 [libopts-optionProcess], page 106), and then either invoke the code specified with the guile-main attribute, or else export the parsed options to Guile symbols and invoke the scm_shell() function from the Guile library. This latter will render the program nearly identical to the stock guile(1) program.

7.5.3.2 shell-process: emit Bourne shell results

This will produce a main() procedure that parses the command line options and emits to stdout Bourne shell commands that puts the option state into environment variables. This can be used within a shell script as follows:

        unset OPTION_CT
        eval "`opt_parser \"$@\"`"
        test -z "${OPTION_CT}" && exit 1
        test ${OPTION_CT} -gt 0 && shift ${OPTION_CT}

If the option parsing code detects an error or a request for usage, it will not emit an assignment to OPTION CT and the script should just exit. If the options are set consistently, then something along the lines of the following will be written to stdout and evaled:

            OPTION_CT=4
            export OPTION_CT
            MYPROG_SECOND='first'
            export MYPROG_SECOND
            MYPROG_ANOTHER=1 # 0x1
            export MYPROG_ANOTHER

If the arguments are to be reordered, however, then the resulting set of operands will be emitted and OPTION_CT gets set to zero. For example, the following would be appended to the above:

            set -- 'operand1' 'operand2' 'operand3'
            OPTION_CT=0

OPTION_CT is set to zero since it is not necessary to shift off any options.

7.5.3.3 shell-parser: emit Bourne shell script

This will produce a main() procedure that emits a shell script that will parse the command line options. That script can be emitted to stdout or inserted or substituted into a preexisting shell script file. Improbable markers are used to identify previously inserted parsing text:

        # # # # # # # # # # -- do not modify this marker --

The program is also pretty insistent upon starting its parsing script on the second line.

7.5.3.4 main: user supplied main procedure

You must supply a value for the main-text attribute. You may also supply a value for option-code. If you do, then the optionProcess invocation will not be emitted into the code. AutoOpts will wrap the main-text inside of:

        int
        main( int argc, char** argv )
        {
            {
                int ct = optionProcess( &<<prog-name>>Options, argc, argv );
                argc -= ct;
                argv += ct;
            }
        <<your text goes here>>
        }

so you can most conveniently set the value with a “here string” (see Section 2.2.7 [herestring], page 8):

        code = <<- _EndOfMainProc_
        <<your text goes here>>
        _EndOfMainProc_;

7.5.3.5 include: code emitted from included template

You must write a template to produce your main procedure. You specify the name of the template with the tpl attribute and it will be incorporated at the point where AutoOpts is ready to emit the main() procedure.

This can be very useful if, in your working environment, you have many programs with highly similar main() procedures. All you need to do is parameterize the variations and specify which variant is needed within the main AutoOpts specification. Since you are coding the template for this, the attributes needed for this variation would be dictated by your template.

7.5.3.6 invoke: code emitted from AutoGen macro

You must write a template to produce your main procedure. That template must contain a definition for the function specified with the func attribute to this main() procedure specification. Typically, this template will be incorporated by using the --lib-template option (see Section 5.4 [autogen lib-template], page 58) in the AutoGen invocation. Otherwise, this variation operates in much the same way as “include” (see Section 7.5.3.5 [main include], page 78) method.

7.5.3.7 for-each: perform function on each argument

This produces a main procedure that invokes a procedure once for each operand on the command line (non-option arguments), OR once for each non-blank, non-comment stdin input line. Leading and trailing white space is trimmed from the input line and comment lines are lines that are empty or begin with a comment character, defaulting to a hash (’#’) character.

NB: The argument program attribute (see Section 7.5.1 [program attributes], page 73) must begin with the [ character, to indicate that there are command operands, but that they are optional.

There are a number of attributes to main that may be used:

handler-proc

This attribute is required. It is used to name the procedure to call. That procedure is presumed to be external, but if you provide the code for it, then the procedure is emitted as a static procedure in the generated code. This procedure should return 0 on success, a cumulative error code on warning and exit without returning on an unrecoverable error. As the cumulative warning codes are or-ed together, the codes should be some sort of bit mask in order to be ultimately decipherable (if you need to do that). If the called procedure needs to cause a fail-exit, it is expected to call exit(3) directly. If you want to cause a warning exit code, then this handler function should return a non-zero status. That value will be OR-ed into a result integer for computing the final exit code. E.g., here is part of the emitted code:

          int res = 0;
          if (argc > 0) {
             do  {
                 res |= my_handler( *(argv++) );
             } while (--argc > 0);
          } else { ...
handler-type

If you do not supply this attribute, your handler procedure must be the default type. The profile of the procedure must be:

        int my_handler( char const *pz_entry );

However, if you do supply this attribute, you may select any of three alternate flavors:

‘name-of-file’

This is essentially the same as the default handler type, except that before your procedure is invoked, the generated code has verified that the string names an existing file. The profile is unchanged.

‘file-X’ Before calling your procedure, the file is f-opened according to the “X”, where “X” may be any of the legal modes for fopen(3C). In this case, the profile for your procedure must be:

                  int my_handler( char const* pz_fname, FILE* entry_fp );
‘text-of-file’
‘some-text-of-file’

Before calling your procedure, the contents of the file are read into memory. (Excessively large files may cause problems.) The “‘some-text-of-file’” disallows empty files. Both require regular files. In this case, the profile for your procedure must be:

                  int my_handler( char const* pz_fname, char* file_text,
                                  size_t text_size );

Note that though the file_text is not const, any changes made to it are not written back to the original file. It is merely a memory image of the file contents. Also, the memory allocated to hold the text is text_size + 1 bytes long and the final byte is always NUL. The file contents need not be text, as the data are read with the read(2) system call.

my_handler-code

With this attribute, you provide the code for your handler procedure in the option definition file. In this case, your main() procedure specification might look something like this:

        main = {
          main-type    = for-each;
          handler-proc = my_handler;
          my_handler-code = <<- EndOfMyCode
        /* whatever you want to do */
        EndOfMyCode;
        };

and instead of an emitted external reference, a procedure will be emitted that looks like this:

        static int
        my_handler( char const* pz_entry )
        {
            int res = 0;
            <<my_handler-code goes here>>
            return res;
        }
main-init

This is code that gets inserted after the options have been processed, but before the handler procs get invoked.

main-fini

This is code that gets inserted after all the entries have been processed, just before returning from main().

comment-char

If you wish comment lines to start with a character other than a hash (#) character, then specify one character with this attribute. If that character is the NUL byte, then only blank lines will be considered comments.

7.5.4 Program Information Attributes

These attributes are used to define how and what information is displayed to the user of the program.

‘copyright’

The copyright is a structured value containing three to five values. If copyright is used, then the first three are required.

1. ‘date’ - the list of applicable dates for the copyright.

2. ‘owner’ - the name of the copyright holder.

3. ‘type’ - specifies the type of distribution license. AutoOpts/AutoGen will automatically support the text of the GNU Public License (‘GPL’), the GNU General Public License with Library extensions (‘LGPL’), the Free BSD license (‘BSD’), and a write-it-yourself copyright notice (‘NOTE’). Only these values are recognized.

4. ‘text’ - the text of the copyright notice. It is only needed if ‘type’ is set to ‘NOTE’.

5. ‘author’ - in case the author name is to appear in the documentation and is different from the copyright owner.

6. ‘eaddr’ - email address for receiving praises and complaints. Typically that of the author or copyright holder.

An example of this might be:

        copyright = {
            date  = "1992-2004";
            owner = "Bruce Korb";
            eaddr = 'bkorb@gnu.org';
            type  = GPL;
        };

‘detail’ This string is added to the usage output when the HELP option is selected.

‘explain’ Gives additional information whenever the usage routine is invoked..

‘package’ The name of the package the program belongs to. This will appear parenthetically after the program name in the version and usage output, e.g.: autogen

(GNU autogen) - The Automated Program Generator.
‘preserve-case’

This attribute will not change anything except appearance. Normally, the option names are all documented in lower case. However, if you specify this attribute, then they will display in the case used in their specification. Command line options will still be matched without case sensitivity.

‘prog-desc and’
‘opts-ptr’

These define global pointer variables that point to the program descriptor and the first option descriptor for a library option. This is intended for use by certain libraries that need command line and/or initialization file option processing. These definitions have no effect on the option template output, but are used for creating a library interface file. Normally, the first "option" for a library will be a documentation option that cannot be specified on the command line, but is marked as settable. The library client program will invoke the SET_OPTION macro which will invoke a handler function that will finally set these global variables.

‘usage’ Optionally names the usage procedure, if the library routine optionUsage() does not work for you. If you specify my_usage as the value of this attribute, for example, you will use a procedure by that name for displaying usage. Of course, you will need to provide that procedure and it must conform to this profile:

        void my_usage( tOptions* pOptions, int exitCode )
‘gnu-usage’

Normally, the default format produced by the optionUsage procedure is AutoOpts Standard. By specifying this attribute, the default format will be GNUish style. Either default may be overridden by the user with the AUTOOPTS_ USAGE environment variable. If it is set to gnu or autoopts, it will alter the style appropriately. This attribute will conflict with the usage attribute.

‘reorder-args’

Some applications traditionally require that the command operands be intermixed with the command options. In order to handle that, the arguments must be reordered. If you are writing such an application, specify this global option. All of the options (and any associated option arguments) will be brought to the beginning of the argument list. New applications should not use this feature, if at all possible. This feature is disabled if POSIXLY_CORRECT is defined in the environment.

7.5.5 Option Attributes

For each option you wish to specify, you must have a block macro named flag defined. There are two required attributes: name and descrip. If any options do not have a value (traditional flag character) attribute, then the long-opts program attribute must also be defined. As a special exception, if no options have a value and long-opts is not defined and argument is not defined, then all arguments to the program are named options. In this case, the - and -- command line option markers are optional.

7.5.5.1 Required Attributes

Every option must have exactly one copy of both of these attributes.

‘name’ Long name for the option. Even if you are not accepting long options and are only accepting flags, it must be provided. AutoOpts generates private, named storage that requires this name. This name also causes a #define-d name to be emitted. It must not conflict with any other names you may be using in your program. For example, if your option name is, debug or munged-up, you must not use the #define names DEBUG (or MUNGED_UP) in your program for non-AutoOpts related purposes. They are now used by AutoOpts. Sometimes (most especially under Windows), you may get a surprise. For example, INTERFACE is apparently a user space name that one should be free to use. Windows usurps this name. To solve this, you must do one of the following:

1. Change the name of your option

2. add the program attribute (see Section 7.5.1 [program attributes], page 73):

           export = '#undef INTERFACE';

3. add the program attribute:

           guard-option-names;

‘descrip’ Except for documentation options, a very brief description of the option. About 40 characters on one line, maximum. It appears on the usage() output next to the option name. If, however, the option is a documentation option, it will appear on one or more lines by itself. It is thus used to visually separate and comment upon groups of options in the usage text.

7.5.5.2 Common Option Attributes

These option attributes are optional. Any that do appear in the definition of a flag, may appear only once.

‘value’ The flag character to specify for traditional option flags, e.g., -L.

‘max’ Maximum occurrence count (invalid if disable present). The default maximum is 1. NOLIMIT can be used for the value, otherwise it must be a number or a #define that evaluates to a number.

‘min’ Minimum occurrence count. If present, then the option must appear on the command line. Do not define it with the value zero (0).

‘must-set’

If an option must be specified, but it need not be specified on the command line, then specify this attribute for the option.

‘enable’ Long-name prefix for enabling the option (invalid if disable not present). Only useful if long option names are being processed.

‘disable’ Prefix for disabling (inverting sense of) the option. Only useful if long option names are being processed.

‘enabled’ If default is for option being enabled. (Otherwise, the OPTST DISABLED bit is set at compile time.) Only useful if the option can be disabled.

‘ifdef’

‘ifndef’ If an option is relevant on certain platforms or when certain features are enabled or disabled, you can specify the compile time flag used to indicate when the option should be compiled in or out. For example, if you have a configurable feature, mumble that is indicated with the compile time define, WITH_MUMBLING, then add:

        ifdef = WITH_MUMBLING;

Take care when using these. There are several caveats:

• The case and spelling must match whatever is specified.

• Do not confuse these attributes with the AutoGen directives of the same names, See Section 2.5 [Directives], page 10. These cause C preprocessing directives to be inserted into the generated C text.

• Only one of these attributes may apply to any given option.

  • The VALUE_OPT_values are #define-d. If WITH_MUMBLING is not

defined, then the associated VALUE_OPT_value will not be #define-d either. So, if you have an option named, MUMBLING that is active

    only if WITH_MUMBLING is #define-d, then VALUE_OPT_MUMBLING will
    be #define-d iff WITH_MUMBLING is #define-d. Watch those switch

statements.

7.5.5.3 Special Option Handling

These option attributes do not fit well with other categories.

‘no-preset’

If presetting this option is not allowed. (Thus, environment variables and values set in configuration files will be ignored.)

‘settable’

If the option can be set outside of option processing. If this attribute is defined, special macros for setting this particular option will be inserted into the interface file. For example, TEMPL_DIRS is a settable option for AutoGen, so a macro named SET_OPT_TEMPL_DIRS(a) appears in the interface file. This attribute interacts with the documentation attribute.

‘equivalence’

Generally, when several options are mutually exclusive and basically serve the purpose of selecting one of several processing modes, these options can be considered an equivalence class. Sometimes, it is just easier to deal with them as such. All members of the equivalence class must contain the same equivalencedto option, including the equivalenced-to option itself. Thus, it must be a class member. For an option equivalence class, there is a single occurrence counter for the class. It can be referenced with the interface macro, COUNT_OPT(BASE_OPTION), where “BASE OPTION” is the equivalenced-to option name. Also, please take careful note: since the options are mapped to the equivalencedto option descriptor, any option argument values are mapped to that descriptor also. Be sure you know which “equivalent option” was selected before getting an option argument value! During the presetting phase of option processing (see Section 7.9 [Presetting Options], page 113), equivalenced options may be specified. However, if different equivalenced members are specified, only the last instance will be recognized and the others will be discarded. A conflict error is indicated only when multiple different members appear on the command line itself. As an example of where equivalenced options might be useful, cpio(1) has three options -o, -i, and -p that define the operational mode of the program (create, extract and pass-through, respectively). They form an equivalence class from which one and only one member must appear on the command line. If cpio were an AutoOpt-ed program, then each of these option definitions would contain:

        equivalence = create;

and the program would be able to determine the operating mode with code that worked something like this:

        switch (WHICH_IDX_CREATE) {
        case INDEX_OPT_CREATE:       ...
        case INDEX_OPT_EXTRACT:      ...
        case INDEX_OPT_PASS_THROUGH: ...
        default:    /* cannot happen */
        }
‘documentation’

This attribute means the option exists for the purpose of separating option description text in the usage output. Libraries may choose to make it settable so that the library can determine which command line option is the first one that pertains to the library. If present, this option disables all other attributes except settable, call-

proc and flag_-ode. settable must be and is only specified if call-proc,

extract-code or flag-code has been specified. When present, the descrip attribute will be displayed only when the --help option has been specified. It will be displayed flush to the left hand margin and may consist of one or more lines of text. The name of the option will not be printed. Documentation options are for clarifying the usage text and will not appear in generated man pages or in the generated invoking texinfo doc.

7.5.5.4 Immediate Action Attributes

Certain options may need to be processed early. For example, in order to suppress the processing of configuration files, it is necessary to process the command line option --no-load-opts before the config files are processed. To accommodate this, certain options may have their enabled or disabled forms marked for immediate processing. The consequence of this is that they are processed ahead of all other options in the reverse of normal order.

Normally, the first options processed are the options specified in the first homerc file, followed by then next homerc file through to the end of config file processing. Next, environment variables are processed and finally, the command line options. The later options override settings processed earlier. That actually gives them higher priority. Command line immediate action options actually have the lowest priority of all. They would be used only if they are to have an effect on the processing of subsequent options.

‘immediate’

Use this option attribute to specify that the enabled form of the option is to be processed immediately. The help and more-help options are so specified. They will also call exit() upon completion, so they do have an effect on the processing of the remaining options :-).

‘immed-disable’

Use this option attribute to specify that the disabled form of the option is to be processed immediately. The load-opts option is so specified. The --no-load-opts command line option will suppress the processing of config files and environment variables. Contrariwise, the --load-opts command line option is processed normally. That means that the options specified in that file will be processed after all the homerc files and, in fact, after options that precede it on the command line.

‘also’ If either the immediate or the immed-disable attributes are set to the string, “also”, then the option will actually be processed twice: first at the immediate processing phase and again at the “normal” time.

7.5.5.5 Option Conflict Attributes

These attributes may be used as many times as you need. They are used at the end of the option processing to verify that the context within which each option is found does not conflict with the presence or absence of other options.

This is not a complete cover of all possible conflicts and requirements, but it simple to implement and covers the more common situations.

‘flags-must’

one entry for every option that must be present when this option is present

‘flags-cant’

one entry for every option that cannot be present when this option is present

7.5.5.6 Option Argument Specification

Command line options come in three flavors: options that do not take arguments, those that do and those that may. Without an "arg-type" attribute, AutoOpts will not process an argument to an option. If "arg-type" is specified and "arg-optional" is also specified, then the next command line token will be taken to be an argument, unless it looks like the name of another option.

If the argument type is specified to be anything other than "str[ing]", then AutoOpts will specify a callback procedure to handle the argument. Some of these procedures will be created and inserted into the generated .c file, and others are already built into the ‘libopts’ library. Therefore, if you write your own callback procedure (see Section 7.5.5.7 [Option Argument Handling], page 89), then you must either not specify an "arg-type" attribute, or else specify it to be of type "str[ing]". Your callback function will be able to place its own restrictions on what that string may contain or represent.

‘arg-type’

This specifies the type of argument the option will take. If not present, the option cannot take an argument. If present, it must be one of the following five. The bracketed part of each name is optional.

          ‘str[ing]’

The argument may be any arbitrary string, though your program or option callback procedure may place additional constraints upon it.

          ‘num[ber]’

The argument must be a correctly formed integer, without any trailing U’s or L’s. AutoOpts contains a library procedure to convert the string to a number. If you specify range checking with arg-range, then AutoOpts produces a special purpose procedure for this option.

‘bool[ean]’

The argument will be interpreted and always yield either AG TRUE or AG FALSE. False values are the empty string, the number zero, or a string that starts with f, F, n or N (representing False or No). Anything else will be interpreted as True.

‘key[word]’

The argument must match a specified list of strings. Assuming you have named the option, optn-name, the strings will be converted into an enumeration of type te_Optn_Name with the values OPTN_ NAME_KEYWORD. If you have not specified a default value, the value OPTN_NAME_UNDEFINED will be inserted with the value zero. The option will be initialized to that value. You may now use this in your code as follows:

        te_Optn_Name opt = OPT_VALUE_OPTN_NAME;
        switch (opt) {
        case OPTN_NAME_UNDEFINED:  /* undefined things */ break;
        case OPTN_NAME_KEYWORD:    /* `keyword' things */ break;
        default: /* utterly impossible */ ;
        }

AutoOpts produces a special purpose procedure for this option. If you have need for the string name of the selected keyword, you may obtain this with the macro, OPT_OPTN_NAME_VAL2STR(val). The value you pass would normally be OPT_VALUE_OPTN_NAME, but anything with numeric value that is legal for te_Optn_Name may be passed. Anything out of range will result in the string, "*INVALID*" being returned. The strings are read only. It may be used as in:

        te_Optn_Name opt = OPT_VALUE_OPTN_NAME;
        printf( "you selected the %s keyword\n",
                OPT_OPTN_NAME_VAL2STR(opt) );
‘set[-membership]’

The argument must be a list of names each of which must match the strings "all", "none" or one of the keywords specified for this option. all will turn on all membership bits and none will turn them all off. Specifying one of the keywords will turn on the corresponding set membership bit. Literal numbers may also be used and may, thereby, set or clear more than one bit. Preceding a keyword or literal number with a bang (! - exclamation point) will turn the bit(s) off. The number of keywords allowed is constrained by the number of bits in a pointer, as the bit set is kept in a void*. If, for example, you specified first in your list of keywords, then you can use the following code to test to see if either first or all was specified:

                  uintptr_t opt = OPT_VALUE_OPTN_NAME;
                  if (opt & OPTN_NAME_FIRST)
                      /* OPTN_NAME_FIRST bit was set */ ;

AutoOpts produces a special purpose procedure for this option.

‘keyword’ If the arg-type is keyword or set-membership, then you must specify the

list of keywords by a series of keyword entries. The interface file will contain values for <OPTN_NAME>_<KEYWORD> for each keyword entry. keyword option types will have an enumeration and set-membership option types will have a set of unsigned long bits #define-d. If there are more than 32 bits defined, the #define will set unsigned long long values and you best be running on a 64 bit platform.

‘arg-optional’

This attribute indicates that the user does not have to supply an argument for the option. This is only valid if the arg-type is string or keyword. If it is keyword, then this attribute may also specify the default keyword to assume when the argument is not supplied. If left empty, arg-default or the zero-valued keyword will be used.

‘arg-default’

This specifies the default value to be used when the option is not specified or preset.

‘default’ If your program processes its arguments in named option mode (See “long-opts” in Section 7.5.1 [program attributes], page 73), then you may select one of your options to be the default option. Do so with this attribute. The option so specified must have an arg-type specified, but not the arg-optional attribute. That is to say, the option argument must be required. If you have done this, then any arguments that do not match an option name and do not contain an equal sign (=) will be interpreted as an option argument to the default option.

‘arg-range’

If the arg-type is number, then arg-ranges may be specified, too. If you specify one or more of these option attributes, then AutoOpts will create a callback procedure for handling it. The argument value supplied for the option must match one of the range entries. Each arg-range should consist of either an integer by itself or an integer range. The integer range is specified by one or two integers separated by the two character sequence, ->. Be sure to quote the entire range string. The definitions parser will not accept the range syntax as a single string token. The generated procedure imposes the range constraints as follows:

• A number by itself will match that one value.

• The high end of the range may not be INT_MIN, both for obvious reasons and because that value is used to indicate a single-valued match.

• An omitted lower value implies a lower bound of INT MIN.

• An omitted upper value implies a upper bound of INT MAX.

• The argument value is required. It may not be optional.

• The value must match one of the entries. If it can match more than one, then you have redundancies, but no harm will come of it.

7.5.5.7 Option Argument Handling

AutoOpts will either specify or automatically generate callback procedures for options that take specialized arguments. The only option argument types that are not specialized are plain string arguments and no argument at all. For options that fall into one of those two categories, you may specify your own callback function, as specified below. If the option takes a string argument, then you may specify that the option is to be handled by the libopts library procedures stackOptArg() or unstackOptArg() (see below). Finally, documentation options (Section 7.5.5.3 [Special Option Handling], page 84) may also be marked as settable and have special callback functions (either flag-code, extract-code,

or call-proc).
‘flag-code’

statements to execute when the option is encountered. The generated procedure will look like this:

        static void
        doOpt<name>( tOptions* pOptions, tOptDesc* pOptDesc )
        {
        <flag_code>
        }

Only certain fields within the tOptions and tOptDesc structures may be accessed. See Section 7.6.1 [Option Processing Data], page 94.

‘extract-code’

This is effectively identical to flag_code, except that the source is kept in the output file instead of the definitions file. A long comment is used to demarcate the code. You must not modify that marker. Before regenerating the option code file, the old file is renamed from MUMBLE.c to MUMBLE.c.save. The template will be looking there for the text to copy into the new output file.

‘call-proc’

external procedure to call when option is encountered. The calling sequence must conform to the sequence defined above for the generated procedure, doOpt<name>. It has the same restrictions regarding the fields within the structures passed in as arguments. See Section 7.6.1 [Option Processing Data], page 94.

‘flag-proc’

Name of another option whose flag_code can be executed when this option is encountered.

‘stack-arg’

Call a special library routine to stack the option’s arguments. Special macros in the interface file are provided for determining how many of the options were found (STACKCT_OPT(NAME)) and to obtain a pointer to a list of pointers to the argument values (STACKLST_OPT(NAME)). Obviously, for a stackable argument, the max attribute needs to be set higher than 1. If this stacked argument option has a disablement prefix, then the entire stack of arguments will be cleared by specifying the option with that disablement prefix.

‘unstack-arg’

Call a special library routine to remove (“unstack”) strings from a stack-arg option stack. This attribute must name the option that is to be “unstacked”. Neither this option nor the stacked argument option it references may be equivalenced to another option.

7.5.6 Man and Info doc Attributes

AutoOpts includes AutoGen templates for producing abbreviated man pages and for producing the invoking section of an info document. To take advantage of these templates, you must add several attributes to your option definitions.

‘doc’ First, every flag definition other than “documentation” definitions, must have a doc attribute defined. If the option takes an argument, then it will need an arg-name attribute as well. The doc text should be in plain sentences with minimal formatting. The Texinfo commands @code, and @var will have its enclosed text made into \fB entries in the man page, and the @file text will be made into \fI entries. The arg-name attribute is used to display the option’s argument in the man page. Options marked with the “documentation” attribute are for documenting the usage text. All other options should have the “doc” attribute in order to document the usage of the option in the generated man pages.

‘arg-name’

If an option has an argument, the argument should have a name for documentation purposes. It will default to arg-type, but it will likely be clearer with something else like, file-name instead of string (the type).

‘prog-man-descrip’
‘prog-info-descrip’

Then, you need to supply a brief description of what your program does. If you already have a detail definition, this may be sufficient. If not, or if you need special formatting for one of the manual formats, then you will need ei-

ther a definition for prog-man-descrip or prog-info-descrip or both. These

will be inserted verbatim in the man page document and the info document, respectively.

‘man-doc’ Finally, if you need to add man page sections like SEE ALSO or USAGE or other, put that text in a man-doc definition. This text will be inserted verbatim in the man page after the OPTIONS section and before the AUTHOR section.

7.5.7 Automatically Supported Options

AutoOpts provides automated support for five options. help and more-help are always provided. version is provided if version is defined in the option definitions See Section 7.5.1 [program attributes], page 73. save-opts and load-opts are provided if at least one homerc is defined See Section 7.5.1 [program attributes], page 73.

Below are the option names and flag values. The flags are activated if and only if at least one user-defined option also uses a flag value. These flags may be deleted or changed to characters of your choosing by specifying xxx-value = "y";, where xxx is one of the five names below and y is either empty or the character of your choice. For example, to change the help flag from ? to h, specify help-value = "h";; and to require that save-opts be specified only with its long option name, specify save-opts-value = "";.

‘help -?’ This option will immediately invoke the USAGE() procedure and display the usage line, a description of each option with its description and option usage information. This is followed by the contents of the definition of the detail text macro.

‘more-help -!’

This option is identical to the help option, except that the output is passed through a pager program. (more by default, or the program identified by the PAGER environment variable.)

‘version -v’

This will print the program name, title and version. If it is followed by the letter c and a value for copyright and owner have been provided, then the copyright will be printed, too. If it is followed by the letter n, then the full copyright notice (if available) will be printed.

‘save-opts ->’

This option will cause the option state to be printed in the configuration file format when option processing is done but not yet verified for consistency. The program will terminate successfully without running when this has completed. Note that for most shells you will have to quote or escape the flag character to restrict special meanings to the shell. The output file will be the configuration file name (default or provided by rcfile) in the last directory named in a homerc definition. This option may be set from within your program by invoking the "SET_OPT_ SAVE_OPTS(filename)" macro (see Section 7.6.16 [SET OPT name], page 97). Invoking this macro will set the file name for saving the option processing state, but the state will not actually be saved. You must call optionSaveFile to do that (see Section 7.6.28.13 [libopts-optionSaveFile], page 107). CAVEAT: if, after invoking this macro, you call optionProcess, the option processing state will be saved to this file and optionProcess will not return. You may wish to invoke CLEAR_OPT( SAVE_OPTS ) (see Section 7.6.2 [CLEAR OPT], page 96) beforehand.

‘load-opts -<’

This option will load options from the named file. They will be treated exactly as if they were loaded from the normally found configuration files, but will not be loaded until the option is actually processed. This can also be used within another configuration file, causing them to nest.

It is ultimately intended that specifying the option, no-load-opts will suppress the processing of configuration files and environment variables. To do this, AutoOpts must first implement pre-scanning of the options, environment and config files.

7.5.8 Library of Standard Options

AutoOpts has developed a set of standardized options. You may incorporate these options in your program simply by first adding a #define for the options you want, and then the line,

#include stdoptions.def

in your option definitions. The supported options are specified thus:

#define DEBUG
#define DIRECTORY
#define DRY_RUN
#define INPUT
#define INTERACTIVE
#define OUTPUT
#define WARN
#define SILENT
#define QUIET
#define BRIEF
#define VERBOSE

By default, only the long form of the option will be available. To specify the short (flag) form, suffix these names with _FLAG. e.g.,

#define DEBUG_FLAG

--silent, --quiet, --brief and --verbose are related in that they all indicate some level of diagnostic output. These options are all designed to conflict with each other. Instead of four different options, however, several levels can be incorporated by #defineing VERBOSE_ENUM. In conjunction with VERBOSE, it incorporates the notion of 5 levels in an enumeration: silent, quiet, brief, informative and verbose; with the default being

brief.

Here is an example program that uses the following set of definitions:

AutoGen Definitions options;
prog-name  = default-test;
prog-title = 'Default Option Example';
homerc     = '$$/../share/default-test', '$HOME', '.';
environrc;
long-opts;
gnu-usage;
version    = '1.0';
main = {
  main-type = shell-process;
};

#define DEBUG_FLAG
#define WARN_FLAG
#define WARN_LEVEL
#define VERBOSE_FLAG
#define VERBOSE_ENUM
#define DRY_RUN_FLAG
#define OUTPUT_FLAG
#define INPUT_FLAG
#define DIRECTORY_FLAG
#define INTERACTIVE_FLAG
#include stdoptions.def

Running a few simple commands on that definition file:

autogen default-test.def
copts="-DTEST_DEFAULT_TEST_OPTS `autoopts-config cflags`"
lopts="`autoopts-config ldflags`"
cc -o default-test ${copts} default-test.c ${lopts}

Yields a program which, when run with ‘--help’, prints out:

default-test - Default Option Example - Ver. 1.0
USAGE:  default-test [ -<flag> [<val>] | --<name>[{=| }<val>] ]...

The following options are commonly used and are provided and supported
by AutoOpts:
   -D, --debug                run program with debugging info
   -V, --verbose=KWd          run program with progress info
   -w, --warn=num             specify a warning-level threshhold
                                - disabled as --no-warn
   -d, --dry-run              program will make no changes
   -I, --interactive=str      prompt for confirmation
   -i, --input=str            redirect input from file
   -o, --output=str           redirect output to file
   -d, --directory=str        use specified dir for I/O
version and help options:
   -v, --version[=arg]        Output version information and exit
   -?, --help                 Display usage information and exit
   -!, --more-help            Extended usage information passed thru pager
   ->, --save-opts[=arg]      Save the option state to a config file
   -<, --load-opts=str        Load options from a config file
                                - disabled as --no-load-opts
                                - may appear multiple times
Options are specified by doubled hyphens and their name
or by a single hyphen and the flag character.

        The following option preset mechanisms are supported:
         - reading file $$/../share/default-test
         - reading file $HOME
         - reading file /home/bkorb/ag/ag/doc/.default_testrc
         - examining environment variables named DEFAULT_TEST_*
        The valid "verbose" option keywords are:
          silent quiet brief informative verbose

7.6 Programmatic Interface

The user interface for access to the argument information is completely defined in the generated header file and in the portions of the distributed file "options.h" that are marked "public".

In the following macros, text marked <NAME> or name is the name of the option in upper case and segmented with underscores _. The macros and enumerations defined in the options header (interface) file are used as follows:

To see how these #define macros are used in a program, the reader is referred to the several ‘opts.h’ files included with the AutoGen sources.

7.6.1 Data for Option Processing

This section describes the data that may be accessed from within the option processing callback routines. The following fields may be used in the following ways and may be used for read only. The first set is addressed from the tOptDesc* pointer:

‘optIndex’
‘optValue’

These may be used by option procedures to determine which option they are working on (in case they handle several options).

‘optActualIndex’
‘optActualValue’

These may be used by option procedures to determine which option was used to set the current option. This may be different from the above if the options are members of an equivalence class.

‘optOccCt’

If AutoOpts is processing command line arguments, then this value will contain the current occurrence count. During the option preset phase (reading configuration files and examining environment variables), the value is zero.

‘fOptState’

The field may be tested for the following bit values (prefix each name with

          OPTST_, e.g. OPTST_INIT):

‘INIT’ Initial compiled value. As a bit test, it will always yield FALSE.

‘SET’ The option was set via the SET_OPT() macro.

‘PRESET’ The option was set via a configuration file.

‘DEFINED’ The option was set via a command line option.

‘SET_MASK’

This is a mask of flags that show the set state, one of the above four values.

‘EQUIVALENCE’

This bit is set when the option was selected by an equivalenced option.

‘DISABLED’

This bit is set if the option is to be disabled. (Meaning it was a long option prefixed by the disablement prefix, or the option has not been specified yet and initializes as disabled.)

As an example of how this might be used, in AutoGen I want to allow template writers to specify that the template output can be left in a writable or readonly state. To support this, there is a Guile function named set-writable (see Section 3.4.36 [SCM set-writable], page 29). Also, I provide for command options --writable and --not-writable. I give precedence to command line and RC file options, thus:

        switch (STATE_OPT( WRITABLE )) {
        case OPTST_DEFINED:
        case OPTST_PRESET:
            fprintf( stderr, zOverrideWarn, pCurTemplate->pzFileName,
                     pCurMacro->lineNo );
            break;
        default:
            if (gh_boolean_p( set ) && (set == SCM_BOOL_F))
                CLEAR_OPT( WRITABLE );
            else
                SET_OPT_WRITABLE;
        }
‘pzLastArg’

Pointer to the latest argument string. BEWARE If the argument type is numeric, an enumeration or a bit mask, then this will be the argument value and not a pointer to a string.

The following two fields are addressed from the tOptions* pointer:

‘pzProgName’

Points to a NUL-terminated string containing the current program name, as retrieved from the argument vector.

‘pzProgPath’

Points to a NUL-terminated string containing the full path of the current program, as retrieved from the argument vector. (If available on your system.)

Note these fields get filled in during the first call to optionProcess(). All other fields are private, for the exclusive use of AutoOpts code and are subject to change.

7.6.2 CLEAR OPT( <NAME> ) - Clear Option Markings

Make as if the option had never been specified. HAVE_OPT(<NAME>) will yield FALSE after invoking this macro.

7.6.3 COUNT OPT( <NAME> ) - Definition Count

This macro will tell you how many times the option was specified on the command line. It does not include counts of preset options.

        if (COUNT_OPT( NAME ) != desired-count) {
            make-an-undesirable-message.
        }

7.6.4 DESC( <NAME> ) - Option Descriptor

This macro is used internally by other AutoOpt macros. It is not for general use. It is used to obtain the option description corresponding to its UPPER CASED option name argument. This is primarily used in other macro definitions.

7.6.5 DISABLE OPT name - Disable an option

This macro is emitted if it is both settable and it can be disabled. If it cannot be disabled, it may always be CLEAR-ed (see above).

The form of the macro will actually depend on whether the option is equivalenced to another, and/or has an assigned handler procedure. Unlike the SET_OPT macro, this macro does not allow an option argument.

        DISABLE_OPT_NAME;

7.6.6 ENABLED OPT( <NAME> ) - Is Option Enabled?

Yields true if the option defaults to disabled and ISUNUSED_OPT() would yield true. It also yields true if the option has been specified with a disablement prefix, disablement value or the DISABLE_OPT_NAME macro was invoked.

7.6.7 ERRSKIP OPTERR - Ignore Option Errors

When it is necessary to continue (return to caller) on option errors, invoke this option. It is reversible. See Section 7.6.8 [ERRSTOP OPTERR], page 96.

7.6.8 ERRSTOP OPTERR - Stop on Errors

After invoking this macro, if optionProcess() encounters an error, it will call exit(1) rather than return. This is the default processing mode. It can be overridden by specifying allow-errors in the definitions file, or invoking the macro See Section 7.6.7 [ERRSKIP OPTERR], page 96.

7.6.9 HAVE OPT( <NAME> ) - Have this option?

This macro yields true if the option has been specified in any fashion at all. It is used thus:

        if (HAVE_OPT( NAME )) {
            <do-things-associated-with-opt-name>;
        }

7.6.10 ISSEL OPT( <NAME> ) - Is Option Selected?

This macro yields true if the option has been specified either on the command line or via a SET/DISABLE macro.

7.6.11 ISUNUSED OPT( <NAME> ) - Never Specified?

This macro yields true if the option has never been specified, or has been cleared via the

CLEAR_OPT() macro.

7.6.12 OPTION CT - Full Count of Options

The full count of all options, both those defined and those generated automatically by AutoOpts. This is primarily used to initialize the program option descriptor structure.

7.6.13 OPT ARG( <NAME> ) - Option Argument String

The option argument value as a pointer to string. Note that argument values that have been specified as numbers are stored as numbers or keywords. For such options, use instead the OPT_VALUE_name define. It is used thus:

        if (HAVE_OPT( NAME )) {
            char* p = OPT_ARG( NAME );
            <do-things-with-opt-name-argument-string>;
        }

7.6.14 OPT VALUE name - Option Argument Value

This macro gets emitted only for options that take numeric, keyword or set membership arguments. The macro yields a word-sized integer containing the enumeration or numeric value of the option argument.

        int opt_val = OPT_VALUE_NAME;

7.6.15 RESTART OPT( n ) - Resume Option Processing

If option processing has stopped (either because of an error or something was encountered that looked like a program argument), it can be resumed by providing this macro with the index n of the next option to process and calling optionProcess() again.

7.6.16 SET OPT name - Force an option to be set

This macro gets emitted only when the given option has the settable attribute specified.

The form of the macro will actually depend on whether the option is equivalenced to another, has an option argument and/or has an assigned handler procedure. If the option has an argument, then this macro will too. Beware that the argument is not reallocated, so the value must not be on the stack or deallocated in any other way for as long as the value might get referenced.

If you have supplied at least one ‘homerc’ file (see Section 7.5.1 [program attributes], page 73), this macro will be emitted for the --save-opts option.

        SET_OPT_SAVE_OPTS( "filename" );

See Section 7.5.7 [automatic options], page 90, for a discussion of the implications of using this particular example.

7.6.17 STACKCT OPT( <NAME> ) - Stacked Arg Count

When the option handling attribute is specified as stack_arg, this macro may be used to determine how many of them actually got stacked.

Do not use this on options that have not been stacked or has not been specified (the stack_arg attribute must have been specified, and HAVE_OPT(<NAME>) must yield TRUE). Otherwise, you will likely seg fault.

        if (HAVE_OPT( NAME )) {
            int     ct = STACKCT_OPT(  NAME );
            char**  pp = STACKLST_OPT( NAME );
            do  {
                char* p = *pp++;
                do-things-with-p;
            } while (--ct > 0);
        }

7.6.18 STACKLST OPT( <NAME> ) - Argument Stack

The address of the list of pointers to the option arguments. The pointers are ordered by the order in which they were encountered in the option presets and command line processing.

Do not use this on options that have not been stacked or has not been specified (the stack_arg attribute must have been specified, and HAVE_OPT(<OPTION>) must yield TRUE). Otherwise, you will likely seg fault.

        if (HAVE_OPT( NAME )) {
            int     ct = STACKCT_OPT(  NAME );
            char**  pp = STACKLST_OPT( NAME );
            do  {
                char* p = *pp++;
                do-things-with-p;
            } while (--ct > 0);
        }

7.6.19 START OPT - Restart Option Processing

This is just a shortcut for RESTART OPT(1) (See Section 7.6.15 [RESTART OPT], page 97.)

7.6.20 STATE OPT( <NAME> ) - Option State

If you need to know if an option was set because of presetting actions (configuration file processing or environment variables), versus a command line entry versus one of the SET/DISABLE macros, then use this macro. It will yield one of four values: OPTST_INIT,

OPTST_SET, OPTST_PRESET or OPTST_DEFINED. It is used thus:
        switch (STATE_OPT( NAME )) {
            case OPTST_INIT:
                not-preset, set or on the command line.  (unless CLEAR-ed)

            case OPTST_SET:
                option set via the SET_OPT_NAME() macro.
            case OPTST_PRESET:
                option set via an configuration file or environment variable
            case OPTST_DEFINED:
                option set via a command line option.
            default:
                cannot happen :)
        }

7.6.21 USAGE( exit-code ) - Usage invocation macro

This macro invokes the procedure registered to display the usage text. Normally, this will be optionUsage from the AutoOpts library, but you may select another procedure by specifying usage = "proc_name" program attribute. This procedure must take two arguments first, a pointer to the option descriptor, and second the exit code. The macro supplies the option descriptor automatically. This routine is expected to call exit(3) with the provided exit code.

The optionUsage routine also behaves differently depending on the exit code. If the exit code is zero, it is assumed that assistance has been requested. Consequently, a little more information is provided than when displaying usage and exiting with a non-zero exit code.

7.6.22 VALUE OPT name - Option Flag Value

This is a #define for the flag character used to specify an option on the command line. If value was not specified for the option, then it is a unique number associated with the option. option value refers to this value, option argument refers to the (optional) argument to the option.

        switch (WHICH_OPT_OTHER_OPT) {
        case VALUE_OPT_NAME:
            this-option-was-really-opt-name;
        case VALUE_OPT_OTHER_OPT:
            this-option-was-really-other-opt;
        }

7.6.23 VERSION - Version and Full Version

If the version attribute is defined for the program, then a stringified version will be #defined as PROGRAM VERSION and PROGRAM FULL VERSION. PRO-GRAM FULL VERSION is used for printing the program version in response to the version option. The version option is automatically supplied in response to this attribute, too.

You may access PROGRAM VERSION via programOptions.pzFullVersion.

7.6.24 WHICH IDX name - Which Equivalenced Index

This macro gets emitted only for equivalenced-to options. It is used to obtain the index for the one of the several equivalence class members set the equivalenced-to option.

        switch (WHICH_IDX_OTHER_OPT) {
        case INDEX_OPT_NAME:
            this-option-was-really-opt-name;
        case INDEX_OPT_OTHER_OPT:
            this-option-was-really-other-opt;
        }

7.6.25 WHICH OPT name - Which Equivalenced Option

This macro gets emitted only for equivalenced-to options. It is used to obtain the value code for the one of the several equivalence class members set the equivalenced-to option.

        switch (WHICH_OPT_OTHER_OPT) {
        case VALUE_OPT_NAME:
            this-option-was-really-opt-name;
        case VALUE_OPT_OTHER_OPT:
            this-option-was-really-other-opt;
        }

7.6.26 teOptIndex - Option Index and Enumeration

This enum defines the complete set of options, both user specified and automatically provided. This can be used, for example, to distinguish which of the equivalenced options was actually used.

        switch (pOptDesc->optActualIndex) {
        case INDEX_OPT_FIRST:
            stuff;
        case INDEX_OPT_DIFFERENT:
            different-stuff;
        default:
            unknown-things;
        }

7.6.27 OPTIONS STRUCT VERSION - active version

You will not actually need to reference this value, but you need to be aware that it is there. It is the first value in the option descriptor that you pass to optionProcess. It contains a magic number and version information. Normally, you should be able to work with a more recent option library than the one you compiled with. However, if the library is changed incompatibly, then the library will detect the out of date magic marker, explain the difficulty and exit. You will then need to rebuild and recompile your option definitions. This has rarely been necessary.

7.6.28 libopts External Procedures

These are the routines that libopts users may call directly from their code. There are several other routines that can be called by code generated by the libopts option templates, but they are not to be called from any other user code. The ‘options.h’ header is fairly clear about this, too.

This subsection was automatically generated by AutoGen using extracted information and the aginfo3.tpl template.

7.6.28.1 ao string tokenize

tokenize an input string

Usage:

        token_list_t* res = ao_string_tokenize( string );

Where the arguments are:

Name  Type  Description
—–  —–  ————-
string  char const*  string to be tokenized
returns  token list t*  pointer to a structure that lists each token

This function will convert one input string into a list of strings. The list of strings is derived by separating the input based on white space separation. However, if the input contains either single or double quote characters, then the text after that character up to a matching quote will become the string in the list.

The returned pointer should be deallocated with free(3C) when are done using the data. The data are placed in a single block of allocated memory. Do not deallocate individual token/strings.

The structure pointed to will contain at least these two fields:

‘tkn_ct’ The number of tokens found in the input string.

‘tok_list’

An array of tkn_ct + 1 pointers to substring tokens, with the last pointer set to NULL.

There are two types of quoted strings: single quoted (') and double quoted ("). Singly quoted strings are fairly raw in that escape characters (\\) are simply another character, except when preceding the following characters:

        \\  double backslashes reduce to one
        '   incorporates the single quote into the string
        \n  suppresses both the backslash and newline character

Double quote strings are formed according to the rules of string constants in ANSI-C programs.

NULL is returned and errno will be set to indicate the problem:

• EINVAL - There was an unterminated quoted string.

• ENOENT - The input string was empty.

• ENOMEM - There is not enough memory.

7.6.28.2 configFileLoad

parse a configuration file Usage:

        const tOptionValue* res = configFileLoad( pzFile );

Where the arguments are:

Name  Type  Description
—–  —–  ————-
pzFile  char const*  the file to load

returns const tOptionValue*

An allocated, compound value structure

This routine will load a named configuration file and parse the text as a hierarchically valued option. The option descriptor created from an option definition file is not used via this interface. The returned value is "named" with the input file name and is of type "OPARG_

TYPE_HIERARCHY". It may be used in calls to optionGetValue(), optionNextValue() and
optionUnloadNested().

If the file cannot be loaded or processed, NULL is returned and errno is set. It may be set by a call to either open(2) mmap(2) or other file system calls, or it may be:

• ENOENT - the file was empty.

• EINVAL - the file contents are invalid – not properly formed.

• ENOMEM - not enough memory to allocate the needed structures.

7.6.28.3 optionFileLoad

Load the locatable config files, in order Usage:

        int res = optionFileLoad( pOpts, pzProg );

Where the arguments are:

Name  Type  Description
—–  —–  ————-
pOpts  tOptions*  program options descriptor
pzProg  char const*  program name
returns  int  0 -> SUCCESS, -1 -> FAILURE

This function looks in all the specified directories for a configuration file ("rc" file or "ini" file) and processes any found twice. The first time through, they are processed in reverse order (last file first). At that time, only "immediate action" configurables are processed. For example, if the last named file specifies not processing any more configuration files, then no more configuration files will be processed. Such an option in the first named directory will have no effect.

Once the immediate action configurables have been handled, then the directories are handled in normal, forward order. In that way, later config files can override the settings of earlier config files.

See the AutoOpts documentation for a thorough discussion of the config file format. Configuration files not found or not decipherable are simply ignored.

Returns the value, "-1" if the program options descriptor is out of date or indecipherable. Otherwise, the value "0" will always be returned.

7.6.28.4 optionFindNextValue

find a hierarcicaly valued option instance Usage:

  const tOptionValue* res = optionFindNextValue( pOptDesc, pPrevVal, name, value );

Where the arguments are:

Name  Type  Description
—–  —–  ————-
pOptDesc const tOptDesc*  an option with a nested arg type

pPrevVal

           const
           tOptionValue*

the last entry

name  char const*  name of value to find
value  char const*  the matching value

returns const tOptionValue*

a compound value structure

This routine will find the next entry in a nested value option or configurable. It will search through the list and return the next entry that matches the criteria.

The returned result is NULL and errno is set:

• EINVAL - the pOptValue does not point to a valid hierarchical option value.

• ENOENT - no entry matched the given name.

7.6.28.5 optionFindValue

find a hierarcicaly valued option instance Usage:

  const tOptionValue* res = optionFindValue( pOptDesc, name, value );

Where the arguments are:

Name  Type  Description
—–  —–  ————-
pOptDesc const tOptDesc*  an option with a nested arg type
name  char const*  name of value to find
value  char const*  the matching value

returns const tOptionValue*

a compound value structure

This routine will find an entry in a nested value option or configurable. It will search through the list and return a matching entry.

The returned result is NULL and errno is set:

• EINVAL - the pOptValue does not point to a valid hierarchical option value.

• ENOENT - no entry matched the given name.

7.6.28.6 optionFree

free allocated option processing memory Usage:

        optionFree( pOpts );

Where the arguments are:

Name  Type  Description
—–  —–  ————-
pOpts  tOptions*  program options descriptor

AutoOpts sometimes allocates memory and puts pointers to it in the option state structures. This routine deallocates all such memory.

As long as memory has not been corrupted, this routine is always successful.

7.6.28.7 optionGetValue

get a specific value from a hierarcical list Usage:

        const tOptionValue* res = optionGetValue( pOptValue, valueName );

Where the arguments are:

Name  Type  Description
—–  —–  ————-

pOptValue

                 const
                 tOptionValue*

a hierarchcal value

valueName char const*  name of value to get

returns const tOptionValue*

a compound value structure

This routine will find an entry in a nested value option or configurable. If "valueName" is NULL, then the first entry is returned. Otherwise, the first entry with a name that exactly matches the argument will be returned.

The returned result is NULL and errno is set:

• EINVAL - the pOptValue does not point to a valid hierarchical option value.

• ENOENT - no entry matched the given name.

7.6.28.8 optionLoadLine

process a string for an option name and value Usage:

        optionLoadLine( pOpts, pzLine );

Where the arguments are:

Name  Type  Description
—–  —–  ————-
pOpts  tOptions*  program options descriptor
pzLine  char const*  NUL-terminated text

This is a client program callable routine for setting options from, for example, the contents of a file that they read in. Only one option may appear in the text. It will be treated as a normal (non-preset) option.

When passed a pointer to the option struct and a string, it will find the option named by the first token on the string and set the option argument to the remainder of the string. The caller must NUL terminate the string. Any embedded new lines will be included in the option argument. If the input looks like one or more quoted strings, then the input will be "cooked". The "cooking" is identical to the string formation used in AutoGen definition files (see Section 3.3.2 [basic expression], page 20), except that you may not use backquotes.

Invalid options are silently ignored. Invalid option arguments will cause a warning to print, but the function should return.

7.6.28.9 optionNextValue

get the next value from a hierarchical list Usage:

        const tOptionValue* res = optionNextValue( pOptValue, pOldValue );

Where the arguments are:

Name  Type  Description
—–  —–  ————-

pOptValue

                 const
                 tOptionValue*

a hierarchcal list value

pOldValue

                 const
                 tOptionValue*

a value from this list

returns const tOptionValue*

a compound value structure

This routine will return the next entry after the entry passed in. At the end of the list, NULL will be returned. If the entry is not found on the list, NULL will be returned and "errno" will be set to EINVAL. The "pOldValue" must have been gotten from a prior call

to this routine or to "opitonGetValue()".

The returned result is NULL and errno is set:

• EINVAL - the pOptValue does not point to a valid hierarchical option value or pOldValue does not point to a member of that option value.

• ENOENT - the supplied pOldValue pointed to the last entry.

7.6.28.10 optionOnlyUsage

Print usage text for just the options Usage:

        optionOnlyUsage( pOpts, ex_code );

Where the arguments are:

Name  Type  Description
—–  —–  ————-
pOpts  tOptions*  program options descriptor

ex code  int  exit code for calling exit(3)

This routine will print only the usage for each option. This function may be used when the emitted usage must incorporate information not available to AutoOpts.

7.6.28.11 optionProcess

this is the main option processing routine Usage:

        int res = optionProcess( pOpts, argc, argv );

Where the arguments are:

Name  Type  Description
—–  —–  ————-
pOpts  tOptions*  program options descriptor
argc  int  program arg count
argv  char**  program arg vector
returns  int  the count of the arguments processed

This is the main entry point for processing options. It is intended that this procedure be called once at the beginning of the execution of a program. Depending on options selected earlier, it is sometimes necessary to stop and restart option processing, or to select completely different sets of options. This can be done easily, but you generally do not want to do this.

The number of arguments processed always includes the program name. If one of the arguments is "–", then it is counted and the processing stops. If an error was encountered and errors are to be tolerated, then the returned value is the index of the argument causing the error. A hyphen by itself ("-") will also cause processing to stop and will not be counted among the processed arguments. A hyphen by itself is treated as an operand. Encountering an operand stops option processing.

Errors will cause diagnostics to be printed. exit(3) may or may not be called. It depends upon whether or not the options were generated with the "allow-errors" attribute, or if the ERRSKIP OPTERR or ERRSTOP OPTERR macros were invoked.

7.6.28.12 optionRestore

restore option state from memory copy Usage:

        optionRestore( pOpts );

Where the arguments are:

Name  Type  Description
—–  —–  ————-
pOpts  tOptions*  program options descriptor

Copy back the option state from saved memory. The allocated memory is left intact, so this routine can be called repeatedly without having to call optionSaveState again. If you are restoring a state that was saved before the first call to optionProcess(3AO), then you may change the contents of the argc/argv parameters to optionProcess.

If you have not called optionSaveState before, a diagnostic is printed to stderr and exit is called.

7.6.28.13 optionSaveFile

saves the option state to a file Usage:

        optionSaveFile( pOpts );

Where the arguments are:

Name  Type  Description
—–  —–  ————-
pOpts  tOptions*  program options descriptor

This routine will save the state of option processing to a file. The name of that file can be specified with the argument to the --save-opts option, or by appending the rcfile attribute to the last homerc attribute. If no rcfile attribute was specified, it will default to .programnamerc. If you wish to specify another file, you should invoke the SET_OPT_

SAVE_OPTS( filename ) macro.

If no homerc file was specified, this routine will silently return and do nothing. If the output file cannot be created or updated, a message will be printed to stderr and the routine will return.

7.6.28.14 optionSaveState

saves the option state to memory Usage:

        optionSaveState( pOpts );

Where the arguments are:

Name  Type  Description
—–  —–  ————-
pOpts  tOptions*  program options descriptor

This routine will allocate enough memory to save the current option processing state. If this routine has been called before, that memory will be reused. You may only save one copy of the option state. This routine may be called before optionProcess(3AO). If you do call it before the first call to optionProcess, then you may also change the contents of argc/argv after you call optionRestore(3AO)

If it fails to allocate the memory, it will print a message to stderr and exit. Otherwise, it will always succeed.

7.6.28.15 optionUnloadNested

Deallocate the memory for a nested value Usage:

        optionUnloadNested( pOptVal );

Where the arguments are:

Name  Type  Description
—–  —–  ————-

pOptVal

                 const
                 tOptionValue*

the hierarchical value

A nested value needs to be deallocated. The pointer passed in should have been gotten from a call to configFileLoad() (See see Section 7.6.28.2 [libopts-configFileLoad], page 102).

7.6.28.16 optionVersion

return the compiled AutoOpts version number Usage:

        char const* res = optionVersion();

Where the arguments are:

Name  Type  Description
—–  —–  ————-
returns  char const*  the version string in constant memory

Returns the full version string compiled into the library. The returned string cannot be modified.

7.6.28.17 pathfind

fild a file in a list of directories Usage:

        char* res = pathfind( path, file, mode );

Where the arguments are:

Name  Type  Description
—–  —–  ————-
path  char const*  colon separated list of search directories
file  char const*  the name of the file to look for
mode  char const*  the mode bits that must be set to match
returns  char*  the path to the located file

pathfind looks for a a file with name "FILE" and "MODE" access along colon delimited "PATH", and returns the full pathname as a string, or NULL if not found. If "FILE" contains a slash, then it is treated as a relative or absolute path and "PATH" is ignored.

NOTE: this function is compiled into ‘libopts’ only if it is not natively supplied.

The "MODE" argument is a string of option letters chosen from the list below:

        Letter    Meaning
        r         readable
        w         writable
        x         executable
        f         normal file       (NOT IMPLEMENTED)
        b         block special     (NOT IMPLEMENTED)
        c         character special (NOT IMPLEMENTED)
        d         directory         (NOT IMPLEMENTED)

        p         FIFO (pipe)       (NOT IMPLEMENTED)
        u         set user ID bit   (NOT IMPLEMENTED)
        g         set group ID bit  (NOT IMPLEMENTED)
        k         sticky bit        (NOT IMPLEMENTED)
        s         size nonzero      (NOT IMPLEMENTED)

returns NULL if the file is not found.

7.6.28.18 strequate

map a list of characters to the same value Usage:

        strequate( ch_list );

Where the arguments are:

Name  Type  Description
—–  —–  ————-
ch list  char const*  characters to equivalence

Each character in the input string get mapped to the first character in the string. This function name is mapped to option strequate so as to not conflict with the POSIX name space.

none.

7.6.28.19 streqvcmp

compare two strings with an equivalence mapping Usage:

        int res = streqvcmp( str1, str2 );

Where the arguments are:

Name  Type  Description
—–  —–  ————-
str1  char const*  first string
str2  char const*  second string
returns  int  the difference between two differing characters

Using a character mapping, two strings are compared for "equivalence". Each input character is mapped to a comparison character and the mapped-to characters are compared for the two NUL terminated input strings. This function name is mapped to option streqvcmp so as to not conflict with the POSIX name space.

none checked. Caller responsible for seg faults.

7.6.28.20 streqvmap

Set the character mappings for the streqv functions Usage:

        streqvmap( From, To, ct );

Where the arguments are:

Name  Type  Description

—–  —–  ————-
From  char  Input character
To  char  Mapped-to character
ct  int  compare length

Set the character mapping. If the count (ct) is set to zero, then the map is cleared by setting all entries in the map to their index value. Otherwise, the "From" character is mapped to the "To" character. If ct is greater than 1, then From and To are incremented and the process repeated until ct entries have been set. For example,

        streqvmap( 'a', 'A', 26 );

will alter the mapping so that all English lower case letters will map to upper case.

This function name is mapped to option streqvmap so as to not conflict with the POSIX name space.

none.

7.6.28.21 strneqvcmp

compare two strings with an equivalence mapping Usage:

        int res = strneqvcmp( str1, str2, ct );

Where the arguments are:

Name  Type  Description
—–  —–  ————-
str1  char const*  first string
str2  char const*  second string
ct  int  compare length
returns  int  the difference between two differing characters

Using a character mapping, two strings are compared for "equivalence". Each input character is mapped to a comparison character and the mapped-to characters are compared for the two NUL terminated input strings. The comparison is limited to ct bytes. This function name is mapped to option strneqvcmp so as to not conflict with the POSIX name space.

none checked. Caller responsible for seg faults.

7.6.28.22 strtransform

convert a string into its mapped-to value Usage:

        strtransform( dest, src );

Where the arguments are:

Name  Type  Description
—–  —–  ————-

dest  char*  output string
src  char const*  input string

Each character in the input string is mapped and the mapped-to character is put into the output. This function name is mapped to option strtransform so as to not conflict with the POSIX name space.

none.

7.7 Option Descriptor File

This is the module that is to be compiled and linked with your program. It contains internal data and procedures subject to change. Basically, it contains a single global data structure containing all the information provided in the option definitions, plus a number of static strings and any callout procedures that are specified or required. You should never have need for looking at this, except, perhaps, to examine the code generated for implementing the flag_code construct.

7.8 Using AutoOpts

There are actually several levels of “using” autoopts. Which you choose depends upon how you plan to distribute (or not) your application.

7.8.1 local-only use

To use AutoOpts in your application where you do not have to worry about distribution issues, your issues are simple and few.

• Create a file ‘myopts.def’, according to the documentation above. It is probably easiest to start with the example in Section 7.3 [Quick Start], page 71 and edit it into the form you need.

• Run AutoGen to create the option interface file (myopts.h) and the option descriptor

    code (myopts.c):
           autogen myopts.def

• In all your source files where you need to refer to option state, #include "myopts.h".

• In your main routine, code something along the lines of:

           #define ARGC_MIN some-lower-limit
           #define ARGC_MAX some-upper-limit
           main( int argc, char** argv )
           {
               {
                   int arg_ct = optionProcess( &myprogOptions, argc, argv );
                   argc -= arg_ct;
                   if ((argc < ARGC_MIN) || (argc > ARGC_MAX)) {
                       fprintf( stderr, "%s ERROR:  remaining args (%d) "
                                "out of range\n", myprogOptions.pzProgName,
                                argc );
                       USAGE( EXIT_FAILURE );

                   }
                   argv += arg_ct;
               }
               if (HAVE_OPT(OPTN_NAME))
                   respond_to_optn_name();
               ...
           }

• Compile ‘myopts.c’ and link your program with the following additional arguments:

           myopts.c -I$prefix/include -L $prefix/lib -lopts

These values can be derived from the “autoopts-config” script:

           myopts.c `autoopts-config cflags` `autoopts-config ldflags`

7.8.2 binary distro, AutoOpts not installed

If you will be distributing (or copying) your project to a system that does not have AutoOpts installed, you will need to statically link the AutoOpts library, “libopts” into your program. Add the output from the following to your link command:

        autoopts-config static-libs

7.8.3 binary distro, AutoOpts pre-installed

If you will be distributing (or copying) your project to a system that does have AutoOpts (or only “libopts”) installed, you will still need to ensure that the library is findable at program load time, or you will still have to statically link. The former can be accomplished by linking your project with --rpath or by setting the LD_LIBRARY_PATH appropriately. Otherwise, See Section 7.8.2 [binary not installed], page 112.

7.8.4 source distro, AutoOpts pre-installed

If you will be distributing your project to a system that will build your product but it may not be pre-installed with AutoOpts, you will need to do some configuration checking before you start the build. Assuming you are willing to fail the build if AutoOpts has not been installed, you will still need to do a little work.

AutoOpts is distributed with a configuration check M4 script, ‘autoopts.m4’. It will add

an autoconf macro named, AG_PATH_AUTOOPTS. Add this to your ‘configure.ac’ script

and use the following substitution values:

AUTOGEN the name of the autogen executable

AUTOGEN_TPLIB

the directory where AutoGen template library is stored

AUTOOPTS_CFLAGS

the compile time options needed to find the AutoOpts headers

AUTOOPTS_LIBS

the link options required to access the libopts library

7.8.5 source distro, AutoOpts not installed

If you will be distributing your project to a system that will build your product but it may not be pre-installed with AutoOpts, you may wish to incorporate the sources for libopts in your project. To do this, I recommend reading the tear-off libopts library ‘README’ that you can find in the ‘pkg/libopts’ directory. You can also examine an example package (blocksort) that incorporates this tear off library in the autogen distribution directory. There is also a web page that describes what you need to do:

        http://autogen.sourceforge.net/blocksort.html

Alternatively, you can pull the libopts library sources into a build directory and build it for installation along with your package. This can be done approximately as follows:

        tar -xzvf `autoopts-config libsrc`
        cd libopts-*
        ./bootstrap
        configure
        make
        make install

That will install the library, but not the headers or anything else.

7.9 Configuring your program

AutoOpts supports the notion of “presetting” the value or state of an option. The values may be obtained either from environment variables or from configuration files (‘rc’ or ‘ini’ files). In order to take advantage of this, the AutoOpts client program must specify these features in the option descriptor file (see Section 7.5.1 [program attributes], page 73) with the rcfile or environrc attributes.

It is also possible to configure your program without using the command line option parsing code. This is done by using only the following four functions from the ‘libopts’ library:

‘configFileLoad’

(see Section 7.6.28.2 [libopts-configFileLoad], page 102) will parse the contents of a config file and return a pointer to a structure representing the hierarchical value. The values are sorted alphabetically by the value name and all entries with the same name will retain their original order. Insertion sort is used.

‘optionGetValue’

(see Section 7.6.28.7 [libopts-optionGetValue], page 104) will find the first value within the hierarchy with a name that matches the name passed in.

‘optionNextValue’

(see Section 7.6.28.9 [libopts-optionNextValue], page 105) will return the next value that follows the value passed in as an argument. If you wish to get all the values for a particular name, you must take note when the name changes.

‘optionUnloadNested’

(see Section 7.6.28.15 [libopts-optionUnloadNested], page 107). The pointer passed in must be of type, OPARG_TYPE_HIERARCHY (see the autoopts/options.h header file). configFileLoad will return a tOptionValue pointer of that type. This function will release all the associated memory. AutoOpts generated code uses this function for its own needs. Client code should only call this function with pointers gotten from configFileLoad.

7.9.1 configuration file presets

Configuration files are enabled by specifying the program attribute homerc (see Section 7.5.1 [program attributes], page 73). Any option not marked with the “no-preset” attribute may appear in a configuration file. The files loaded are selected both by the homerc entries and, optionally, via a command line option. The first component of the homerc entry may be an environment variable such as $HOME, or it may also be $$ (two dollar sign characters) to specify the directory of the executable. For example:

        homerc = "$$/../share/autogen";

will cause the AutoOpts library to look in the normal autogen datadir relative to the current installation directory for autogen.

The configuration files are processed in the order they are specified by the homerc attribute, so that each new file will normally override the settings of the previous files. This may be overridden by marking some options for immediate action (see Section 7.5.5.4 [Immediate Action], page 85). Any such options are acted upon in reverse order. The disabled load-opts (--no-load-opts) option, for example, is an immediate action option. Its presence in the last homerc file will prevent the processing of any prior homerc files because its effect is immediate.

Configuration file processing can be completely suppressed by specifying --no-load-opts on the command line, or PROGRAM_LOAD_OPTS=no in the environment (if environrc has been specified).

See the “Configuration File Format” section (see Section 7.10 [Config File Format], page 116) for details on the format of the file.

7.9.2 Saving the presets into a configuration file

When configuration files are enabled for an application, the user is also provided with an automatically supplied --save-opts option. All of the known option state will be written to either the specified output file or, if it is not specified, then to the last specified homerc file.

7.9.3 Creating a sample configuration file

AutoOpts is shipped with a template named, ‘rc-sample.tpl’. If your option definition file specifies the homerc attribute, then you may invoke ‘autogen’ thus:

        autogen -Trc-sample <your-option-def-file>

This will, by default, produce a sample file named, ‘sample-<prog-name>rc’. It will be named differently if you specify your configuration (rc) file name with the rcfile attribute. In that case, the output file will be named, ‘sample-<rcfile-name>’. It will contain all of the program options not marked as no-preset. It will also include information about how they are handled and the text from the doc attribute.

7.9.4 environment variable presets

If the AutoOpts client program specifies environrc in its option descriptor file, then environment variables will be used for presetting option state. Variables will be looked for that are named, PROGRAM_OPTNAME and PROGRAM. PROGRAM is the upper cased C-name of the program, and OPTNAME is the upper cased C-name of a specific option. (The C-names are the regular names with all special characters converted to underscores (_).)

Option specific environment variables are processed after (and thus take precedence over) the contents of the PROGRAM environment variable. The option argument string for these options takes on the string value gotten from the environment. Consequently, you can only have one instance of the OPTNAME.

If a particular option may be disabled, then its disabled state is indicated by setting the PROGRAM_OPTNAME value to the disablement prefix. So, for example, if the disablement prefix were dont, then you can disable the optname option by setting the PROGRAM_OPTNAME’ environment variable to ‘dont’. See Section 7.5.5.2 [Common Attributes], page 83.

The PROGRAM environment string is tokenized and parsed much like a command line. Doubly quoted strings have backslash escapes processed the same way they are processed in C program constant strings. Singly quoted strings are “pretty raw” in that backslashes are honored before other backslashes, apostrophes, newlines and cr/newline pairs. The options must be introduced with hyphens in the same way as the command line.

Note that not all options may be preset. Options that are specified with the no-preset attribute and the --help, --more-help, and --save-opts auto-supported options may not be preset.

7.9.5 Config file only example

If for some reason it is difficult or unworkable to integrate configuration file processing with command line option parsing, the libopts (see Section 7.6.28 [libopts procedures], page 100) library can still be used to process configuration files. Below is a “Hello, World!” greeting program that tries to load a configuration file ‘hello.conf’ to see if it should use an alternate greeting or to personalize the salutation.

        #include <sys/types.h>
        #include <stdio.h>
        #include <pwd.h>
        #include <string.h>
        #include <unistd.h>
        #include <autoopts/options.h>
        int main( int argc, char** argv ) {
          char const* greeting = "Hello";
          char const* greeted  = "World";
          const tOptionValue* pOV = configFileLoad( "hello.conf" );
          if (pOV != NULL) {
            const tOptionValue* pGetV = optionGetValue( pOV, "greeting" );
            if (  (pGetV != NULL)
               && (pGetV->valType == OPARG_TYPE_STRING))
              greeting = strdup( pGetV->v.strVal );
            pGetV = optionGetValue( pOV, "personalize" );
            if (pGetV != NULL) {
              struct passwd* pwe = getpwuid( getuid() );
              if (pwe != NULL)
                greeted = strdup( pwe->pw_gecos );

            }
            optionUnloadNested( pOV ); /* deallocate config data */
          }
          printf( "%s, %s!\n", greeting, greeted );
          return 0;
        }

With that text in a file named “hello.c”, this short script:

        cc -o hello hello.c `autoopts-config cflags ldflags`
        ./hello
        echo 'greeting Buzz off' > hello.conf
        ./hello
        echo personalize > hello.conf
        ./hello

will produce the following output (for me):

        Hello, World!
        Buzz off, World!
        Hello, Bruce Korb!

7.10 Configuration File Format

The configuration file is designed to associate names and values, much like an AutoGen Definition File (see Chapter 2 [Definitions File], page 6). Unfortunately, the file formats are different. Specifically, AutoGen Definitions provide for simpler methods for the precise control of a value string and provides for dynamically computed content. Configuration files have some established traditions in their layout. So, they are different, even though they do both allow for a single name to be associated with multiple values and they both allow for hierarchical values.

7.10.1 assigning a string value to a configurable

The basic syntax is a name followed by a value on a single line. They are separated from each other by either white space, a colon (:) or an equal sign (=). The colon or equal sign may optionally be surrounded by additional white space. If more than one value line is needed, a backslash (\) may be used to continue the value. The backslash (but not the newline) will be erased. Leading and trailing white space is always stripped from the value.

Fundamentally, it looks like this:

        name  value for that name
        name = another \
             multi-line value \
             for that name.
        name: a *third* value for ``name''

If you need more control over the content of the value, you may enclose the value in XML style brackets:

        <name>value </name>

Within these brackets you need not (must not) continue the value data with backslashes. You may also select the string formation rules to use, just add the attribute after the name,

thus: <name keep>.

‘keep’ This mode will keep all text between the brackets and not strip any white space.

‘uncooked’

This mode strips leading and trailing white space, but not do any quote processing. This is the default and need not be specified.

‘cooked’ Strings are formed and concatenated if, after stripping leading and trailing white space, the text begins and ends with either single (’) or double (") quote characters. That processing is identical to the string formation used in AutoGen definition files (see Section 3.3.2 [basic expression], page 20), except that you may not use backquotes.

And here is an example of an XML-styled value:

        <name cooked>
            "This is\n\tanother multi-line\n"
            "\tstring example."
        </name>

The string value associated with “name” will be exactly the text enclosed in quotes with the escaped characters “cooked” as you would expect (three text lines with the last line not ending with a newline, but ending with a period).

7.10.2 integer values

A name can be specified as having an integer value. To do this, you must use the XML-ish format and specify a “type” attribute for the name:

        <name type=integer> 1234 </name>

Boolean, enumeration and set membership types will be added as time allows. “type=string” is also supported, but also is the default.

7.10.3 hierarchical values

In order to specify a hierarchical value, you *must* use XML-styled formatting, specifying a type that is shorter and easier to spell:

        <structured-name type=nested>
            [[....]]
        </structured-name>

The ellipsis may be filled with any legal configuration file name/value assignments.

7.10.4 configuration file sections

Configuration files may be sectioned. If, for example, you have a collection of programs that work closely together and, likely, have a common set of options, these programs may use a single, sectioned, configuration file. The file may be sectioned in either of two ways. The two ways may not be intermixed in a single configuration file. All text before the first segmentation line is processed, then only the segment that applies:

‘[PROG_NAME]’

The file is partitioned by lines that contains an square open bracket ([), the upper-cased c-variable-syntax program name and a square close bracket (]). For example, if the prog-name program had a sectioned configuration file, then a line containing exactly ‘[PROG_NAME]’ would be processed.

‘<?program prog-name>’

The <? marker indicates an XML directive. The program directive is interpreted by the configuration file processor to segment the file in the same way as the ‘[PROG_NAME]’ sectioning is done. Any other XML directives are treated as comments.

Segmentation does not apply if the config file is being parsed with the

configFileLoad(3AutoOpts) function.

7.10.5 comments in the configuration file

Comments are lines beginning with a hash mark (#), XML-style comments (<!-- arbitrary text -->), and unrecognized XML directives.

        # this is a comment
        <!-- this is also
             a comment -->
        <?this is
          a bad comment ;->

7.11 AutoOpts for Shell Scripts

AutoOpts may be used with shell scripts either by automatically creating a complete program that will process command line options and pass back the results to the invoking shell by issuing shell variable assignment commands, or it may be used to generate portable shell code that can be inserted into your script.

The functionality of these features, of course, is somewhat constrained compared with the normal program facilities. Specifically, you cannot invoke callout procedures with either of these methods. Additionally, if you generate a shell script to do the parsing:

1. You cannot obtain options from configuration files.

2. You cannot obtain options from environment variables.

3. You cannot save the option state to an option file.

4. Option conflict/requirement verification is disabled.

Both of these methods are enabled by running AutoGen on the definitions file with the additional global attribute:

        test-main [ = proc-to-call ] ;

If you do not supply a proc-to-call, it will default to optionPutShell. That will produce a program that will process the options and generate shell text for the invoking shell to interpret (see Section 7.11.1 [binary-parser], page 119). If you supply the name, optionParseShell, then you will have a program that will generate a shell script that can parse the options (see Section 7.11.2 [script-parser], page 120). If you supply a different procedure name, you will have to provide that routine and it may do whatever you like.

7.11.1 Parsing with an Executable

The following commands are approximately all that is needed to build a shell script command line option parser from an option definition file:

autogen -L <opt-template-dir> test-errors.def
cc -o test-errors -L <opt-lib-dir> -I <opt-include-dir> \
        -DTEST_PROGRAM_OPTS test-errors.c -lopts

The resulting program can then be used within your shell script as follows:

eval `./test-errors "$@"`
if [ -z "${OPTION_CT}" ] ; then exit 1 ; fi
test ${OPTION_CT} -gt 0 && shift ${OPTION_CT}

Here is the usage output example from AutoOpts error handling tests. The option definition has argument reordering enabled:

test_errors - Test AutoOpts for errors
USAGE:  errors [ -<flag> [<val>] | --<name>[{=| }<val>] ]... arg ...
  Flg Arg Option-Name    Description
   -o no  option         The option option descrip
   -s Str second         The second option descrip
                                - may appear up to 10 times
   -X no  another        Another option descrip
                                - may appear up to 5 times
   -? no  help           Display usage information and exit
   -! no  more-help      Extended usage information passed thru pager
   -> opt save-opts      Save the option state to a config file
   -< Str load-opts      Load options from a config file
                                - disabled as --no-load-opts
                                - may appear multiple times
Options are specified by doubled hyphens and their name
or by a single hyphen and the flag character.
Operands and options may be intermixed.  They will be reordered.
The following option preset mechanisms are supported:
 - reading file errorsRC

Using the invocation,

  test-errors operand1 -s first operand2 -X -- -s operand3

you get the following output for your shell script to evaluate:

OPTION_CT=4
export OPTION_CT
TEST_ERRORS_SECOND='first'
export TEST_ERRORS_SECOND
TEST_ERRORS_ANOTHER=1 # 0x1
export TEST_ERRORS_ANOTHER
set -- 'operand1' 'operand2' '-s' 'operand3'
OPTION_CT=0

7.11.2 Parsing with a Portable Script

If you had used test-main = optionParseShell instead, then you can, at this point, merely run the program and it will write the parsing script to standard out. You may also provide this program with command line options to specify the shell script file to create or edit, and you may specify the shell program to use on the first shell script line. That program’s usage text would look something like the following and the script parser itself would be very verbose:

genshellopt - Generate Shell Option Processing Script - Ver. 1
USAGE:  genshellopt [ -<flag> [<val>] | --<name>[{=| }<val>] ]...
  Flg Arg Option-Name    Description
   -o Str script         Output Script File
   -s Str shell          Shell name (follows "#!" magic)
                                - disabled as --no-shell
                                - enabled by default
   -v opt version        Output version information and exit
   -? no  help           Display usage information and exit
   -! no  more-help      Extended usage information passed thru pager
Options are specified by doubled hyphens and their name
or by a single hyphen and the flag character.
Note that `shell' is only useful if the output file does not already
exist.  If it does, then the shell name and optional first argument
will be extracted from the script file.
If the script file already exists and contains Automated Option Processing
text, the second line of the file through the ending tag will be replaced
by the newly generated text.  The first `#!' line will be regenerated.
please send bug reports to:  autogen-users@lists.sourceforge.net
= = = = = = = =
This incarnation of genshell will produce
a shell script to parse the options for getdefs:
getdefs (GNU AutoGen) - AutoGen Definition Extraction Tool - Ver. 1.4
USAGE:  getdefs [ <option-name>[{=| }<val>] ]...
   Arg Option-Name    Description
   Str defs-to-get    Regexp to look for after the "/*="
   opt ordering       Alphabetize or use named file
   Num first-index    The first index to apply to groups
   Str input          Input file to search for defs
   Str subblock       subblock definition names
   Str listattr       attribute with list of values
   opt filelist       Insert source file names into defs

Str assign         Global assignments
Str common-assign  Assignments common to all blocks
Str copy           File(s) to copy into definitions
opt srcfile        Insert source file name into each def
opt linenum        Insert source line number into each def
Str output         Output file to open
opt autogen        Invoke AutoGen with defs
Str template       Template Name
Str agarg          AutoGen Argument
Str base-name      Base name for output file(s)
opt version        Output version information and exit
no  help           Display usage information and exit
no  more-help      Extended usage information passed thru pager
opt save-opts      Save the option state to a config file
Str load-opts      Load options from a config file
All arguments are named options.
If no ``input'' argument is provided or is set to simply "-", and if
``stdin'' is not a ``tty'', then the list of input files will be
read from ``stdin''.
please send bug reports to:  autogen-users@lists.sourceforge.net

Resulting in the following script:

#! /bin/sh
# # # # # # # # # # -- do not modify this marker --
#
#  DO NOT EDIT THIS SECTION OF ./.ag-qskc4F/genshellopt.sh
#
#  From here to the next `-- do not modify this marker --',
#  the text has been generated Saturday September 30, 2006 at 12:34:33 PM PDT
#  From the GETDEFS option definitions
#
GETDEFS_LONGUSAGE_TEXT='getdefs (GNU AutoGen) - AutoGen Definition Extraction Tool - Ver. 1.4
USAGE:  getdefs [ <option-name>[{=| }<val>] ]...
Arg Option-Name    Description
Str defs-to-get    Regexp to look for after the "/*="
opt ordering       Alphabetize or use named file
                             - disabled as --no-ordering
                             - enabled by default
Num first-index    The first index to apply to groups
Str input          Input file to search for defs
                             - may appear multiple times
                             - default option for unnamed options
Str subblock       subblock definition names
                             - may appear multiple times

Str listattr       attribute with list of values
                             - may appear multiple times
opt filelist       Insert source file names into defs
Definition insertion options
Arg Option-Name    Description
Str assign         Global assignments
                             - may appear multiple times
Str common-assign  Assignments common to all blocks
                             - may appear multiple times
Str copy           File(s) to copy into definitions
                             - may appear multiple times
opt srcfile        Insert source file name into each def
opt linenum        Insert source line number into each def
Definition output disposition options:
Arg Option-Name    Description
Str output         Output file to open
                             - an alternate for autogen
opt autogen        Invoke AutoGen with defs
                             - disabled as --no-autogen
                             - enabled by default
Str template       Template Name
Str agarg          AutoGen Argument
                             - prohibits these options:
                             output
                             - may appear multiple times
Str base-name      Base name for output file(s)
                             - prohibits these options:
                             output
version and help options:
Arg Option-Name    Description
opt version        Output version information and exit
no  help           Display usage information and exit
no  more-help      Extended usage information passed thru pager
opt save-opts      Save the option state to a config file
Str load-opts      Load options from a config file
                             - disabled as --no-load-opts
                             - may appear multiple times
All arguments are named options.
If no ``input'\'''\'' argument is provided or is set to simply "-", and if

``stdin'\'''\'' is not a ``tty'\'''\'', then the list of input files will be
read from ``stdin'\'''\''.
The following option preset mechanisms are supported:
 - reading file /dev/null
This program extracts AutoGen definitions from a list of source files.
Definitions are delimited by `/*=<entry-type> <entry-name>\n'\'' and
`=*/\n'\''.  From that, this program creates a definition of the
following form:
    #line nnn "source-file-name"
    entry_type = {
        name = entry_name;
        ...
    };
The ellipsis '\''...'\'' is filled in by text found between the two
delimiters, with everything up through the first sequence of
asterisks deleted on every line.
There are two special ``entry types'\'''\'':
*  The entry_type enclosure and the name entry will be omitted
   and the ellipsis will become top-level definitions.
-- The contents of the comment must be a single getdefs option.
   The option name must follow the double hyphen and its argument
   will be everything following the name.  This is intended for use
   with the ``subblock'\'''\'' and ``listattr'\'''\'' options.
please send bug reports to:  autogen-users@lists.sourceforge.net'
GETDEFS_USAGE_TEXT='getdefs (GNU AutoGen) - AutoGen Definition Extraction Tool - Ver. 1.4
USAGE:  getdefs [ <option-name>[{=| }<val>] ]...
   Arg Option-Name    Description
   Str defs-to-get    Regexp to look for after the "/*="
   opt ordering       Alphabetize or use named file
   Num first-index    The first index to apply to groups
   Str input          Input file to search for defs
   Str subblock       subblock definition names
   Str listattr       attribute with list of values
   opt filelist       Insert source file names into defs
   Str assign         Global assignments
   Str common-assign  Assignments common to all blocks
   Str copy           File(s) to copy into definitions
   opt srcfile        Insert source file name into each def

   opt linenum        Insert source line number into each def
   Str output         Output file to open
   opt autogen        Invoke AutoGen with defs
   Str template       Template Name
   Str agarg          AutoGen Argument
   Str base-name      Base name for output file(s)
   opt version        Output version information and exit
   no  help           Display usage information and exit
   no  more-help      Extended usage information passed thru pager
   opt save-opts      Save the option state to a config file
   Str load-opts      Load options from a config file
All arguments are named options.
If no ``input'\'''\'' argument is provided or is set to simply "-", and if
``stdin'\'''\'' is not a ``tty'\'''\'', then the list of input files will be
read from ``stdin'\'''\''.
please send bug reports to:  autogen-users@lists.sourceforge.net'

GETDEFS_DEFS_TO_GET="${GETDEFS_DEFS_TO_GET}"
GETDEFS_DEFS_TO_GET_set=false
export GETDEFS_DEFS_TO_GET
GETDEFS_ORDERING="${GETDEFS_ORDERING}"
GETDEFS_ORDERING_set=false
export GETDEFS_ORDERING
GETDEFS_FIRST_INDEX="${GETDEFS_FIRST_INDEX-'0'}"
GETDEFS_FIRST_INDEX_set=false
export GETDEFS_FIRST_INDEX
if test -z "${GETDEFS_INPUT}"
then
  GETDEFS_INPUT_CT=0
else
  GETDEFS_INPUT_CT=1
  GETDEFS_INPUT_1="${GETDEFS_INPUT}"
fi
export GETDEFS_INPUT_CT
if test -z "${GETDEFS_SUBBLOCK}"
then
  GETDEFS_SUBBLOCK_CT=0
else
  GETDEFS_SUBBLOCK_CT=1
  GETDEFS_SUBBLOCK_1="${GETDEFS_SUBBLOCK}"

fi
export GETDEFS_SUBBLOCK_CT
if test -z "${GETDEFS_LISTATTR}"
then
  GETDEFS_LISTATTR_CT=0
else
  GETDEFS_LISTATTR_CT=1
  GETDEFS_LISTATTR_1="${GETDEFS_LISTATTR}"
fi
export GETDEFS_LISTATTR_CT
GETDEFS_FILELIST="${GETDEFS_FILELIST}"
GETDEFS_FILELIST_set=false
export GETDEFS_FILELIST
if test -z "${GETDEFS_ASSIGN}"
then
  GETDEFS_ASSIGN_CT=0
else
  GETDEFS_ASSIGN_CT=1
  GETDEFS_ASSIGN_1="${GETDEFS_ASSIGN}"
fi
export GETDEFS_ASSIGN_CT
if test -z "${GETDEFS_COMMON_ASSIGN}"
then
  GETDEFS_COMMON_ASSIGN_CT=0
else
  GETDEFS_COMMON_ASSIGN_CT=1
  GETDEFS_COMMON_ASSIGN_1="${GETDEFS_COMMON_ASSIGN}"
fi
export GETDEFS_COMMON_ASSIGN_CT
if test -z "${GETDEFS_COPY}"
then
  GETDEFS_COPY_CT=0
else
  GETDEFS_COPY_CT=1
  GETDEFS_COPY_1="${GETDEFS_COPY}"
fi
export GETDEFS_COPY_CT
GETDEFS_SRCFILE="${GETDEFS_SRCFILE}"
GETDEFS_SRCFILE_set=false
export GETDEFS_SRCFILE
GETDEFS_LINENUM="${GETDEFS_LINENUM}"
GETDEFS_LINENUM_set=false
export GETDEFS_LINENUM
GETDEFS_OUTPUT="${GETDEFS_OUTPUT}"

GETDEFS_OUTPUT_set=false
export GETDEFS_OUTPUT
GETDEFS_AUTOGEN="${GETDEFS_AUTOGEN}"
GETDEFS_AUTOGEN_set=false
export GETDEFS_AUTOGEN
GETDEFS_TEMPLATE="${GETDEFS_TEMPLATE}"
GETDEFS_TEMPLATE_set=false
export GETDEFS_TEMPLATE
if test -z "${GETDEFS_AGARG}"
then
  GETDEFS_AGARG_CT=0
else
  GETDEFS_AGARG_CT=1
  GETDEFS_AGARG_1="${GETDEFS_AGARG}"
fi
export GETDEFS_AGARG_CT
GETDEFS_BASE_NAME="${GETDEFS_BASE_NAME}"
GETDEFS_BASE_NAME_set=false
export GETDEFS_BASE_NAME
OPT_ARG="$1"
while [ $# -gt 0 ]
do
    OPT_ELEMENT=''
    OPT_ARG_VAL=''
    OPT_ARG="${1}"
        OPT_CODE=`echo "X${OPT_ARG}"|sed 's/^X-*//'`
        shift
        OPT_ARG="$1"
        case "${OPT_CODE}" in *=* )
            OPT_ARG_VAL=`echo "${OPT_CODE}"|sed 's/^[^=]*=//'`
            OPT_CODE=`echo "${OPT_CODE}"|sed 's/=.*$//'` ;; esac
        case "${OPT_CODE}" in
        'de' | \
        'def' | \
        'defs' | \
        'defs-' | \
        'defs-t' | \
        'defs-to' | \
        'defs-to-' | \

'defs-to-g' | \
'defs-to-ge' | \
'defs-to-get' )
    if [ -n "${GETDEFS_DEFS_TO_GET}" ] && ${GETDEFS_DEFS_TO_GET_set} ; then
        echo Error:  duplicate DEFS_TO_GET option >&2
        echo "$GETDEFS_USAGE_TEXT"
        exit 1 ; fi
    GETDEFS_DEFS_TO_GET_set=true
    OPT_NAME='DEFS_TO_GET'
    OPT_ARG_NEEDED=YES
    ;;
'or' | \
'ord' | \
'orde' | \
'order' | \
'orderi' | \
'orderin' | \
'ordering' )
    if [ -n "${GETDEFS_ORDERING}" ] && ${GETDEFS_ORDERING_set} ; then
        echo Error:  duplicate ORDERING option >&2
        echo "$GETDEFS_USAGE_TEXT"
        exit 1 ; fi
    GETDEFS_ORDERING_set=true
    OPT_NAME='ORDERING'
    eval GETDEFS_ORDERING${OPT_ELEMENT}=true
    export GETDEFS_ORDERING${OPT_ELEMENT}
    OPT_ARG_NEEDED=OK
    ;;
'no-o' | \
'no-or' | \
'no-ord' | \
'no-orde' | \
'no-order' | \
'no-orderi' | \
'no-orderin' | \
'no-ordering' )
    if [ -n "${GETDEFS_ORDERING}" ] && ${GETDEFS_ORDERING_set} ; then
        echo Error:  duplicate ORDERING option >&2
        echo "$GETDEFS_USAGE_TEXT"
        exit 1 ; fi
    GETDEFS_ORDERING_set=true
    GETDEFS_ORDERING='no'
    export GETDEFS_ORDERING
    OPT_NAME='ORDERING'
    OPT_ARG_NEEDED=NO

    ;;
'fir' | \
'firs' | \
'first' | \
'first-' | \
'first-i' | \
'first-in' | \
'first-ind' | \
'first-inde' | \
'first-index' )
    if [ -n "${GETDEFS_FIRST_INDEX}" ] && ${GETDEFS_FIRST_INDEX_set} ; then
        echo Error:  duplicate FIRST_INDEX option >&2
        echo "$GETDEFS_USAGE_TEXT"
        exit 1 ; fi
    GETDEFS_FIRST_INDEX_set=true
    OPT_NAME='FIRST_INDEX'
    OPT_ARG_NEEDED=YES
    ;;
'in' | \
'inp' | \
'inpu' | \
'input' )
    GETDEFS_INPUT_CT=`expr ${GETDEFS_INPUT_CT} + 1`
    OPT_ELEMENT="_${GETDEFS_INPUT_CT}"
    OPT_NAME='INPUT'
    OPT_ARG_NEEDED=YES
    ;;
'su' | \
'sub' | \
'subb' | \
'subbl' | \
'subblo' | \
'subbloc' | \
'subblock' )
    GETDEFS_SUBBLOCK_CT=`expr ${GETDEFS_SUBBLOCK_CT} + 1`
    OPT_ELEMENT="_${GETDEFS_SUBBLOCK_CT}"
    OPT_NAME='SUBBLOCK'
    OPT_ARG_NEEDED=YES
    ;;
'lis' | \
'list' | \
'lista' | \
'listat' | \

'listatt' | \
'listattr' )
    GETDEFS_LISTATTR_CT=`expr ${GETDEFS_LISTATTR_CT} + 1`
    OPT_ELEMENT="_${GETDEFS_LISTATTR_CT}"
    OPT_NAME='LISTATTR'
    OPT_ARG_NEEDED=YES
    ;;
'fil' | \
'file' | \
'filel' | \
'fileli' | \
'filelis' | \
'filelist' )
    if [ -n "${GETDEFS_FILELIST}" ] && ${GETDEFS_FILELIST_set} ; then
        echo Error:  duplicate FILELIST option >&2
        echo "$GETDEFS_USAGE_TEXT"
        exit 1 ; fi
    GETDEFS_FILELIST_set=true
    OPT_NAME='FILELIST'
    eval GETDEFS_FILELIST${OPT_ELEMENT}=true
    export GETDEFS_FILELIST${OPT_ELEMENT}
    OPT_ARG_NEEDED=OK
    ;;
'as' | \
'ass' | \
'assi' | \
'assig' | \
'assign' )
    GETDEFS_ASSIGN_CT=`expr ${GETDEFS_ASSIGN_CT} + 1`
    OPT_ELEMENT="_${GETDEFS_ASSIGN_CT}"
    OPT_NAME='ASSIGN'
    OPT_ARG_NEEDED=YES
    ;;
'com' | \
'comm' | \
'commo' | \
'common' | \
'common-' | \
'common-a' | \
'common-as' | \
'common-ass' | \
'common-assi' | \
'common-assig' | \
'common-assign' )

GETDEFS_COMMON_ASSIGN_CT=`expr ${GETDEFS_COMMON_ASSIGN_CT} + 1`
OPT_ELEMENT="_${GETDEFS_COMMON_ASSIGN_CT}"
OPT_NAME='COMMON_ASSIGN'
OPT_ARG_NEEDED=YES
;;
'cop' | \
'copy' )
GETDEFS_COPY_CT=`expr ${GETDEFS_COPY_CT} + 1`
OPT_ELEMENT="_${GETDEFS_COPY_CT}"
OPT_NAME='COPY'
OPT_ARG_NEEDED=YES
;;
'sr' | \
'src' | \
'srcf' | \
'srcfi' | \
'srcfil' | \
'srcfile' )
if [ -n "${GETDEFS_SRCFILE}" ] && ${GETDEFS_SRCFILE_set} ; then
    echo Error:  duplicate SRCFILE option >&2
    echo "$GETDEFS_USAGE_TEXT"
    exit 1 ; fi
GETDEFS_SRCFILE_set=true
OPT_NAME='SRCFILE'
eval GETDEFS_SRCFILE${OPT_ELEMENT}=true
export GETDEFS_SRCFILE${OPT_ELEMENT}
OPT_ARG_NEEDED=OK
;;
'lin' | \
'line' | \
'linen' | \
'linenu' | \
'linenum' )
if [ -n "${GETDEFS_LINENUM}" ] && ${GETDEFS_LINENUM_set} ; then
    echo Error:  duplicate LINENUM option >&2
    echo "$GETDEFS_USAGE_TEXT"
    exit 1 ; fi
GETDEFS_LINENUM_set=true
OPT_NAME='LINENUM'
eval GETDEFS_LINENUM${OPT_ELEMENT}=true
export GETDEFS_LINENUM${OPT_ELEMENT}
OPT_ARG_NEEDED=OK
;;

'ou' | \
'out' | \
'outp' | \
'outpu' | \
'output' )
    if [ -n "${GETDEFS_OUTPUT}" ] && ${GETDEFS_OUTPUT_set} ; then
        echo Error:  duplicate OUTPUT option >&2
        echo "$GETDEFS_USAGE_TEXT"
        exit 1 ; fi
    GETDEFS_OUTPUT_set=true
    OPT_NAME='OUTPUT'
    OPT_ARG_NEEDED=YES
    ;;
'au' | \
'aut' | \
'auto' | \
'autog' | \
'autoge' | \
'autogen' )
    if [ -n "${GETDEFS_AUTOGEN}" ] && ${GETDEFS_AUTOGEN_set} ; then
        echo Error:  duplicate AUTOGEN option >&2
        echo "$GETDEFS_USAGE_TEXT"
        exit 1 ; fi
    GETDEFS_AUTOGEN_set=true
    OPT_NAME='AUTOGEN'
    eval GETDEFS_AUTOGEN${OPT_ELEMENT}=true
    export GETDEFS_AUTOGEN${OPT_ELEMENT}
    OPT_ARG_NEEDED=OK
    ;;
'no-a' | \
'no-au' | \
'no-aut' | \
'no-auto' | \
'no-autog' | \
'no-autoge' | \
'no-autogen' )
    if [ -n "${GETDEFS_AUTOGEN}" ] && ${GETDEFS_AUTOGEN_set} ; then
        echo Error:  duplicate AUTOGEN option >&2
        echo "$GETDEFS_USAGE_TEXT"
        exit 1 ; fi
    GETDEFS_AUTOGEN_set=true
    GETDEFS_AUTOGEN='no'
    export GETDEFS_AUTOGEN
    OPT_NAME='AUTOGEN'
    OPT_ARG_NEEDED=NO

    ;;
'te' | \
'tem' | \
'temp' | \
'templ' | \
'templa' | \
'templat' | \
'template' )
    if [ -n "${GETDEFS_TEMPLATE}" ] && ${GETDEFS_TEMPLATE_set} ; then
        echo Error:  duplicate TEMPLATE option >&2
        echo "$GETDEFS_USAGE_TEXT"
        exit 1 ; fi
    GETDEFS_TEMPLATE_set=true
    OPT_NAME='TEMPLATE'
    OPT_ARG_NEEDED=YES
    ;;
'ag' | \
'aga' | \
'agar' | \
'agarg' )
    GETDEFS_AGARG_CT=`expr ${GETDEFS_AGARG_CT} + 1`
    OPT_ELEMENT="_${GETDEFS_AGARG_CT}"
    OPT_NAME='AGARG'
    OPT_ARG_NEEDED=YES
    ;;
'ba' | \
'bas' | \
'base' | \
'base-' | \
'base-n' | \
'base-na' | \
'base-nam' | \
'base-name' )
    if [ -n "${GETDEFS_BASE_NAME}" ] && ${GETDEFS_BASE_NAME_set} ; then
        echo Error:  duplicate BASE_NAME option >&2
        echo "$GETDEFS_USAGE_TEXT"
        exit 1 ; fi
    GETDEFS_BASE_NAME_set=true
    OPT_NAME='BASE_NAME'
    OPT_ARG_NEEDED=YES
    ;;
've' | \
'ver' | \

'vers' | \
'versi' | \
'versio' | \
'version' )
    echo "$GETDEFS_LONGUSAGE_TEXT"
    exit 0
    ;;
'he' | \
'hel' | \
'help' )
    echo "$GETDEFS_LONGUSAGE_TEXT"
    exit 0
    ;;
'mo' | \
'mor' | \
'more' | \
'more-' | \
'more-h' | \
'more-he' | \
'more-hel' | \
'more-help' )
    echo "$GETDEFS_LONGUSAGE_TEXT" | ${PAGER-more}
    exit 0
    ;;
'sa' | \
'sav' | \
'save' | \
'save-' | \
'save-o' | \
'save-op' | \
'save-opt' | \
'save-opts' )
    echo 'Warning:  Cannot save options files' >&2
    OPT_ARG_NEEDED=OK
    ;;
'lo' | \
'loa' | \
'load' | \
'load-' | \
'load-o' | \
'load-op' | \
'load-opt' | \
'load-opts' )

echo 'Warning:  Cannot load options files' >&2
OPT_ARG_NEEDED=YES
;;
'no-l' | \
'no-lo' | \
'no-loa' | \
'no-load' | \
'no-load-' | \
'no-load-o' | \
'no-load-op' | \
'no-load-opt' | \
'no-load-opts' )
echo 'Warning:  Cannot suppress the loading of options files' >&2
OPT_ARG_NEEDED=NO
;;
* )
echo Unknown option: "${OPT_CODE}" >&2
echo "$GETDEFS_USAGE_TEXT"
exit 1
;;
esac
case "${OPT_ARG_NEEDED}" in
NO )
OPT_ARG_VAL=''
;;
YES )
if [ -z "${OPT_ARG_VAL}" ]
then
    if [ $# -eq 0 ]
    then
        echo No argument provided for ${OPT_NAME} option >&2
        echo "$GETDEFS_USAGE_TEXT"
        exit 1
    fi
    OPT_ARG_VAL="${OPT_ARG}"
    shift
    OPT_ARG="$1"
fi
;;
OK )
if [ -z "${OPT_ARG_VAL}" ] && [ $# -gt 0 ]

            then
                case "${OPT_ARG}" in -* ) ;; * )
                    OPT_ARG_VAL="${OPT_ARG}"
                    shift
                    OPT_ARG="$1" ;; esac
            fi
            ;;
        esac
    if [ -n "${OPT_ARG_VAL}" ]
    then
        eval GETDEFS_${OPT_NAME}${OPT_ELEMENT}="'${OPT_ARG_VAL}'"
        export GETDEFS_${OPT_NAME}${OPT_ELEMENT}
    fi
done
unset OPT_PROCESS || :
unset OPT_ELEMENT || :
unset OPT_ARG || :
unset OPT_ARG_NEEDED || :
unset OPT_NAME || :
unset OPT_CODE || :
unset OPT_ARG_VAL || :
# # # # # # # # # #
#
#  END OF AUTOMATED OPTION PROCESSING
#
# # # # # # # # # # -- do not modify this marker --
env | egrep GETDEFS_

7.12 Automated Info Docs

AutoOpts provides two templates for producing ‘.texi’ documentation. ‘aginfo.tpl’ for the invoking section, and ‘aginfo3.tpl’ for describing exported library functions and macros.

For both types of documents, the documentation level is selected by passing a ‘-DLEVEL=<level-name>’ argument to AutoGen when you build the document. (See the example invocation below.)

Two files will be produced, a ‘.texi’ file and a ‘.menu’ file. You should include the ‘.menu’ file in your document where you wish to reference the ‘invoking’ chapter, section or subsection.

The ‘.texi’ file will contain an introductory paragraph, a menu and a subordinate section for the invocation usage and for each documented option. The introductory paragraph is normally the boiler plate text, along the lines of:

This chapter documents the @file{AutoOpts} generated usage text

        and option meanings for the @file{your-program} program.

or:

        These are the publicly exported procedures from the libname library.
        Any other functions mentioned in the header file are for the private use
        of the library.

7.12.1 “invoking” info docs

Using the option definitions for an AutoOpt client program, the ‘aginfo.tpl’ template will produce texinfo text that documents the invocation of your program. The text emitted is designed to be included in the full texinfo document for your product. It is not a standalone document. The usage text for the Section 5.1 [autogen usage], page 56, Section 8.5.1 [getdefs usage], page 146 and Section 8.4.1 [columns usage], page 142 programs, are included in this document and are all generated using this template.

If your program’s option definitions include a ‘prog-info-descrip’ section, then that text will replace the boilerplate introductory paragraph.

These files are produced by invoking the following command:

        autogen -L ${prefix}/share/autogen -T aginfo.tpl \
                -DLEVEL=section your-opts.def

Where ‘${prefix}’ is the AutoGen installation prefix and ‘your-opts.def’ is the name of your product’s option definition file.

7.12.2 library info docs

The ‘texinfo’ doc for libraries is derived from mostly the same information as is used for producing man pages See Section 7.13.2 [man3], page 137. The main difference is that there is only one output file and the individual functions are referenced from a .texi menu. There is also a small difference in the global attributes used:

lib description A description of the library. This text appears before the menu. If not provided, the standard boilerplate version will be inserted.

see also The SEE ALSO functionality is not supported for the ‘texinfo’ documentation, so any see_also attribute will be ignored.

These files are produced by invoking the following commands:

        getdefs linenum srcfile template=aginfo3.tpl output=libexport.def \
               <source-file-list>
        autogen -L ${prefix}/share/autogen -DLEVEL=section libexport.def

Where ‘${prefix}’ is the AutoGen installation prefix and ‘libexport.def’ is some name that suits you.

An example of this can be seen in this document, See Section 7.6.28 [libopts procedures], page 100.

7.13 Automated Man Pages

AutoOpts provides two templates for producing man pages. The command (‘man1’) pages are derived from the options definition file, and the library (‘man3’) pages are derived from stylized comments (see Section 8.5 [getdefs Invocation], page 145).

7.13.1 command line man pages

Using the option definitions for an AutoOpts client program, the ‘agman1.tpl’ template will produce an nroff document suitable for use as a ‘man(1)’ page document for a command line command. The description section of the document is either the ‘prog-man-descrip’ text, if present, or the ‘detail’ text.

Each option in the option definitions file is fully documented in its usage. This includes all the information documented above for each option (see Section 7.5.5 [option attributes], page 82), plus the ‘doc’ attribute is appended. Since the ‘doc’ text is presumed to be designed for texinfo documentation, sed is used to convert some constructs from texi to nroff-for-man-pages. Specifically,

        convert @code, @var and @samp into \fB...\fP phrases
        convert @file into \fI...\fP phrases
        Remove the '@' prefix from curly braces
        Indent example regions
        Delete the example commands
        Replace `end example' command with ".br"
        Replace the `@*' command with ".br"

This document is produced by invoking the following command:

        autogen -L ${prefix}/share/autogen -T agman1.tpl options.def

Where ‘${prefix}’ is the AutoGen installation prefix and ‘options.def’ is the name of your product’s option definition file. I do not use this very much, so any feedback or improvements would be greatly appreciated.

7.13.2 library man pages

Two global definitions are required, and then one library man page is produced for each export_func definition that is found. It is generally convenient to place these definitions as ‘getdefs’ comments (see Section 8.5 [getdefs Invocation], page 145) near the procedure definition, but they may also be a separate AutoGen definitions file (see Chapter 2 [Definitions File], page 6). Each function will be cross referenced with their sister functions in a ‘SEE ALSO’ section. A global see_also definition will be appended to this cross referencing text. The two global definitions required are:

library This is the name of your library, without the ‘lib’ prefix. The AutoOpts library is named ‘libopts.so...’, so the library attribute would have the value opts.

header Generally, using a library with a compiled program entails #include-ing a header file. Name that header with this attribute. In the case of AutoOpts, it is generated and will vary based on the name of the option definition file. Consequently, ‘your-opts.h’ is specified.

The export_func definition should contain the following attributes:

name  The name of the procedure the library user may call.
what  A brief sentence describing what the procedure does.

doc A detailed description of what the procedure does. It may ramble on for as long as necessary to properly describe it.

err  A short description of how errors are handled.

ret type The data type returned by the procedure. Omit this for void procedures.

ret desc  Describe what the returned value is, if needed.

private If specified, the function will not be documented. This is used, for example, to produce external declarations for functions that are not available for public use, but are used in the generated text.

arg  This is a compound attribute that contains:
arg type  The data type of the argument.
arg name A short name for it.
arg desc  A brief description.

As a ‘getdefs’ comment, this would appear something like this:

        /*=--subblock=arg=arg_type,arg_name,arg_desc =*/
        /*=*
         * library: opts
         * header:  your-opts.h
        =*/
        /*=export_func optionProcess
         *
         * what: this is the main option processing routine
         * arg:  + tOptions* + pOpts + program options descriptor +
         * arg:  + int       + argc  + program arg count  +
         * arg:  + char**    + argv  + program arg vector +
         * ret_type:  int
         * ret_desc:  the count of the arguments processed
         *
         * doc:  This is what it does.
         * err:  When it can't, it does this.
        =*/

Note the subblock and library comments. subblock is an embedded ‘getdefs’ option (see Section 8.5.6 [getdefs subblock], page 148) that tells it how to parse the arg attribute. The library and header entries are global definitions that apply to all the documented functions.

7.14 Using getopt(3C)

There is now a template named, “getopt.tpl” that is distributed with autoopts. With it, you will have another source file generated for you that will utilize either the standard getopt(3C) or the GNU getopt_long(3GNU) function for parsing the command line arguments. Which is used is selected by the presence or absence of the long-opts program attribute. It will save you from being dependent upon the libopts library and it produces code ready for internationalization. However, it also carries with it some limitations on the use of AutoOpts features:

1. You cannot automatically take advantage of environment variable options or rc (ini) files.

2. You cannot use set membership, enumerated, range checked or stacked argument type options. In fact, you cannot use anything that depends upon the libopts library. You are constrained to options that take “string” arguments, though you may handle the option argument with a callback procedure.

3. You must specify every option as “settable” because the emitted code depends upon the SET_OPT_XXX macros having been defined.

4. You must specify a main procedure of type “main”. The ‘getopt.tpl’ template depends upon being able to compile the traditional .c file into a program and get it to emit the usage text.

5. For the same reason, the traditional option parsing table code must be emitted before the ‘getopt.tpl’ template gets expanded.

6. The usage text is, therefore, statically defined.

7. You must supply some compile and link options via environment variables.

‘srcdir’ In case the option definition file lives in a different directory.

‘CFLAGS’ Any special flags required to compile. This should minimally include the output from running the autoopts-config cflags script.

‘LDFLAGS’ Any special flags required to link. This should minimally include the output

              from running the autoopts-config ldflags script.

‘CC’ Set this only if “cc” cannot be found in $PATH (or it is not the one you want).

To use this, set the exported environment variables and then invoke autogen twice, in the following order:

        autogen myprog-opts.def
        autogen -T getopt.tpl myprog-opts.def

and you will have three new files: ‘myprog-opts.h’, ‘myprog-opts.c’, and ‘getopt-progname.c’, where “progname” is the name specified with the global prog-name attribute in the option definition file.

7.15 Internationalizing AutoOpts

The generated code for AutoOpts will enable and disable the translation of AutoOpts run time messages. If ENABLE_NLS is defined at compile time, then the _() macro may be used to specify a translation function. If undefined, it will default to gettext(3GNU). This define will also enable a callback function that optionProcess invokes at the beginning of option processing. The AutoOpts libopts library will always check for this “compiled with NLS” flag, so libopts does not need to be specially compiled. The strings returned by the translation function will be strdup(3)-ed and kept. They will not be re-translated, even if the locale changes, but they will also not be dependent upon reused or unmappable memory.

To internationalize option processing, you should first internationalize your program. Then, the option processing strings can be added to your translation text by processing the AutoOpts-generated ‘my-opts.c’ file and adding the distributed ‘po/usage-txt.pot’ file. (Also by extracting the strings yourself from the ‘usage-txt.h’ file.) When you call optionProcess, all of the user visible AutoOpts strings will be passed through the localization procedure established with the _() preprocessing macro.

7.16 Naming Conflicts

AutoOpts generates a header file that contains many C preprocessing macros and several external names. For the most part, they begin with either opt_or option, or else they end with _opt. If this happens to conflict with other macros you are using, or if you are compiling multiple option sets in the same compilation unit, the conflicts can be avoided. You may specify an external name prefix (see Section 7.5.1 [program attributes], page 73) for all of the names generated for each set of option definitions.

Among these macros, several take an option name as a macro argument. Sometimes, this will inconveniently conflict. For example, if you specify an option named, debug, the emitted code will presume that DEBUG is not a preprocessing name. Or also, if you are building on a Windows platform, you may find that MicroSoft has usurped a number of user space names in its header files. Consequently, you will get a preprocessing error if you use, for example, HAVE_OPT(DEBUG) or HAVE_OPT(INTERNAL) (see Section 7.6.9 [HAVE OPT], page 96) in your code. You may trigger an obvious warning for such conflicts by specifying the guard-option-names attribute (see Section 7.5.1 [program attributes], page 73). That emitted code will also #undef-ine the conflicting name.

8 Add-on packages for AutoGen

This chapter includes several programs that either work closely with AutoGen (extracting definitions or providing special formatting functions), or leverage off of AutoGen technology. There is also a formatting library that helps make AutoGen possible.

AutoOpts ought to appear in this list as well, but since it is the primary reason why many people would even look into AutoGen at all, I decided to leave it in the list of chapters.

8.1 Automated Finite State Machine

The templates to generate a finite state machine in C or C++ is included with AutoGen. The documentation is not. The documentation is in HTML format for viewing

(http://www.gnu.org/software/autogen/autofsm.html), or you can download FSM
(http://download.sourceforge.net/autogen/).

8.2 Combined RPC Marshalling

The templates and NFSv4 definitions are not included with AutoGen in any way. The folks that designed NFSv4 noticed that much time and bandwidth was wasted sending queries and responses when many of them could be bundled. The protocol bundles the data, but there is no support for it in rpcgen. That means you have to write your own code to do that. Until now. Download this and you will have a large, complex example of how to use AutoXDR for generating the marshaling and unmarshaling of combined RPC calls. There is

a brief example on the web (http://www.gnu.org/software/autogen/xdr/index.html),
but you should download AutoXDR (http://download.sourceforge.net/autogen/).

8.3 Automated Event Management

Large software development projects invariably have a need to manage the distribution and display of state information and state changes. In other words, they need to manage their software events. Generally, each such project invents its own way of accomplishing this and then struggles to get all of its components to play the same way. It is a difficult process and not always completely successful. This project helps with that.

AutoEvents completely separates the tasks of supplying the data needed for a particular event from the methods used to manage the distribution and display of that event. Consequently, the programmer writing the code no longer has to worry about that part of the problem. Likewise the persons responsible for designing the event management and distribution no longer have to worry about getting programmers to write conforming code.

   This is a work in progress. See my web page (http://www.gnu.org/software/autogen/autoevents.html)

on the subject, if you are interested. I have some useful things put together, but it is not ready to call a product.

8.4 Invoking columns

This program has no explanation.

This program was designed for the purpose of generating compact, columnized tables. It will read a list of text items from standard in or a specified input file and produce a columnized listing of all the non-blank lines. Leading white space on each line is preserved, but trailing white space is stripped. Methods of applying per-entry and per-line embellishments are provided. See the formatting and separation arguments below.

This program is used by AutoGen to help clean up and organize its output.

   See ‘autogen/agen5/fsm.tpl’ and the generated output ‘pseudo-fsm.h’.

This function was not implemented as an expression function because either it would have to be many expression functions, or a provision would have to be added to provide options to expression functions. Maybe not a bad idea, but it is not being implemented at the moment.

A side benefit is that you can use it outside of AutoGen to columnize input, a la the ls command.

This section was generated by AutoGen, the aginfo template and the option descriptions for the columns program. It documents the columns usage text and option meanings.

This software is released under the GNU General Public License.

8.4.1 columns usage help (-?)

This is the automatically generated usage text for columns:

columns (GNU AutoGen) - Columnize Input Text - Ver. 1.1
USAGE:  columns [ -<flag> [<val>] | --<name>[{=| }<val>] ]...
  Flg Arg Option-Name    Description
   -W Num width          Maximum Line Width
   -c Num columns        Desired number of columns
   -w Num col-width      Set width of each column
      Num spread         maximum spread added to column width
   -I Str indent         Line prefix or indentation
      Str first-indent   First line prefix
                                - requires these options:
                                indent
      Num tab-width      tab width
   -s opt sort           Sort input text
   -f Str format         Formatting string for each input
   -S Str separation     Separation string - follows all but last
      Str line-separation string at end of all lines but last
      no  by-columns     Print entries in column order
   -i Str input          Input file (if not stdin)
   -v opt version        Output version information and exit
   -? no  help           Display usage information and exit
   -! no  more-help      Extended usage information passed thru pager
Options are specified by doubled hyphens and their name

or by a single hyphen and the flag character.
This program was designed for the purpose of generating compact,
columnized tables.  It will read a list of text items from standard
in or a specified input file and produce a columnized listing of
all the non-blank lines.  Leading white space on each line is
preserved, but trailing white space is stripped.  Methods of
applying per-entry and per-line embellishments are provided.
See the formatting and separation arguments below.
This program is used by AutoGen to help clean up and organize
its output.
please send bug reports to:  autogen-users@lists.sourceforge.net

8.4.2 width option (-W)

This is the “maximum line width” option. This option specifies the full width of the output line, including any start-of-line indentation. The output will fill each line as completely as possible, unless the column width has been explicitly specified. If the maximum width is less than the length of the widest input, you will get a single column of output.

8.4.3 columns option (-c)

This is the “desired number of columns” option. Use this option to specify exactly how many columns to produce. If that many columns will not fit within line width, then the count will be reduced to the number that fit.

8.4.4 col-width option (-w)

This is the “set width of each column” option. Use this option to specify exactly how many characters are to be allocated for each column. If it is narrower than the widest entry, it will be over-ridden with the required width.

8.4.5 spread option

This is the “maximum spread added to column width” option. Use this option to specify exactly how many characters may be added to each column. It allows you to prevent columns from becoming too far apart.

8.4.6 indent option (-I)

This is the “line prefix or indentation” option. If a number, then this many spaces will be inserted at the start of every line. Otherwise, it is a line prefix that will be inserted at the start of every line.

8.4.7 first-indent option

This is the “first line prefix” option.

This option has some usage constraints. It:

• must appear in combination with the following options: indent.

If a number, then this many spaces will be inserted at the start of the first line. Otherwise, it is a line prefix that will be inserted at the start of that line.

8.4.8 tab-width option

This is the “tab width” option. If an indentation string contains tabs, then this value is used to compute the ending column of the prefix string.

8.4.9 sort option (-s)

This is the “sort input text” option. Causes the input text to be sorted. If an argument is supplied, it is presumed to be a pattern and the sort is based upon the matched text. If the pattern starts with or consists of an asterisk (*), then the sort is case insensitive.

8.4.10 format option (-f)

This is the “formatting string for each input” option. If you need to reformat each input text, the argument to this option is interpreted as an sprintf(3) format that is used to produce each output entry.

8.4.11 separation option (-S)

This is the “separation string - follows all but last” option. Use this option if, for example, you wish a comma to appear after each entry except the last.

8.4.12 line-separation option

This is the “string at end of all lines but last” option. Use this option if, for example, you wish a backslash to appear at the end of every line, except the last.

8.4.13 by-columns option

This is the “print entries in column order” option. Normally, the entries are printed out in order by rows and then columns. This option will cause the entries to be ordered within columns. The final column, instead of the final row, may be shorter than the others.

8.4.14 input option (-i)

This is the “input file (if not stdin)” option. This program normally runs as a filter, reading from standard input, columnizing and writing to standard out. This option redirects input to a file.

8.5 Invoking getdefs

If no input argument is provided or is set to simply "-", and if stdin is not a tty, then the list of input files will be read from stdin. This program extracts AutoGen definitions from a list of source files. Definitions are delimited by ‘/*=<entry-type> <entry-name>\n’ and ‘=*/\n’. From that, this program creates a definition of the following form:

#line nnn "source-file-name" entry type = {

name = entry name; ...

};

The ellipsis “...” is filled in by text found between the two delimiters, using the following rules:

1. Each entry is located by the pattern "\n[^*\n]*\\*[ \t]*([a-z][a-z0-9 ]*):". Fundamentally, it finds a line that, after the first asterisk on the line, contains whitespace then a name and is immediately followed by a colon. The name becomes the name of the attribute and what follows, up to the next attribute, is its value.

2. If the first character of the value is either a single or double quote, then you are responsible for quoting the text as it gets inserted into the output definitions.

3. All the leading text on a line is stripped from the value. The leading text is everything before the first asterisk, the asterisk and all the whitespace characters that immediately follow it. If you want whitespace at the beginnings of the lines of text, you must do something like this:

         * mumble:
         * "  this is some\n"
         * "  indented text."

4. If the ‘<entry-name>’ is followed by a comma, the word ‘ifdef’ (or ‘ifndef’) and a name

‘if name’, then the above entry will appear as:

        #ifdef if_name
        #line nnn "source-file-name"
        entry_type = {
            name = entry_name;
            ...
        };
        #endif

5. If you use of the subblock option, you can specify a nested value, See Section 8.5.6

[getdefs subblock], page 148. That is, this text:

         * arg:  int, this, what-it-is

with the ‘–subblock=arg=type,name,doc’ option would yield:

        arg = { type = int; name = this; doc = what-it-is; };

This section was generated by AutoGen, the aginfo template and the option descriptions for the getdefs program. It documents the getdefs usage text and option meanings.

This software is released under the GNU General Public License.

8.5.1 getdefs usage help

This is the automatically generated usage text for getdefs:

getdefs (GNU AutoGen) - AutoGen Definition Extraction Tool - Ver. 1.4
USAGE:  getdefs [ <option-name>[{=| }<val>] ]...
Arg Option-Name    Description
Str defs-to-get    Regexp to look for after the "/*="
opt ordering       Alphabetize or use named file
                             - disabled as --no-ordering
                             - enabled by default
Num first-index    The first index to apply to groups
Str input          Input file to search for defs
                             - may appear multiple times
                             - default option for unnamed options
Str subblock       subblock definition names
                             - may appear multiple times
Str listattr       attribute with list of values
                             - may appear multiple times
opt filelist       Insert source file names into defs
Definition insertion options
Arg Option-Name    Description
Str assign         Global assignments
                             - may appear multiple times
Str common-assign  Assignments common to all blocks
                             - may appear multiple times
Str copy           File(s) to copy into definitions
                             - may appear multiple times
opt srcfile        Insert source file name into each def
opt linenum        Insert source line number into each def
Definition output disposition options:
Arg Option-Name    Description
Str output         Output file to open
                             - an alternate for autogen
opt autogen        Invoke AutoGen with defs
                             - disabled as --no-autogen
                             - enabled by default
Str template       Template Name
Str agarg          AutoGen Argument
                             - prohibits these options:
                             output
                             - may appear multiple times
Str base-name      Base name for output file(s)
                             - prohibits these options:

                                output
version and help options:
   Arg Option-Name    Description
   opt version        Output version information and exit
   no  help           Display usage information and exit
   no  more-help      Extended usage information passed thru pager
   opt save-opts      Save the option state to a config file
   Str load-opts      Load options from a config file
                                - disabled as --no-load-opts
                                - may appear multiple times
All arguments are named options.
If no ``input'' argument is provided or is set to simply "-", and if
``stdin'' is not a ``tty'', then the list of input files will be
read from ``stdin''.
The following option preset mechanisms are supported:
 - reading file /dev/null
This program extracts AutoGen definitions from a list of source files.
Definitions are delimited by `/*=<entry-type> <entry-name>\n' and
`=*/\n'.  From that, this program creates a definition of the
following form:
    #line nnn "source-file-name"
    entry_type = {
        name = entry_name;
        ...
    };
The ellipsis '...' is filled in by text found between the two
delimiters, with everything up through the first sequence of
asterisks deleted on every line.
There are two special ``entry types'':
*  The entry_type enclosure and the name entry will be omitted
   and the ellipsis will become top-level definitions.
-- The contents of the comment must be a single getdefs option.
   The option name must follow the double hyphen and its argument
   will be everything following the name.  This is intended for use
   with the ``subblock'' and ``listattr'' options.

please send bug reports to:  autogen-users@lists.sourceforge.net

8.5.2 defs-to-get option

This is the “regexp to look for after the "/*="” option. If you want definitions only from a particular category, or even with names matching particular patterns, then specify this regular expression for the text that must follow the /*=.

8.5.3 ordering option

This is the “alphabetize or use named file” option.

This option has some usage constraints. It:

• is enabled by default.

By default, ordering is alphabetical by the entry name. Use, no-ordering if order is unimportant. Use ordering with no argument to order without case sensitivity. Use ordering=<file-name> if chronological order is important. getdefs will maintain the text content of file-name. file-name need not exist.

8.5.4 first-index option

This is the “the first index to apply to groups” option. By default, the first occurrence of a named definition will have an index of zero. Sometimes, that needs to be a reserved value. Provide this option to specify a different starting point.

8.5.5 input option

This is the “input file to search for defs” option.

This option has some usage constraints. It:

• may appear an unlimited number of times.

All files that are to be searched for definitions must be named on the command line or read from stdin. If there is only one input option and it is the string, "-", then the input file list is read from stdin. If a command line argument is not an option name and does not contain an assignment operator (=), then it defaults to being an input file name. At least one input file must be specified.

8.5.6 subblock option

This is the “subblock definition names” option.

This option has some usage constraints. It:

• may appear an unlimited number of times.

This option is used to create shorthand entries for nested definitions. For example, with:

using subblock thus

          --subblock=arg=argname,type,null

and defining an arg thus

          arg: this, char *

will then expand to:

          arg = { argname = this; type = "char *"; };

The "this, char *" string is separated at the commas, with the white space removed. You may use characters other than commas by starting the value string with a punctuation character other than a single or double quote character. You may also omit intermediate values by placing the commas next to each other with no intervening white space. For example, "+mumble++yes+" will expand to:

arg = { argname = mumble; null = "yes"; };.

8.5.7 listattr option

This is the “attribute with list of values” option.

This option has some usage constraints. It:

• may appear an unlimited number of times.

This option is used to create shorthand entries for definitions that generally appear several times. That is, they tend to be a list of values. For example, with: listattr=foo defined, the text:

foo: this, is, a, multi-list will then expand to:
foo = 'this', 'is', 'a', 'multi-list';

The texts are separated by the commas, with the white space removed. You may use characters other than commas by starting the value string with a punctuation character other than a single or double quote character.

8.5.8 filelist option

This is the “insert source file names into defs” option. Inserts the name of each input file into the output definitions. If no argument is supplied, the format will be:

        infile = '%s';

If an argument is supplied, that string will be used for the entry name instead of infile.

8.5.9 assign option

This is the “global assignments” option.

This option has some usage constraints. It:

• may appear an unlimited number of times.

The argument to each copy of this option will be inserted into the output definitions, with only a semicolon attached.

8.5.10 common-assign option

This is the “assignments common to all blocks” option.

This option has some usage constraints. It:

• may appear an unlimited number of times.

The argument to each copy of this option will be inserted into each output definition, with only a semicolon attached.

8.5.11 copy option

This is the “file(s) to copy into definitions” option.

This option has some usage constraints. It:

• may appear an unlimited number of times.

The content of each file named by these options will be inserted into the output definitions.

8.5.12 srcfile option

This is the “insert source file name into each def” option. Inserts the name of the input file where a definition was found into the output definition. If no argument is supplied, the format will be:

        srcfile = '%s';

If an argument is supplied, that string will be used for the entry name instead of srcfile.

8.5.13 linenum option

This is the “insert source line number into each def” option. Inserts the line number in the input file where a definition was found into the output definition. If no argument is supplied, the format will be:

        linenum = '%s';

If an argument is supplied, that string will be used for the entry name instead of linenum.

8.5.14 output option

This is the “output file to open” option.

This option has some usage constraints. It:

• is a member of the autogen class of options.

If you are not sending the output to an AutoGen process, you may name an output file instead.

8.5.15 autogen option

This is the “invoke autogen with defs” option.

This option has some usage constraints. It:

• is enabled by default.

• is a member of the autogen class of options.

This is the default output mode. Specifying no-autogen is equivalent to output=-. If you supply an argument to this option, that program will be started as if it were AutoGen and its standard in will be set to the output definitions of this program.

8.5.16 template option

This is the “template name” option. Specifies the template name to be used for generating the final output.

8.5.17 agarg option

This is the “autogen argument” option.

This option has some usage constraints. It:

• may appear an unlimited number of times.

• must not appear in combination with any of the following options: output.

This is a pass-through argument. It allows you to specify any arbitrary argument to be passed to AutoGen.

8.5.18 base-name option

This is the “base name for output file(s)” option.

This option has some usage constraints. It:

• must not appear in combination with any of the following options: output.

When output is going to AutoGen, a base name must either be supplied or derived. If this option is not supplied, then it is taken from the template option. If that is not provided either, then it is set to the base name of the current directory.

8.6 Invoking xml2ag

This program will convert any arbitrary XML file into equivalent AutoGen definitions, and invoke AutoGen. The template used will be derived from either:

• The –override-tpl command line option

• A top level XML attribute named, "template"

One or the other must be provided, or the program will exit with a failure message. The base-name for the output will similarly be either:

• The –base-name command line option.

• The base name of the ‘.xml’ file.

The definitions derived from XML generally have an extra layer of definition. Specifically, this XML input:

        <mumble attr="foo">
          mumble-1
          <grumble>
          grumble, grumble, grumble.
        </grumble>mumble, mumble
        </mumble>

Will get converted into this:

        mumble = {
          grumble = {
            text = 'grumble, grumble, grumble';
          };
          text = 'mumble-1';
          text = 'mumble, mumble';
        };

Please notice that some information is lost. AutoGen cannot tell that "grumble" used to lie between the mumble texts. Also please note that you cannot assign:

        grumble = 'grumble, grumble, grumble.';

because if another "grumble" has an attribute or multiple texts, it becomes impossible to have the definitions be the same type (compound or text values).

This section was generated by AutoGen, the aginfo template and the option descriptions for the xml2ag program. It documents the xml2ag usage text and option meanings.

This software is released under the GNU General Public License.

8.6.1 xml2ag usage help (-?)

This is the automatically generated usage text for xml2ag:

xml2ag (GNU AutoGen) - XML to AutoGen Definiton Converter - Ver. 5.8.6
USAGE:  xml2ag [ -<flag> [<val>] | --<name>[{=| }<val>] ]... [ <def-file> ]
  Flg Arg Option-Name    Description
   -O Str output         Output file in lieu of AutoGen processing
   -L Str templ-dirs     Template search directory list
                                - may appear multiple times

-T Str override-tpl   Override template file
-l Str lib-template   Library template file
                             - may appear multiple times
-b Str base-name      Base name for output file(s)
   Str definitions    Definitions input file
-S Str load-scheme    Scheme code file to load
-F Str load-functions Load scheme function library
-s Str skip-suffix    Omit the file with this suffix
                             - may appear multiple times
-o opt select-suffix  specify this output suffix
                             - may appear multiple times
   no  source-time    set mod times to latest source
-m no  no-fmemopen    Do not use in-mem streams
   Str equate         characters considered equivalent
   no  writable       Allow output files to be writable
                             - disabled as --not-writable
   Num loop-limit     Limit on increment loops
                               it must lie in one of the ranges:
                               -1 exactly, or
                               1 to 16777216
-t Num timeout        Time limit for servers
                               it must lie in the range: 0 to 3600
   KWd trace          tracing level of detail
   Str trace-out      tracing output file or filter
   no  show-defs      Show the definition tree
-D Str define         name to add to definition list
                             - may appear multiple times
-U Str undefine       definition list removal pattern
                             - an alternate for define
-v opt version        Output version information and exit
-? no  help           Display usage information and exit
-! no  more-help      Extended usage information passed thru pager
Options are specified by doubled hyphens and their name
or by a single hyphen and the flag character.
This program will convert any arbitrary XML file into equivalent
AutoGen definitions, and invoke AutoGen.
The valid "trace" option keywords are:
nothing server-shell templates block-macros expressions everything
The template will be derived from either:
*  the ``--override-tpl'' command line option
*  a top level XML attribute named, "template"
The ``base-name'' for the output will similarly be either:

*  the ``--base-name'' command line option
*  the base name of the .xml file
please send bug reports to:  autogen-users@lists.sourceforge.net

8.6.2 output option (-O)

This is the “output file in lieu of autogen processing” option. By default, the output is handed to an AutoGen for processing. However, you may save the definitions to a file instead.

8.6.3 templ-dirs option (-L)

This is the “template search directory list” option.

This option has some usage constraints. It:

• may appear an unlimited number of times.

Pass-through AutoGen argument

8.6.4 override-tpl option (-T)

This is the “override template file” option. Pass-through AutoGen argument

8.6.5 lib-template option (-l)

This is the “library template file” option.

This option has some usage constraints. It:

• may appear an unlimited number of times.

Pass-through AutoGen argument

8.6.6 base-name option (-b)

This is the “base name for output file(s)” option. Pass-through AutoGen argument

8.6.7 definitions option

This is the “definitions input file” option. Pass-through AutoGen argument

8.6.8 load-scheme option (-S)

This is the “scheme code file to load” option. Pass-through AutoGen argument

8.6.9 load-functions option (-F)

This is the “load scheme function library” option.

This option has some usage constraints. It:

• must be compiled in by defining HAVE_DLOPEN during the compilation.

Pass-through AutoGen argument

8.6.10 skip-suffix option (-s)

This is the “omit the file with this suffix” option.

This option has some usage constraints. It:

• may appear an unlimited number of times.

Pass-through AutoGen argument

8.6.11 select-suffix option (-o)

This is the “specify this output suffix” option.

This option has some usage constraints. It:

• may appear an unlimited number of times.

Pass-through AutoGen argument

8.6.12 source-time option

This is the “set mod times to latest source” option. Pass-through AutoGen argument

8.6.13 no-fmemopen option (-m)

This is the “do not use in-mem streams” option.

This option has some usage constraints. It:

• must be compiled in by defining ENABLE_FMEMOPEN during the compilation.

Pass-through AutoGen argument

8.6.14 equate option

This is the “characters considered equivalent” option. Pass-through AutoGen argument

8.6.15 writable option

This is the “allow output files to be writable” option. Pass-through AutoGen argument

8.6.16 loop-limit option

This is the “limit on increment loops” option. Pass-through AutoGen argument

8.6.17 timeout option (-t)

This is the “time limit for servers” option. Pass-through AutoGen argument

8.6.18 trace option

This is the “tracing level of detail” option.

This option has some usage constraints. It:

• This option takes a keyword as its argument. The argument sets an enumeration value that can be tested by comparing the option value macro (OPT VALUE TRACE). The available keywords are:

           nothing      server-shell templates
           block-macros expressions  everything

Pass-through AutoGen argument

8.6.19 trace-out option

This is the “tracing output file or filter” option. Pass-through AutoGen argument

8.6.20 show-defs option

This is the “show the definition tree” option. Pass-through AutoGen argument

8.6.21 define option (-D)

This is the “name to add to definition list” option.

This option has some usage constraints. It:

• may appear an unlimited number of times.

Pass-through AutoGen argument

8.6.22 undefine option (-U)

This is the “definition list removal pattern” option.

This option has some usage constraints. It:

• may appear an unlimited number of times.

Pass-through AutoGen argument

8.7 Replacement for Stdio Formatting Library

Using the ‘printf’ formatting routines in a portable fashion has always been a pain, and this package has been way more pain than anyone ever imagined. Hopefully, with this release of snprintfv, the pain is now over for all time.

The issues with portable usage are these:

1. Argument number specifiers are often either not implemented or are buggy. Even GNU libc, version 1 got it wrong.

2. ANSI/ISO "forgot" to provide a mechanism for computing argument lists for vararg procedures.

3. The argument array version of printf (‘printfv()’) is not generally available, does not work with the native printf, and does not have a working argument number specifier in the format specification. (Last I knew, anyway.)

4. You cannot fake varargs by calling ‘vprintf()’ with an array of arguments, because ANSI does not require such an implementation and some vendors play funny tricks because they are allowed to.

These four issues made it impossible for AutoGen to ship without its own implementation of the ‘printf’ formatting routines. Since we were forced to do this, we decided to make the formatting routines both better and more complete :-). We addressed these issues and added the following features to the common printf API:

5. The formatted output can be written to • a string allocated by the formatting function (‘asprintf()’).

• a file descriptor instead of a file stream (‘dprintf()’).

• a user specified stream (‘stream printf()’).

6. The formatting functions can be augmented with your own functions. These functions are allowed to consume more than one character from the format, but must commence with a unique character. For example,

        "%{struct stat}\n"

might be used with ’{’ registered to a procedure that would look up "struct stat" in a symbol table and do appropriate things, consuming the format string through the ’}’ character.

Gary V. Vaughan was generous enough to supply this implementation. Many thanks!! For further details, the reader is referred to the snprintfv documentation. These functions

are also available in the template processing as ‘sprintf’ (see Section 3.5.29 [SCM sprintf], page 39), ‘printf’ (see Section 3.5.24 [SCM printf], page 37), ‘fprintf’ (see Section 3.5.9 [SCM fprintf], page 33), and ‘shellf’ (see Section 3.5.28 [SCM shellf], page 39).

Chapter 9: Some ideas for the future. 158

9 Some ideas for the future.

Here are some things that might happen in the distant future.

• Fix up current tools that contain miserably complex perl, shell, sed, awk and m4 scripts to instead use this tool.

Concept Index 159

Concept Index

#

#assert . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 #define . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 #elif. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 #else . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 #endif . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 #endmac . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 #endshell . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 #error . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 #if . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 #ifdef . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 #ifndef . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 #include . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 #line . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 #macdef . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 #option. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 #shell . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 #undef . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12

.

.def file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 .tpl file. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17

A

allow-errors . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 Alternate Definition. . . . . . . . . . . . . . . . . . . . . . . . . . 15 arg-default . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 88 arg-optional . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 88 arg-range . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 88 arg-type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86 argument . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 assert directive . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 Augmenting AutoGen . . . . . . . . . . . . . . . . . . . . . . . . 54 AutoEvents . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 141 AutoFSM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 141 AutoGen Definition Extraction Tool. . . . . . . . . . 145 autogen usage . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 autogen-base-name. . . . . . . . . . . . . . . . . . . . . . . . . . . 59 autogen-define . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 autogen-definitions. . . . . . . . . . . . . . . . . . . . . . . . . . . 59 autogen-equate . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 autogen-lib-template . . . . . . . . . . . . . . . . . . . . . . . . . 58 autogen-load-functions . . . . . . . . . . . . . . . . . . . . . . . 59 autogen-load-scheme . . . . . . . . . . . . . . . . . . . . . . . . . 59 autogen-loop-limit . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 autogen-no-fmemopen . . . . . . . . . . . . . . . . . . . . . . . . 61 autogen-override-tpl. . . . . . . . . . . . . . . . . . . . . . . . . . 58 autogen-select-suffix. . . . . . . . . . . . . . . . . . . . . . . . . . 60 autogen-show-defs . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 autogen-skip-suffix . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 autogen-source-time. . . . . . . . . . . . . . . . . . . . . . . . . . 60 autogen-templ-dirs . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 autogen-timeout . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 autogen-trace. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 autogen-trace-out . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 autogen-undefine . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 autogen-writable. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 AutoInfo . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 135 AutoMan pages. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 137 automatic options . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 autoopts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 AutoOpts API. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94 AutoXDR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 141

C

call-proc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 Columnize Input Text. . . . . . . . . . . . . . . . . . . . . . . 142 columns usage . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 142 columns-by-columns . . . . . . . . . . . . . . . . . . . . . . . . 144 columns-col-width . . . . . . . . . . . . . . . . . . . . . . . . . . 143 columns-columns . . . . . . . . . . . . . . . . . . . . . . . . . . . 143 columns-first-indent . . . . . . . . . . . . . . . . . . . . . . . . . 143 columns-format . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 144 columns-indent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 143 columns-input . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 144 columns-line-separation . . . . . . . . . . . . . . . . . . . . . 144 columns-separation . . . . . . . . . . . . . . . . . . . . . . . . . 144 columns-sort . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 144 columns-spread . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 143 columns-tab-width . . . . . . . . . . . . . . . . . . . . . . . . . . 144 columns-width. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 143 comments . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 Common Option Attributes . . . . . . . . . . . . . . . . . . 83 compound definitions . . . . . . . . . . . . . . . . . . . . . . . . . 7 concat-string . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 conditional emit . . . . . . . . . . . . . . . . . . . . . . . . . . 50, 52 config-header . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 configuration file . . . . . . . . . . . . . . 74, 84, 91, 94, 118 Configuration File . . . . . . . . . . . . . . . . . . . . . . 115, 116 Configuration File example . . . . . . . . . . . . . . . . . . 115 configuring . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 copyright. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81

D

default . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 88 define directive . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 define macro . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 Definition Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 definitions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 definitions file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 descrip . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83 design goals . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 detail . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 directives . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 disable . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83 diversion . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 documentation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 documentation attributes. . . . . . . . . . . . . . . . . . . . . 90 Dynamic Definition Text . . . . . . . . . . . . . . . . . . . . . 10

E

elif directive . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 else directive . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 enable . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83 enabled . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83 endif directive . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 endmac directive. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 endshell directive . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 environrc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74, 114 equivalence . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84 error directive . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 example, simple AutoGen . . . . . . . . . . . . . . . . . . . . . 2 example, simple AutoOpts. . . . . . . . . . . . . . . . . . . . 71 explain. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 export . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 expression syntax . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 extract-code . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89

F

features . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 finite state machine . . . . . . . . . . . . . . . . . . . . . . . . . 141 flag-code . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 flag-proc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 flags-cant . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86 flags-must . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86 fOptState . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94 for loop . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 futures . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 158

G

getdefs usage . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 146 getdefs-agarg . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 150 getdefs-assign . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 149 getdefs-autogen . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 150 getdefs-base-name . . . . . . . . . . . . . . . . . . . . . . . . . . 151 getdefs-common-assign . . . . . . . . . . . . . . . . . . . . . . 149 getdefs-copy . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 149 getdefs-defs-to-get . . . . . . . . . . . . . . . . . . . . . . . . . . 148 getdefs-filelist . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 149 getdefs-first-index. . . . . . . . . . . . . . . . . . . . . . . . . . . 148 getdefs-input . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 148 getdefs-linenum. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 150 getdefs-listattr. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 149 getdefs-ordering . . . . . . . . . . . . . . . . . . . . . . . . . . . . 148 getdefs-output . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 150 getdefs-srcfile . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 150 getdefs-subblock . . . . . . . . . . . . . . . . . . . . . . . . . . . . 148 getdefs-template . . . . . . . . . . . . . . . . . . . . . . . . . . . . 150 getopt long . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 138 gnu-usage . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82 guard-option-names . . . . . . . . . . . . . . . . . . . . . . . . . . 74

H

here-string . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 homerc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74

I

identification . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 if directive . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 if test . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 ifdef directive. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 ifndef directive . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 immed-disable . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 immediate. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 immediate action . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 include. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 include directive . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 information attributes. . . . . . . . . . . . . . . . . . . . . . . . 80 Installing. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 Internationalizing AutoOpts . . . . . . . . . . . . . . . . . 139 Introduction. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

K

keyword. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 88

L

library attributes . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 Licensing. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 line directive . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 long-opts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 looping, for. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49

M

m4 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 macdef directive . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 macro syntax . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 macro, pseudo . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 main procedure . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 man-doc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 max . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83 min. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83 must-set . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83

N

name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82 named option mode . . . . . . . . . . . . . . . . . . . . . . . . . . 74 Naming Conflicts . . . . . . . . . . . . . . . . . . . . . . . . . . . 140 naming values . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 native macros . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 no-preset . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84

O

optActualIndex . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94 optActualValue . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94 optIndex . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94 Option Argument Handling. . . . . . . . . . . . . . . . . . . 89 option argument name . . . . . . . . . . . . . . . . . . . . . . . 90 Option Arguments . . . . . . . . . . . . . . . . . . . . . . . . . . . 86 option attributes. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82 Option Conflict Attributes . . . . . . . . . . . . . . . . . . . 86 Option Definitions . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 option descriptor . . . . . . . . . . . . . . . . . . . . . . . . . . . 111 option directive . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 option documentation . . . . . . . . . . . . . . . . . . . . . . . . 90 Option Processing Data . . . . . . . . . . . . . . . . . . . . . . 94 optOccCt . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94 opts-ptr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 optValue . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94

P

package . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 predefines . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 prefix . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 preserve-case . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 prog-desc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 prog-info-descrip . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 prog-man-descrip . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 prog-name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 prog-title. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 program attributes. . . . . . . . . . . . . . . . . . . . . . . . . . . 73 pseudo macro . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 pzLastArg. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95 pzProgName . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95 pzProgPath . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95

R

rcfile . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74, 114, 115 Redirecting Output . . . . . . . . . . . . . . . . . . . . . . . . . . 52 remote procedure call . . . . . . . . . . . . . . . . . . . . . . . 141 reorder-args . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82 Required Attributes. . . . . . . . . . . . . . . . . . . . . . . . . . 82 RPC. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 141 rpcgen . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 141

S

sample rcfile. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 114 sectioned config file . . . . . . . . . . . . . . . . . . . . . . . . . 117 settable . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84 shell directive . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 shell options . . . . . . . . . . . . . . . . . . . . . . . . . . . 113, 118 shell-generated string . . . . . . . . . . . . . . . . . . . . . . . . . 8 Signal Names. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 simple definitions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 Special Option Handling . . . . . . . . . . . . . . . . . . . . . 84 stack-arg . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 standard options. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 string, double quote . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 string, shell output. . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 string, single quote. . . . . . . . . . . . . . . . . . . . . . . . . . . . 7

T

template file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6, 17 The Automated Program Generator . . . . . . . . . . . 56

U

undef directive . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 unstack-arg . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 usage . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82 using AutoOpts . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111

V

value . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83 version . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74

W

while test . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52

X

XDR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 141 XML to AutoGen Definiton Converter. . . . . . . . 152 xml2ag usage . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 152 xml2ag-base-name . . . . . . . . . . . . . . . . . . . . . . . . . . 154 xml2ag-define . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 156 xml2ag-definitions . . . . . . . . . . . . . . . . . . . . . . . . . . 154 xml2ag-equate . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 155 xml2ag-lib-template. . . . . . . . . . . . . . . . . . . . . . . . . 154 xml2ag-load-functions . . . . . . . . . . . . . . . . . . . . . . . 154 xml2ag-load-scheme. . . . . . . . . . . . . . . . . . . . . . . . . 154 xml2ag-loop-limit . . . . . . . . . . . . . . . . . . . . . . . . . . . 155 xml2ag-no-fmemopen . . . . . . . . . . . . . . . . . . . . . . . 155 xml2ag-output. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 154 xml2ag-override-tpl . . . . . . . . . . . . . . . . . . . . . . . . . 154 xml2ag-select-suffix . . . . . . . . . . . . . . . . . . . . . . . . . 155 xml2ag-show-defs . . . . . . . . . . . . . . . . . . . . . . . . . . . 156 xml2ag-skip-suffix . . . . . . . . . . . . . . . . . . . . . . . . . . 155 xml2ag-source-time . . . . . . . . . . . . . . . . . . . . . . . . . 155 xml2ag-templ-dirs . . . . . . . . . . . . . . . . . . . . . . . . . . 154 xml2ag-timeout. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 155 xml2ag-trace . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 155 xml2ag-trace-out . . . . . . . . . . . . . . . . . . . . . . . . . . . 156 xml2ag-undefine . . . . . . . . . . . . . . . . . . . . . . . . . . . . 156 xml2ag-writable . . . . . . . . . . . . . . . . . . . . . . . . . . . . 155

Function Index 162

Function Index

*

*= . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 *=* . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 *== . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 *==* . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 *~ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 *~* . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 *~~ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 *~~* . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41

=

= . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 =* . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 == . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 ==* . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42

~

~ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 ~* . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 ~~ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 ~~* . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42

A

ag-fprintf . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 ag-function? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 ao_string_tokenize. . . . . . . . . . . . . . . . . . . . . . . . 101 autogen-version . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30

B

base-name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 bsd . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31

C

c-file-line-fmt . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 c-string. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 CASE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 chdir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 CLEAR_OPT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 96 COMMENT. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 configFileLoad . . . . . . . . . . . . . . . . . . . . . . . . . . . . 102 count . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 COUNT_OPT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 96

D

def-file. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 def-file-line . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 DEFINE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 DESC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 96 DISABLE_OPT_name . . . . . . . . . . . . . . . . . . . . . . . . . . . 96 dne . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23

E

ELIF . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 ELSE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 emit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 emit-string-table . . . . . . . . . . . . . . . . . . . . . . . . . . 32 ENABLED_OPT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 96 ENDDEF . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 ENDFOR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 ENDIF . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 ENDWHILE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 error . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 error-source-line . . . . . . . . . . . . . . . . . . . . . . . . . . 32 ERRSKIP_OPTERR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 96 ERRSTOP_OPTERR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 96 ESAC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 exist? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 EXPR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 extract. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32

F

find-file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 first-for? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 FOR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 for-by . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 for-from. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 for-index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 for-sep. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 for-to . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 format-arg-count . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 fprintf. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33

G

get . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 gperf . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 gpl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34

H

HAVE_OPT. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 96 hide-email . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 high-lim. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 html-escape-encode . . . . . . . . . . . . . . . . . . . . . . . . . 34

I

IF . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 in? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 INCLUDE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 INVOKE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 ISSEL_OPT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97 ISUNUSED_OPT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97

J

join . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35

K

kr-string . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35

L

last-for? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 len . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 lgpl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 license. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 low-lim. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26

M

make-gperf . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 make-header-guard . . . . . . . . . . . . . . . . . . . . . . . . . . 26 makefile-script . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 match-value? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 max . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 min . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37

O

OPT_ARG. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97 OPT_VALUE_name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97 OPTION_CT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97 optionFileLoad . . . . . . . . . . . . . . . . . . . . . . . . . . . . 102 optionFindNextValue. . . . . . . . . . . . . . . . . . . . . . . 103 optionFindValue . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 optionFree . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104 optionGetValue . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104 optionLoadLine . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104 optionNextValue . . . . . . . . . . . . . . . . . . . . . . . . . . . 105 optionOnlyUsage . . . . . . . . . . . . . . . . . . . . . . . . . . . 105 optionProcess . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 106 optionRestore . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 106 optionSaveFile . . . . . . . . . . . . . . . . . . . . . . . . . . . . 107 optionSaveState . . . . . . . . . . . . . . . . . . . . . . . . . . . 107 optionUnloadNested. . . . . . . . . . . . . . . . . . . . . . . . 107 optionVersion . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108 out-delete . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 out-depth . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 out-line. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 out-move. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 out-name. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 out-pop. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 out-push-add . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 out-push-new . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 out-resume . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 out-suspend . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 out-switch . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29

P

pathfind . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108 prefix . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 printf . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37

R

raw-shell-str . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 RESTART_OPT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97

S

SELECT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 set-option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 set-writable . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 SET_OPT_name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97 shell . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 shell-str . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 shellf . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 sprintf. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 stack . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 STACKCT_OPT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 98 STACKLST_OPT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 98 START_OPT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 98 STATE_OPT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 98 strequate . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 streqvcmp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 streqvmap . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 string->c-name! . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 string-capitalize . . . . . . . . . . . . . . . . . . . . . . . . . . 39 string-capitalize! . . . . . . . . . . . . . . . . . . . . . . . . . 39 string-contains-eqv?. . . . . . . . . . . . . . . . . . . . . . . 39 string-contains? . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 string-downcase . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 string-downcase! . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 string-end-eqv-match? . . . . . . . . . . . . . . . . . . . . . 40 string-end-match? . . . . . . . . . . . . . . . . . . . . . . . . . . 40 string-ends-eqv? . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 string-ends-with? . . . . . . . . . . . . . . . . . . . . . . . . . . 40 string-equals? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 string-eqv-match? . . . . . . . . . . . . . . . . . . . . . . . . . . 41 string-eqv? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 string-has-eqv-match? . . . . . . . . . . . . . . . . . . . . . 41 string-has-match? . . . . . . . . . . . . . . . . . . . . . . . . . . 41 string-match? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42

string-start-eqv-match? . . . . . . . . . . . . . . . . . . . 42

string-start-match?. . . . . . . . . . . . . . . . . . . . . . . . 42 string-starts-eqv? . . . . . . . . . . . . . . . . . . . . . . . . . 42 string-starts-with?. . . . . . . . . . . . . . . . . . . . . . . . 42 string-substitute . . . . . . . . . . . . . . . . . . . . . . . . . . 42 string-table-add . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 string-table-new . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 string-tr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 string-tr! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 string-upcase . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 string-upcase! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 strneqvcmp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 110 strtransform . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 110 sub-shell-str . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 suffix . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 sum . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45

T

teOptIndex . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 100 tpl-file. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 tpl-file-line . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30

U

UNKNOWN. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 USAGE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99

V

VALUE_OPT_name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99 VERSION. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99 version-compare . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45

W

WHICH_IDX_name . . . . . . . . . . . . . . . . . . . . . . . . . . . . 100 WHICH_OPT_name . . . . . . . . . . . . . . . . . . . . . . . . . . . . 100 WHILE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52

Table of Contents

1 Introduction. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

1.1 The Purpose of AutoGen . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 1.2 A Simple Example . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 1.3 csh/zsh caveat . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 1.4 A User’s Perspective. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

2 Definitions File . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

2.1 The Identification Definition . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 2.2 Named Definitions. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7

2.2.1 Definition List . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 2.2.2 Double Quote String . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 2.2.3 Single Quote String . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 2.2.4 Shell Output String . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 2.2.5 An Unquoted String . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 2.2.6 Scheme Result String . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 2.2.7 A Here String . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 2.2.8 Concatenated Strings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9

2.3 Assigning an Index to a Definition . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 2.4 Dynamic Text. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 2.5 Controlling What Gets Processed. . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 2.6 Pre-defined Names . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 2.7 Commenting Your Definitions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 2.8 What it all looks like.. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 2.9 Finite State Machine Grammar. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 2.10 Alternate Definition Forms. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15

3 Template File . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17

3.1 Format of the Pseudo Macro . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 3.2 Naming a value . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 3.3 Macro Expression Syntax . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19

3.3.1 Apply Code . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 3.3.2 Basic Expression. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20

3.4 AutoGen Scheme Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 3.4.1 ‘ag-function?’ - test for function. . . . . . . . . . . . . . . . . . . . . . . 22 3.4.2 ‘base-name’ - base output name . . . . . . . . . . . . . . . . . . . . . . . . 22 3.4.3 ‘chdir’ - Change current directory . . . . . . . . . . . . . . . . . . . . . . 22 3.4.4 ‘count’ - definition count . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 3.4.5 ‘def-file’ - definitions file name . . . . . . . . . . . . . . . . . . . . . . . 22 3.4.6 ‘def-file-line’ - get a definition file+line number . . . . . . 23 3.4.7 ‘dne’ - "Do Not Edit" warning . . . . . . . . . . . . . . . . . . . . . . . . . . 23 3.4.8 ‘error’ - display message and exit . . . . . . . . . . . . . . . . . . . . . . 23 3.4.9 ‘exist?’ - test for value name . . . . . . . . . . . . . . . . . . . . . . . . . . 24 3.4.10 ‘find-file’ - locate a file in the search path . . . . . . . . . . . . 24 3.4.11 ‘first-for?’ - detect first iteration . . . . . . . . . . . . . . . . . . . . 24 3.4.12 ‘for-by’ - set iteration step. . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 3.4.13 ‘for-from’ - set initial index. . . . . . . . . . . . . . . . . . . . . . . . . . . 24 3.4.14 ‘for-index’ - get current loop index . . . . . . . . . . . . . . . . . . . 25 3.4.15 ‘for-sep’ - set loop separation string. . . . . . . . . . . . . . . . . . . 25 3.4.16 ‘for-to’ - set ending index . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 3.4.17 ‘get’ - get named value . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 3.4.18 ‘high-lim’ - get highest value index. . . . . . . . . . . . . . . . . . . . 25 3.4.19 ‘last-for?’ - detect last iteration. . . . . . . . . . . . . . . . . . . . . . 26 3.4.20 ‘len’ - get count of values . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 3.4.21 ‘low-lim’ - get lowest value index . . . . . . . . . . . . . . . . . . . . . . 26 3.4.22 ‘make-header-guard’ - make self-inclusion guard. . . . . . . . 26 3.4.23 ‘match-value?’ - test for matching value . . . . . . . . . . . . . . . 27 3.4.24 ‘out-delete’ - delete current output file. . . . . . . . . . . . . . . . 27 3.4.25 ‘out-depth’ - output file stack depth . . . . . . . . . . . . . . . . . . . 27 3.4.26 ‘out-line’ - output file line number. . . . . . . . . . . . . . . . . . . . 27 3.4.27 ‘out-move’ - change name of output file . . . . . . . . . . . . . . . . 27 3.4.28 ‘out-name’ - current output file name . . . . . . . . . . . . . . . . . . 28 3.4.29 ‘out-pop’ - close current output file . . . . . . . . . . . . . . . . . . . . 28 3.4.30 ‘out-push-add’ - append output to file . . . . . . . . . . . . . . . . . 28 3.4.31 ‘out-push-new’ - purge and create output file. . . . . . . . . . . 28 3.4.32 ‘out-resume’ - resume suspended output file . . . . . . . . . . . . 28 3.4.33 ‘out-suspend’ - suspend current output file. . . . . . . . . . . . . 29 3.4.34 ‘out-switch’ - close and create new output . . . . . . . . . . . . . 29 3.4.35 ‘set-option’ - Set a command line option . . . . . . . . . . . . . . 29 3.4.36 ‘set-writable’ - Make the output file be writable . . . . . . 29 3.4.37 ‘stack’ - make list of AutoGen values . . . . . . . . . . . . . . . . . . 29 3.4.38 ‘suffix’ - get the current suffix . . . . . . . . . . . . . . . . . . . . . . . . 29 3.4.39 ‘tpl-file’ - get the template file name. . . . . . . . . . . . . . . . . 30 3.4.40 ‘tpl-file-line’ - get the template file+line number …. 30 3.4.41 ‘autogen-version’ - autogen version number . . . . . . . . . . . 30 3.4.42 format file info as, “#line nn "file"”. . . . . . . . . . . . . . . . . . 30

3.5 Common Scheme Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 3.5.1 ‘ag-fprintf’ - format to autogen stream . . . . . . . . . . . . . . . . 31 3.5.2 ‘bsd’ - BSD Public License . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 3.5.3 ‘c-string’ - emit string for ANSI C. . . . . . . . . . . . . . . . . . . . . 31 3.5.4 ‘emit’ - emit the text for each argument . . . . . . . . . . . . . . . . . 31 3.5.5 ‘emit-string-table’ - output a string table . . . . . . . . . . . . . 32 3.5.6 ‘error-source-line’ - display of file & line . . . . . . . . . . . . . 32 3.5.7 ‘extract’ - extract text from another file . . . . . . . . . . . . . . . . 32 3.5.8 ‘format-arg-count’ - count the args to a format. . . . . . . . . 33 3.5.9 ‘fprintf’ - format to a file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 3.5.10 ‘gperf’ - perform a perfect hash function . . . . . . . . . . . . . . . 34 3.5.11 ‘gpl’ - GNU General Public License . . . . . . . . . . . . . . . . . . . . 34 3.5.12 ‘hide-email’ - convert eaddr to javascript . . . . . . . . . . . . . . 34 3.5.13 ‘html-escape-encode’ - encode html special characters. . 34 3.5.14 ‘in?’ - test for string in list . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 3.5.15 ‘join’ - join string list with separator . . . . . . . . . . . . . . . . . . 35 3.5.16 ‘kr-string’ - emit string for K&R C . . . . . . . . . . . . . . . . . . . 35 3.5.17 ‘lgpl’ - GNU Library General Public License . . . . . . . . . . . 35 3.5.18 ‘license’ - an arbitrary license . . . . . . . . . . . . . . . . . . . . . . . . 35 3.5.19 ‘make-gperf’ - build a perfect hash function program. … 35 3.5.20 ‘makefile-script’ - create makefile script. . . . . . . . . . . . . . 36 3.5.21 ‘max’ - maximum value in list . . . . . . . . . . . . . . . . . . . . . . . . . . 37 3.5.22 ‘min’ - minimum value in list . . . . . . . . . . . . . . . . . . . . . . . . . . 37 3.5.23 ‘prefix’ - prefix lines with a string. . . . . . . . . . . . . . . . . . . . . 37 3.5.24 ‘printf’ - format to stdout . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 3.5.25 ‘raw-shell-str’ - single quote shell string. . . . . . . . . . . . . . 37 3.5.26 ‘shell’ - invoke a shell script . . . . . . . . . . . . . . . . . . . . . . . . . . 38 3.5.27 ‘shell-str’ - double quote shell string . . . . . . . . . . . . . . . . . 38 3.5.28 ‘shellf’ - format a string, run shell . . . . . . . . . . . . . . . . . . . . 39 3.5.29 ‘sprintf’ - format a string . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 3.5.30 ‘string-capitalize’ - capitalize a new string . . . . . . . . . . 39 3.5.31 ‘string-capitalize!’ - capitalize a string . . . . . . . . . . . . . 39 3.5.32 ‘string-contains-eqv?’ - caseless substring . . . . . . . . . . . 39 3.5.33 ‘string-contains?’ - substring match . . . . . . . . . . . . . . . . . 40 3.5.34 ‘string-downcase’ - lower case a new string . . . . . . . . . . . . 40 3.5.35 ‘string-downcase!’ - make a string be lower case. . . . . . . 40 3.5.36 ‘string-end-eqv-match?’ - caseless regex ending . . . . . . . 40 3.5.37 ‘string-end-match?’ - regex match end. . . . . . . . . . . . . . . . 40 3.5.38 ‘string-ends-eqv?’ - caseless string ending . . . . . . . . . . . . 40 3.5.39 ‘string-ends-with?’ - string ending . . . . . . . . . . . . . . . . . . . 40 3.5.40 ‘string-equals?’ - string matching . . . . . . . . . . . . . . . . . . . . 41 3.5.41 ‘string-eqv-match?’ - caseless regex match . . . . . . . . . . . . 41 3.5.42 ‘string-eqv?’ - caseless string match . . . . . . . . . . . . . . . . . . 41 3.5.43 ‘string-has-eqv-match?’ - caseless regex contains. . . . . . 41 3.5.44 ‘string-has-match?’ - contained regex match . . . . . . . . . . 41 3.5.45 ‘string-match?’ - regex match . . . . . . . . . . . . . . . . . . . . . . . . 42 3.5.46 ‘string-start-eqv-match?’ - caseless regex start . . . . . . . 42 3.5.47 ‘string-start-match?’ - regex match start . . . . . . . . . . . . 42 3.5.48 ‘string-starts-eqv?’ - caseless string start . . . . . . . . . . . . 42 3.5.49 ‘string-starts-with?’ - string starting. . . . . . . . . . . . . . . . 42 3.5.50 ‘string-substitute’ - multiple global replacements. . . . . 42 3.5.51 ‘string-table-add’ - Add an entry to a string table …. 43 3.5.52 ‘string-table-new’ - create a string table. . . . . . . . . . . . . . 43 3.5.53 ‘string->c-name!’ - map non-name chars to underscore

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 3.5.54 ‘string-tr’ - convert characters with new result . . . . . . . . 44 3.5.55 ‘string-tr!’ - convert characters . . . . . . . . . . . . . . . . . . . . . . 44 3.5.56 ‘string-upcase’ - upper case a new string . . . . . . . . . . . . . 45 3.5.57 ‘string-upcase!’ - make a string be upper case . . . . . . . . 45 3.5.58 ‘sub-shell-str’ - back quoted (sub-)shell string . . . . . . . . 45 3.5.59 ‘sum’ - sum of values in list . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 3.5.60 ‘version-compare’ - compare two version numbers . . . . . . 45

3.6 AutoGen Native Macros . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 3.6.1 AutoGen Macro Syntax . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 3.6.2 CASE - Select one of several template blocks . . . . . . . . . . . . 47 3.6.3 COMMENT - A block of comment to be ignored . . . . . . . . . 48 3.6.4 DEFINE - Define a user AutoGen macro . . . . . . . . . . . . . . . . 48 3.6.5 ELIF - Alternate Conditional Template Block. . . . . . . . . . . . 49 3.6.6 ELSE - Alternate Template Block. . . . . . . . . . . . . . . . . . . . . . . 49 3.6.7 ENDDEF - Ends a macro definition. . . . . . . . . . . . . . . . . . . . . 49 3.6.8 ENDFOR - Terminates the FOR function template block . . 49 3.6.9 ENDIF - Terminate the IF Template Block . . . . . . . . . . . . . . 49 3.6.10 ENDWHILE - Terminate the WHILE Template Block. . . . . 49 3.6.11 ESAC - Terminate the CASE Template Block . . . . . . . . . . . . 49 3.6.12 EXPR - Evaluate and emit an Expression . . . . . . . . . . . . . . 49 3.6.13 FOR - Emit a template block multiple times . . . . . . . . . . . . 49 3.6.14 IF - Conditionally Emit a Template Block . . . . . . . . . . . . . . 50 3.6.15 INCLUDE - Read in and emit a template block . . . . . . . . . 51 3.6.16 INVOKE - Invoke a User Defined Macro. . . . . . . . . . . . . . . . 51 3.6.17 SELECT - Selection block for CASE function . . . . . . . . . . . 51 3.6.18 UNKNOWN - Either a user macro or a value name. . . . . . 51 3.6.19 WHILE - Conditionally loop over a Template Block . . . . . 52

3.7 Redirecting Output. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52

4 Augmenting AutoGen Features . . . . . . . . . . . . 54

4.1 Shell Output Commands. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 4.2 Guile Macros . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 4.3 Guile Callout Functions. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 4.4 AutoGen Macros . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55

5 Invoking autogen . . . . . . . . . . . . . . . . . . . . . . . . . 56

5.1 autogen usage help (-?) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 5.2 templ-dirs option (-L) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 5.3 override-tpl option (-T) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 5.4 lib-template option (-l) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 5.5 base-name option (-b) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 5.6 definitions option. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 5.7 load-scheme option (-S) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 5.8 load-functions option (-F) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 5.9 skip-suffix option (-s) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 5.10 select-suffix option (-o) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 5.11 source-time option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 5.12 no-fmemopen option (-m) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 5.13 equate option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 5.14 writable option. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 5.15 loop-limit option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 5.16 timeout option (-t) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 5.17 trace option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 5.18 trace-out option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 5.19 show-defs option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 5.20 define option (-D) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62

5.21 undefine option (-U). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63

6 Configuring and Installing . . . . . . . . . . . . . . . . . 64

6.1 Configuring AutoGen . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 6.2 AutoGen as a CGI server . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 6.3 Signal Names . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 6.4 Installing AutoGen . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66

7 Automated Option Processing . . . . . . . . . . . . . 68

7.1 AutoOpts Features . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 7.2 AutoOpts Licensing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 7.3 Quick Start . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 7.4 Multi-Threading . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 7.5 Option Definitions. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73

7.5.1 Program Description Attributes. . . . . . . . . . . . . . . . . . . . . . . . . 73 7.5.2 Options for Library Code. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75

7.5.2.1 AutoOpt-ed Library for AutoOpt-ed Program . . . . . . . 75 7.5.2.2 AutoOpt-ed Library for Regular Program . . . . . . . . . . . 76 7.5.2.3 AutoOpt-ed Program Calls Regular Library . . . . . . . . . 76

7.5.3 Generating main procedures . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 7.5.3.1 guile: main and inner main procedures. . . . . . . . . . . . . . 77 7.5.3.2 shell-process: emit Bourne shell results. . . . . . . . . . . . . . 77 7.5.3.3 shell-parser: emit Bourne shell script . . . . . . . . . . . . . . . 77 7.5.3.4 main: user supplied main procedure . . . . . . . . . . . . . . . . 78 7.5.3.5 include: code emitted from included template . . . . . . . 78 7.5.3.6 invoke: code emitted from AutoGen macro . . . . . . . . . . 78 7.5.3.7 for-each: perform function on each argument . . . . . . . . 78

7.5.4 Program Information Attributes . . . . . . . . . . . . . . . . . . . . . . . . 80 7.5.5 Option Attributes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82

7.5.5.1 Required Attributes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82 7.5.5.2 Common Option Attributes . . . . . . . . . . . . . . . . . . . . . . . . 83 7.5.5.3 Special Option Handling . . . . . . . . . . . . . . . . . . . . . . . . . . . 84 7.5.5.4 Immediate Action Attributes . . . . . . . . . . . . . . . . . . . . . . . 85 7.5.5.5 Option Conflict Attributes . . . . . . . . . . . . . . . . . . . . . . . . . 86 7.5.5.6 Option Argument Specification . . . . . . . . . . . . . . . . . . . . . 86 7.5.5.7 Option Argument Handling . . . . . . . . . . . . . . . . . . . . . . . . 89

7.5.6 Man and Info doc Attributes . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 7.5.7 Automatically Supported Options . . . . . . . . . . . . . . . . . . . . . . . 90 7.5.8 Library of Standard Options. . . . . . . . . . . . . . . . . . . . . . . . . . . . 92

7.6 Programmatic Interface . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94 7.6.1 Data for Option Processing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94 7.6.2 CLEAR OPT( <NAME> ) - Clear Option Markings . . . . . . 96 7.6.3 COUNT OPT( <NAME> ) - Definition Count . . . . . . . . . . . 96 7.6.4 DESC( <NAME> ) - Option Descriptor . . . . . . . . . . . . . . . . . . 96 7.6.5 DISABLE OPT name - Disable an option . . . . . . . . . . . . . . . 96 7.6.6 ENABLED OPT( <NAME> ) - Is Option Enabled? . . . . . . 96 7.6.7 ERRSKIP OPTERR - Ignore Option Errors . . . . . . . . . . . . . 96 7.6.8 ERRSTOP OPTERR - Stop on Errors . . . . . . . . . . . . . . . . . . 96 7.6.9 HAVE OPT( <NAME> ) - Have this option? . . . . . . . . . . . . 96 7.6.10 ISSEL OPT( <NAME> ) - Is Option Selected? . . . . . . . . . . 97 7.6.11 ISUNUSED OPT( <NAME> ) - Never Specified?. . . . . . . . 97 7.6.12 OPTION CT - Full Count of Options . . . . . . . . . . . . . . . . . . 97 7.6.13 OPT ARG( <NAME> ) - Option Argument String . . . . . . 97 7.6.14 OPT VALUE name - Option Argument Value . . . . . . . . . . 97 7.6.15 RESTART OPT( n ) - Resume Option Processing . . . . . . 97 7.6.16 SET OPT name - Force an option to be set . . . . . . . . . . . . 97 7.6.17 STACKCT OPT( <NAME> ) - Stacked Arg Count . . . . . . 98 7.6.18 STACKLST OPT( <NAME> ) - Argument Stack . . . . . . . 98 7.6.19 START OPT - Restart Option Processing . . . . . . . . . . . . . . 98 7.6.20 STATE OPT( <NAME> ) - Option State . . . . . . . . . . . . . . . 98 7.6.21 USAGE( exit-code ) - Usage invocation macro . . . . . . . . . . 99 7.6.22 VALUE OPT name - Option Flag Value . . . . . . . . . . . . . . . 99 7.6.23 VERSION - Version and Full Version. . . . . . . . . . . . . . . . . . . 99 7.6.24 WHICH IDX name - Which Equivalenced Index . . . . . . . . 99 7.6.25 WHICH OPT name - Which Equivalenced Option . . . . . 100 7.6.26 teOptIndex - Option Index and Enumeration . . . . . . . . . . 100 7.6.27 OPTIONS STRUCT VERSION - active version . . . . . . . 100 7.6.28 libopts External Procedures . . . . . . . . . . . . . . . . . . . . . . . . . . 100

7.6.28.1 ao string tokenize . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101 7.6.28.2 configFileLoad . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 102 7.6.28.3 optionFileLoad. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 102 7.6.28.4 optionFindNextValue . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 7.6.28.5 optionFindValue . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 7.6.28.6 optionFree . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104 7.6.28.7 optionGetValue . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104 7.6.28.8 optionLoadLine . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104 7.6.28.9 optionNextValue . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 105 7.6.28.10 optionOnlyUsage. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 105 7.6.28.11 optionProcess. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 106 7.6.28.12 optionRestore. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 106 7.6.28.13 optionSaveFile . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 107 7.6.28.14 optionSaveState. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 107 7.6.28.15 optionUnloadNested. . . . . . . . . . . . . . . . . . . . . . . . . . . . 107 7.6.28.16 optionVersion . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108 7.6.28.17 pathfind . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108 7.6.28.18 strequate . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 7.6.28.19 streqvcmp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 7.6.28.20 streqvmap . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 7.6.28.21 strneqvcmp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 110 7.6.28.22 strtransform . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 110

7.7 Option Descriptor File. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111 7.8 Using AutoOpts. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111

7.8.1 local-only use. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111 7.8.2 binary distro, AutoOpts not installed. . . . . . . . . . . . . . . . . . . 112 7.8.3 binary distro, AutoOpts pre-installed. . . . . . . . . . . . . . . . . . . 112 7.8.4 source distro, AutoOpts pre-installed . . . . . . . . . . . . . . . . . . . 112

7.8.5 source distro, AutoOpts not installed . . . . . . . . . . . . . . . . . . . 112 7.9 Configuring your program. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 113

7.9.1 configuration file presets. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 114 7.9.2 Saving the presets into a configuration file . . . . . . . . . . . . . . 114 7.9.3 Creating a sample configuration file . . . . . . . . . . . . . . . . . . . . 114 7.9.4 environment variable presets. . . . . . . . . . . . . . . . . . . . . . . . . . . 114 7.9.5 Config file only example . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 115

7.10 Configuration File Format . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 116 7.10.1 assigning a string value to a configurable . . . . . . . . . . . . . . 116 7.10.2 integer values . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 117 7.10.3 hierarchical values . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 117 7.10.4 configuration file sections. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 117 7.10.5 comments in the configuration file. . . . . . . . . . . . . . . . . . . . . 118

7.11 AutoOpts for Shell Scripts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 118 7.11.1 Parsing with an Executable . . . . . . . . . . . . . . . . . . . . . . . . . . 119 7.11.2 Parsing with a Portable Script . . . . . . . . . . . . . . . . . . . . . . . . 120

7.12 Automated Info Docs. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 135 7.12.1 “invoking” info docs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 136 7.12.2 library info docs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 136

7.13 Automated Man Pages . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 137 7.13.1 command line man pages. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 137 7.13.2 library man pages . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 137

7.14 Using getopt(3C). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 138 7.15 Internationalizing AutoOpts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 139 7.16 Naming Conflicts. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 140

8 Add-on packages for AutoGen . . . . . . . . . . . . 141

8.1 Automated Finite State Machine . . . . . . . . . . . . . . . . . . . . . . . . . . . 141 8.2 Combined RPC Marshalling. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 141 8.3 Automated Event Management. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 141 8.4 Invoking columns . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 142

8.4.1 columns usage help (-?) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 142 8.4.2 width option (-W) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 143 8.4.3 columns option (-c) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 143 8.4.4 col-width option (-w) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 143 8.4.5 spread option. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 143 8.4.6 indent option (-I) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 143 8.4.7 first-indent option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 143 8.4.8 tab-width option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 144 8.4.9 sort option (-s) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 144 8.4.10 format option (-f) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 144 8.4.11 separation option (-S) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 144 8.4.12 line-separation option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 144 8.4.13 by-columns option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 144 8.4.14 input option (-i) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 144

8.5 Invoking getdefs. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 145 8.5.1 getdefs usage help . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 146 8.5.2 defs-to-get option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 148 8.5.3 ordering option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 148 8.5.4 first-index option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 148 8.5.5 input option. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 148 8.5.6 subblock option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 148 8.5.7 listattr option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 149 8.5.8 filelist option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 149 8.5.9 assign option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 149 8.5.10 common-assign option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 149 8.5.11 copy option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 149 8.5.12 srcfile option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 150 8.5.13 linenum option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 150 8.5.14 output option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 150 8.5.15 autogen option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 150 8.5.16 template option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 150 8.5.17 agarg option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 150 8.5.18 base-name option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 151

8.6 Invoking xml2ag . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 152 8.6.1 xml2ag usage help (-?) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 152 8.6.2 output option (-O). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 154 8.6.3 templ-dirs option (-L) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 154 8.6.4 override-tpl option (-T) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 154 8.6.5 lib-template option (-l) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 154 8.6.6 base-name option (-b). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 154 8.6.7 definitions option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 154 8.6.8 load-scheme option (-S) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 154 8.6.9 load-functions option (-F) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 154 8.6.10 skip-suffix option (-s) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 155 8.6.11 select-suffix option (-o) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 155 8.6.12 source-time option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 155 8.6.13 no-fmemopen option (-m) . . . . . . . . . . . . . . . . . . . . . . . . . . . . 155 8.6.14 equate option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 155 8.6.15 writable option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 155 8.6.16 loop-limit option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 155 8.6.17 timeout option (-t) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 155 8.6.18 trace option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 155 8.6.19 trace-out option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 156 8.6.20 show-defs option. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 156 8.6.21 define option (-D) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 156 8.6.22 undefine option (-U) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 156

8.7 Replacement for Stdio Formatting Library . . . . . . . . . . . . . . . . . . 157

9 Some ideas for the future. . . . . . . . . . . . . . . . . 158

Concept Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 159

Function Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 162